### v1

Abstract
We study natural privacy filters, which enable the exact composition of differentially private (DP) mechanisms with adaptively chosen privacy characteristics. Earlier privacy filters consider only simple privacy parameters such as Rényi-DP or Gaussian DP parameters. Natural filters account for the entire privacy profile of every query, promising greater utility for a given privacy budget. We show that, contrary to other forms of DP, natural privacy filters are not free in general. Indeed, we show that only families of privacy mechanisms that are well-ordered when composed admit free natural privacy filters.
1 Introduction
Differential Privacy (DP) [dwork2006calibrating] is a rigorous notion of data privacy, used to protect against a host of different privacy attacks, including membership inference [homer2008resolving, dwork2015robust, carlini2022membership] and data reconstruction [dinur2003revealing, dwork2007price, carlini2021extracting, balle2022reconstructing, dick2023confidence, nasr2023scalable]. As such, there is a sustained push to integrate and deploy DP in a large number of applications [opendp-registry], from machine learning (ML) and artificial intelligence (AI) models [kairouz2021practical, sinha2025vaultgemma], to analytics collection [rogers2020linkedin, apple-emoji, ding2017collecting], to data releases [adeleye2023publishing, abowd20222020], as well as in general libraries and systems supporting those applications [gaboardi2020programming, wilson2020differentially, jax-privacy2022github, tholoniat2024cookie, ghazi2025differential]. This practical DP push also motivates important theoretical progress.
A prominent direction of theoretical improvement comes from alternative definitions of DP, that enable tighter privacy loss accounting under composition and improve the privacy-accuracy tradeoff. The most successful definitions include $(\varepsilon,\delta)$-DP [dwork2006our] (known as approximate DP), the Rényi divergence family of DP definitions such as zCDP [dwork2016concentrated, bun2016concentrated] and RDP [mironov2017renyi], and privacy definitions that generalize approximate DP by considering all $(\varepsilon,\delta)$-DP values compatible with a given mechanism, such as the privacy loss distribution (PLD) [sommer2019privacy] and $f$-DP [dong2022gaussian]. The latter definitions, PLD and $f$-DP, are theoretically appealing as they tightly capture resilience to membership inference attacks framed as hypothesis tests [3, dong2022gaussian]. However, composition under these definitions is more complex and can be computationally difficult, with analyses of specific algorithms that rely on numerical approaches to handle composition in practice [koskela2020computing, gopi2021numerical].
Two other theoretical improvements enable new forms of composition that are necessary requirements for practical, real world DP usage. First, concurrent composition [vadhan2021concurrent] supports composition of DP mechanisms running over multiple rounds of interaction with the user—such as the sparse vector technique [dwork2009complexity]—that are running in parallel. This is required in practical systems which typically have to handle queries from multiple parties simultaneously, and where it is not possible to enforce a sequential ordering of the queries. Even individual users may want to run DP computations concurrently with long-running, multi-step mechanism (e.g., [kostopoulou2023turbo]).
Second, adaptive composition, the main topic of this paper, enables the composition of DP mechanisms with adaptively chosen privacy parameters (as opposed to regular composition, in which these parameters have to be known in advance). Adaptive composition is formalized through the notion of privacy filters [RRUV16], which accept DP mechanisms as long as their composition remains within a bound fixed in advance (more details in §2). Such adaptive composition is necessary to support long running applications, in which analysts can adaptively query a system with mechanisms and DP parameters of their choice [lecuyer2019privacy, kostopoulou2023turbo, kuchler2024cohere]. Furthermore, filters enable crucial privacy accounting savings through advanced versions of parallel composition [lecuyer2019privacy], as well as individual (personalized) DP [ebadi2015differential], which enables data dependent privacy accounting in ML model training [feldman2021individual, yu2022individual] and measurement systems [tholoniat2024cookie, ghazi2025differential].
A key question in DP composition is whether these various improvements combine with each other. Ideally these more flexible forms of composition come for free: that is, that can we apply composition results under concurrent mechanisms or adaptive budgets using privacy filters, without a penalty in the final privacy loss after composition. Flexible composition for free is important not only because it is efficient privacy-wise, but also because it enables reusing all existing composition results, including tight analyses for specific mechanisms and empirical computational results.
As it turns out, concurrent composition is indeed free for all DP definitions [vadhan2023concurrent, lyu2022composition]. Moreover, concurrent and adaptive filter composition reduces to adaptive filters [haney2023concurrent], and thus has the same composition cost. Therefore, the full expressivity of both concurrency and adaptivity hinges on free privacy filter composition for different notions of DP. Filters, however, are not as well understood. We know that adaptive composition under privacy filters is free for $\varepsilon$-DP [RRUV16] and Rényi-based DP definitions [FZ21, lecuyer2021practical] (zCDP, RDP). Filters are also free for Gaussian mechanisms under $f$-DP and PLD accounting [ST22, KTH22]. However, for $(\varepsilon,\delta)$-DP, the best known result for adaptive composition with filters is only asymptotically free, with worse constants than non-adaptive composition [WRRW23]. Even less is known about filters under $f$-DP and PLD. In particular, we consider natural filters, which leverage lossless composition of the privacy characteristics of individual queries. Natural filters can be realized via PLDs, privacy profiles, or $f$-DP accounting. In this paper, we make three important contributions towards fully understanding natural privacy filters:
  • First, in §3 we present a general theory of adaptive mechanisms that issue adversarial queries with a constrained per-query privacy cost.
  • In §4 we apply this theory to the natural filter and, surprisingly, we show that adaptive composition is NOT free under the PLD nor $f$-DP notions of privacy (see Corollary˜1).
  • Finally, we show in §6 that filters are only free for very restricted families of mechanisms: those that are closed under composition, and admit totally ordered tradeoff functions.
Our work also includes some minor contributions including that PLDs can be assigned a natural ordering satisfying a completeness property enabling one to take a supremum of PLDs (see Proposition˜11). As far as we are aware this property has not been documented in the privacy accounting literature. We alo document in §5 some families of queries and budgets that do and do not admit free filters.
2 Preliminaries
Differential privacy is the study of randomized algorithms (called mechanisms) $M:\mathcal{D}\to\mathcal{Y}$ where $\mathcal{D}$ denotes some space of datasets containing sensitive personal information. Differential privacy requires that an adversary should not gain much statistical advantage for inferring whether the mechanism was run on dataset $D_{1}\in\mathcal{D}$ or a second dataset $D_{2}\in\mathcal{D}$ when the datasets are close to each other in some sense. That is, the distributions of $M(D_{1})$ and $M(D_{2})$ should be difficult to distinguish from one another in the hypothesis testing sense and this should hold uniformly for any close datasets. The measure of closeness for datasets is generally defined by a neighbouring relation but this turns out not to be relevant to our work. In fact, our conclusions hold for an arbitrary pair of datasets. For this reason, we fix throughout a single arbitrary pair of datasets $(D_{1},D_{2})\in\mathcal{D}\times\mathcal{D}$ and view the privacy characteristics of a mechanism $M$ in terms of the hypothesis testing problem $M(D_{1})$ vs. $M(D_{2})$.
In our work we reason about the composition of adaptive mechanisms, which are permitted to inspect outcomes of any previously run mechanism. It is enough to define adaptivity for just two rounds of interaction as we can define more complex adaptively composed mechanisms inductively.
Definition 1.
Let $M_{1}:\mathcal{D}\to\mathcal{Y}_{1}$ be a mechanism and let $M_{2}:\mathcal{D}\times\mathcal{Y}_{1}\to\mathcal{Y}_{2}$ be an adaptive mechanism. The adaptive composition of $M_{1}$ and $M_{2}$ is the mechanism
$(M_{1}\otimes M_{2})(D):=(Y_{1},Y_{2}),Y_{1}\sim M_{1}(D),Y_{2}\sim M_{2}(D;Y_{1}).$
It is natural in general to view the privacy characteristics of any private mechanism as a hypothesis testing problem [3, DRS22]. As such it is desirable to furnish hypothesis testing problem $(P,Q)$ with some kind of ordering so that we can compare the privacy guarantees offered by various mechanisms. We also require an ordering so that we can make sense of privacy filters. Fortunately testing pairs can be equipped with a very natural order, namely the Blackwell order.
Definition 2.
For a pair of distributions $(P,Q)$ defined on a common probability space $\Omega$ and a second pair $(P^{\prime},Q^{\prime})$ defined on $\Omega^{\prime}$, we say that $(P^{\prime},Q^{\prime})$ dominates $(P,Q)$ in the Blackwell order, written $(P,Q)\preceq(P^{\prime},Q^{\prime})$, if we can find a Markov kernel $\phi$ from $\Omega^{\prime}$ to $\Omega$ such that $P=\phi P^{\prime}$ and $Q=\phi Q^{\prime}$. We write $(P,Q)\prec(P^{\prime},Q^{\prime})$ if $(P,Q)\preceq(P^{\prime},Q^{\prime})$ but not vice-versa.
Informally, the Blackwell order says that more information is available for distinguishing $P^{\prime}$ and $Q^{\prime}$ compared to $P$ vs $Q$. For mechanisms, we can interpret this as $(M(D_{1}),M(D_{2}))\preceq(M^{\prime}(D_{1}),M^{\prime}(D_{2}))$ when $M$ can be reconstructed from $M^{\prime}$ by introducing additional randomness through the kernel $\phi$. Note that the Blackwell order is a partial order.
Instead of a carrying around a pair of distributions to represent the privacy characteristics of a mechanism, we can consolidate them into the privacy loss distribution (PLD) [2]. As we will see the PLD possesses very desirable composition properties and is a practical basis for privacy loss computations. To formalize the PLD, we require likelihood ratios. Recall that, for a distribution $P$ absolutely continuous with respect to a distribution $Q$ ($P\ll Q$) on a probability domain $\Omega$, their likelihood ratio is the Radon-Nikodym derivative $\ell(\omega):=\frac{dP}{dQ}(\omega)$. However, in some cases, we would like to consider private mechanisms that violate absolute continuity and still construct the PLD. This can be achieved by taking any measure $\nu$ on $\Omega$ such that $P,Q\ll\nu$ (e.g. $\nu:=P+Q$) and setting $\ell(\omega):=\frac{dP}{d\nu}(\omega)/\frac{dQ}{d\nu}(\omega)$ ($\cdot/0:=\infty$). This construction of the likelihood ratio is unique only outside of a null-measure set [^1] More carefully, assume $P,Q\ll\nu,\nu^{\prime}$ and set $\mu:=\nu+\nu^{\prime}$ so that $P,Q\ll\nu,\nu^{\prime}\ll\mu$. By Radon–Nikodym and the chain rule for measures, we have, for $\omega\sim P$, that $\frac{dP}{d\nu}(\omega)/\frac{dQ}{d\nu}(\omega)=\frac{dP}{d\mu}(\omega)/\frac{dQ}{d\mu}(\omega)=\frac{dP}{d\nu^{\prime}}(\omega)/\frac{dQ}{d\nu^{\prime}}(\omega)$ with probability $1$., which is enough to construct a unique PLD. For this reason, we denote by $\frac{dP}{dQ}$ the likelihood ratio of $P$ with respect to $Q$ even when $P\not\ll Q$.
Definition 3.
Let $P$ and $Q$ be distributions on $\Omega$. The PLD of $(P,Q)$, denoted $\operatorname{PLD}(P\parallel Q)$, is the distribution over $\mathbb{R}\cup\{\infty\}$ of $\log(\frac{dP}{dQ}(\omega)),\omega\sim P$.
For shorthand, we denote by $\operatorname{PLD}(M):=\operatorname{PLD}(M(D_{1})\parallel M(D_{2}))$ the PLD of a mechanism $M$. Note that a distribution can be easily recognized as a PLD as follows.
Proposition 1.
A distribution $L$ on $\mathbb{R}\cup\{\infty\}$ is the PLD of some $(P,Q$) if and only if it satisfies $\mathop{\mathbb{E}}_{Z\sim L}[e^{-Z}]\leq 1$. In this case, one such pair is $(L,L^{\prime})$ where $L^{\prime}$ is the Esscher tilt of $L$, namely $dL^{\prime}(z):=e^{-z}dL(z)$ for $z\in\mathbb{R}$ and $L^{\prime}(\{-\infty\}):=1-\mathop{\mathbb{E}}_{Z\sim L}[e^{-Z}]$.
PLDs also inherit the Blackwell order: For PLDs $L$ and $L^{\prime}$, we say that $L\preceq L^{\prime}$ if there exist pairs $(P,Q)$ and $(P^{\prime},Q^{\prime})$ such that $L:=\operatorname{PLD}(P\parallel Q)$, $L^{\prime}=\operatorname{PLD}(P^{\prime}\parallel Q^{\prime})$, and $(P,Q)\preceq(P^{\prime},Q^{\prime})$. It is straightforward to show that this order is well-defined regardless of the underlying representations $(P,Q)$ and $(P^{\prime},Q^{\prime})$.
PLDs also form a monoid under convolution that is monotone in the Blackwell order.
Proposition 2.
For any PLDs $L_{1},L_{2},L_{3},L^{\prime}_{1},L^{\prime}_{2}$,
  1. 1.
    The convolution $L_{1}\oplus L_{2}$ is also a PLD;
  2. 2.
    The identically zero distribution Id is a PLD such that $L_{1}\oplus\text{Id}=\text{Id}\oplus L_{1}=L_{1}$;
  3. 3.
    $(L_{1}\oplus L_{2})\oplus L_{3}=L_{1}\oplus(L_{2}\oplus L_{3})$; and
  4. 4.
    If $L_{1}\preceq L^{\prime}_{1}$ and $L_{2}\preceq L^{\prime}_{2}$, then $L_{1}\oplus L_{2}\preceq L^{\prime}_{1}\oplus L^{\prime}_{2}$.
A highly prized property of PLDs is that the privacy characteristics of an adaptive mechanism corresponds to convolution of PLDs, provided that each component fixes a privacy budget in advance. This is important because the convolution of a long sequence of distributions can be computed very efficiently by applying the fast Fourier transform [koskela2020computing]. Note that we will require more than this to understand fully adaptive composition where the privacy bound of each component is itself adaptive.
Proposition 3.
Let $M_{1}:\mathcal{D}\to\mathcal{Y}_{1}$ and $M_{2}:\mathcal{D}\times\mathcal{Y}_{1}\to\mathcal{Y}_{2}$ be adaptive mechanisms such that $\operatorname{PLD}(M_{1})\preceq L_{1}$ and $\operatorname{PLD}(M_{2}(\cdot;y_{1}))\preceq L_{2}$ for every $y_{1}\in\mathcal{Y}_{1}$. Then $\operatorname{PLD}(M_{1}\otimes M_{2})\preceq L_{1}\oplus L_{2}$. Moreover, if $\operatorname{PLD}(M_{1})\preceq\operatorname{PLD}(M^{\prime}_{1})$ and $\operatorname{PLD}(M_{2}(\cdot;y_{1}))\preceq\operatorname{PLD}(M^{\prime}_{2}(\cdot;y_{1}))$ for every $y_{1}\in\mathcal{Y}_{1}$, then $\operatorname{PLD}(M_{1}\otimes M_{2})\preceq\operatorname{PLD}(M^{\prime}_{1}\otimes M^{\prime}_{2})$.
Closely related to the PLD is the hockey-stick curve: a convex reparameterization of the privacy profile of a mechanism generalizing the classical notion of $(\varepsilon,\delta)$-DP.
Definition 4.
For a pair of distributions $(P,Q)$ over $\Omega$, their hockey-stick divergence at order $x\in\mathbb{R}^{\times}:=(0,\infty)$ is
$H_{x}(P\parallel Q):=\sup_{E\subseteq\Omega}P(E)-xQ(E).$
Again, for convenience, we write $H_{x}(M):=H_{x}(M(D_{1})\parallel M(D_{2}))$ and sometimes $h_{M}(x):=H_{x}(M)$ for the hockey-stick curve of a mechanism $M$. We will rely on hockey-stick curves in order to reason about adversarial behaviour of adaptive mechanisms. The hockey-stick curve is very closely related to the classical $(\varepsilon,\delta)$ form of differential privacy. In particular, by construction of the hockey-stick divergence a mechanism $M$ satisfies $(\varepsilon,\delta)$-DP exactly when $H_{e^{\varepsilon}}(M)\leq\delta$. In general, hockey-stick curves are characterized by a few simple conditions.
Proposition 4 ([4] Lemma 9).
A curve $h:\mathbb{R}^{\times}\to[0,1]$ is a hockey-stick curve of some pair $(P,Q)$ if and only if $h$ is convex and decreasing such that $\lim_{x\to 0}h(x)=1$ and $h(x)\geq 1-x$ for all $x\in\mathbb{R}^{\times}$.
Hockey-stick curves are also closely related to PLDs and can be fully recovered from the PLD via the following formula.
Proposition 5.
For any mechanism $M$ and $x\in\mathbb{R}^{\times}$,
$h_{M}(x)=\mathop{\mathbb{E}}_{Z\sim\operatorname{PLD}(M)}[(1-xe^{-Z})_{+}]$
where $(t)_{+}:=\max\{0,t\}$.
Proof.
For any pair $(P,Q)$ defined on $\Omega$ with likelihood ratio $\ell$, we have that
$H_{x}(P\parallel Q)$ $=\sup_{E\subseteq\Omega}P(E)-xQ(E)$
$=\sup_{E\subseteq\Omega}\int_{E}\frac{d(P-xQ)}{dP}(\omega)\,dP(\omega)$
$=\sup_{E\subseteq\Omega}\int_{E}1-x/\frac{dP}{dQ}(\omega)\,dP(\omega)$
$=\int_{\Omega}\left(1-x/\frac{dP}{dQ}(\omega)\right)_{+}\,dP(\omega)$
$=\int_{\Omega}(1-xe^{-\log(\ell(\omega))})_{+}\,dP(\omega)$
$=\mathop{\mathbb{E}}_{Z\sim\operatorname{PLD}(P\parallel Q)}[(1-xe^{-Z})_{+}].$
For convenience, we will sometimes write $h_{L}$ to denote the unique hockey-stick curve associated to a PLD $L$, namely $h_{L}(x):=\mathop{\mathbb{E}}_{Z\sim L}[(1-xe^{-Z})_{+}]$. Moreover, like PLDs, hockey-stick curves also capture adaptive composition in a natural way.
Proposition 6.
Let $M_{1}:\mathcal{D}\to\mathcal{Y}_{1}$ and $M_{2}:\mathcal{D}\times\mathcal{Y}_{1}\to\mathcal{Y}_{2}$ be adaptive mechanisms and let $\ell^{1}:=\frac{dM_{1}(D_{1})}{dM_{1}(D_{2})}$ denote the likelihood ratio for $M_{1}$. Then
$h_{M_{1}\otimes M_{2}}(x)=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[h_{M_{2}(\cdot;Y_{1})}(x/\ell^{1}(Y_{1}))].$
Proof.
Let $\ell$ denote the likelihood ratio for $M_{1}\otimes M_{2}$ and let $\ell^{2}_{y_{1}}$ denote the likelihood ratio for $M_{2}(\cdot;y_{1})$. By Bayes’ rule, we have $\ell(y_{1},y_{2})=\ell^{1}(y_{1})\cdot\ell^{2}_{y_{1}}(y_{2})$, so we have
$\operatorname{PLD}(M_{1}\otimes M_{2})\equiv\log(\ell^{1}(Y_{1}))+\log(\ell^{2}_{Y_{1}}(Y_{2})),Y_{1}\sim M_{1}(D_{1}),Y_{2}\sim M_{2}(D_{1};Y_{1}).$
By Proposition˜5 and the law of total expectation, it follows that
$h_{M_{1}\otimes M_{2}}(x)$ $=\mathop{\mathbb{E}}_{Z\sim\operatorname{PLD}(M_{1}\otimes M_{2})}[(1-xe^{-Z})_{+}]$
$=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[\mathop{\mathbb{E}}_{Y_{2}\sim M_{2}(D_{1};Y_{1})}[(1-xe^{-(\log(\ell^{1}(Y_{1}))+\log(\ell^{2}_{Y_{1}}(Y_{2})))})_{+}]]$
$=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[\mathop{\mathbb{E}}_{Y_{2}\sim M_{2}(D_{1};Y_{1})}[(1-x/\ell^{1}(Y_{1})\cdot e^{-\log(\ell^{2}_{Y_{1}}(Y_{2}))})_{+}]]$
$=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[\mathop{\mathbb{E}}_{Z_{2}\sim\operatorname{PLD}(M_{2}(\cdot;Y_{1})}[(1-x/\ell^{1}(Y_{1})\cdot e^{-Z_{2}})_{+}]]$
$=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[h_{M_{2}(\cdot;Y_{1})}(x/\ell^{1}(Y_{1}))].$
Our last characterization of privacy is given the Type I/Type II error tradeoff curve, also known as the $f$-DP framework of privacy [1].
Definition 5.
Let $(P,Q)$ be any distributions on $\Omega$. For any hypothesis test $\phi:\Omega\to\{P,Q\}$ for distinguishing $P$ and $Q$, we call $\alpha_{\phi}:=\mathbb{P}_{\omega\sim P}[\phi(\omega)=Q]$ its Type I error and $\beta_{\phi}:=\mathbb{P}_{\omega\sim Q}[\phi(\omega)=P]$ its Type II error. The tradeoff curve for $(P,Q)$ is defined as
$T_{\alpha}(P\parallel Q):=\inf_{\phi}\{\beta_{\phi}:\alpha_{\phi}\leq\alpha\}.$
For convenience, we sometimes write the tradeoff curve of a mechanism $M$ with respect to the fixed dataset pair $(D_{1},D_{2})$ as $T_{\alpha}(M):=T_{\alpha}(M(D_{1})\parallel M(D_{2}))$ or sometimes as $\tau_{M}(\alpha):=T_{\alpha}(M)$ where appropriate. In general, we can characterize valid tradeoff curves as follows.
Proposition 7 ([1] Proposition 2.2).
A curve $\tau:[0,1]\to[0,1]$ is a tradeoff curve for some testing problem $(P,Q)$ exactly when $\tau$ is continuous, convex, decreasing, and $\tau(\alpha)\leq 1-\alpha$ for all $\alpha\in[0,1]$.
Like hockey-stick curves, it is enough to understand the PLD of a testing problem in order to obtain its tradeoff curve. The following formula can be derived by the well-known Neyman–Pearson lemma.
Proposition 8.
Let $(P,Q)$ be a pair of distributions with tradeoff curve $\tau(\alpha):=T_{\alpha}(P\parallel Q)$ and let $F$ denote the CDF of $\operatorname{PLD}(P\parallel Q)$. Then
$\tau(\alpha)=\mathbb{E}_{Z\sim\operatorname{PLD}(P\parallel Q)}[1(F(Z)>\alpha)\cdot e^{-Z}]$
It is common to represent a pair of distributions by either their hockey-stick curve or their tradeoff function. We show that there is a natural link between these representations via inversion and convex conjugacy. For a tradeoff function $\tau$ of a testing problem $(P,Q)$, we write $\tau^{-1}(\beta):=\inf\{\alpha\in[0,1]:\tau(\alpha)\leq\beta\}$ for its inverse, namely the tradeoff curve for the testing problem $(Q,P)$. For a convex function $g$, we denote its convex conjugate by $g^{*}(y):=\sup_{x\in\mathbb{R}}xy-g(x)$. Note that we can extend tradeoff functions to the real-line by setting them to $\infty$ outside their support $[0,1]$. Note that a special case of the following formula appears in [1] (Prop 2.12) for symmetric tradeoff functions but this generalization is new to the best of our knowledge.
Proposition 9.
Let $(P,Q)$ be a pair of distributions with tradeoff function $\tau(\alpha):=T_{\alpha}(P\parallel Q)$ and hockey-stick curve $h(x):=H_{x}(P\parallel Q)$. Then
$h(x)=1+(\tau^{-1})^{*}(-x)$
Proof.
First, let $L$ denote the PLD of $(P,Q)$ and let $F$ be its CDF. By Proposition˜8,
$\tau(\alpha)=\mathbb{E}_{Z\sim L}[1(Z>F^{-1}(\alpha))\cdot e^{-Z}].$
Therefore, by Proposition˜5
$h(x)$ $=\mathbb{E}_{Z\sim L}[(1-xe^{-Z})_{+}]$
$=\mathbb{E}_{Z\sim L}[1(Z>\log(x))\cdot\underbrace{(1-xe^{-Z})}_{>0\iff Z>\log(x)}]$
$=\sup_{\varepsilon\in\mathbb{R}}\mathbb{E}_{Z\sim L}[1(Z>\varepsilon)\cdot(1-xe^{-Z})]$
$=\sup_{\alpha\in[0,1]}\mathbb{E}_{Z\sim L}[1(Z>F^{-1}(\alpha))\cdot(1-xe^{-Z})]$
$=\sup_{\alpha\in[0,1]}1-\alpha-x\tau(\alpha)$
$=\sup_{\alpha^{\prime}\in[0,1]}1-\tau^{-1}(\alpha^{\prime})-x\alpha^{\prime}$ ($\alpha^{\prime}=\tau^{-1}(\alpha)$)
$=1+(\tau^{-1})^{*}(-x).$
It follows as a natural consequence that the Blackwell order on pairs, the ordering induced by the tradeoff function, as well as the hockey-stick induced ordering are all equivalent.
Proposition 10.
Let $(P,Q)$ and $(P^{\prime},Q^{\prime})$ be pairs of distributions with tradeoff functions $\tau$ and $\tau^{\prime}$ respectively as well as hockey-stick curves $h$ and $h^{\prime}$ respectively. The following are equivalent
  1. (i)
    $(P,Q)\preceq(P^{\prime},Q^{\prime})$
  2. (ii)
    $\tau\succeq\tau^{\prime}$
  3. (iii)
    $h\preceq h^{\prime}$
The equivalence of (i) and (ii) is given by a celebrated theorem of Blackwell [Bla51, Theorem 10]. As for the equivalence of (ii) and (iii), this now follows by applying Fenchel–Moreau duality to Proposition˜9 and recalling that convex conjugation is an order-reversing operation.
As a consequence, it turns out that PLDs endowed with the Blackwell order possess the same remarkable completeness property that enables analysis in the real numbers.
Proposition 11.
For any non-empty family of PLDs $\mathcal{L}$ dominated by at least one PLD, there exists a unique PLD $\sup\mathcal{L}$ that dominates $\mathcal{L}$ but is dominated by every upper bound for $\mathcal{L}$. In this case,
  1. (i)
    $h_{\sup\mathcal{L}}(x)=\sup\{h_{L}(x):L\in\mathcal{L}\}$
  2. (ii)
    $\tau_{\sup\mathcal{L}}=\operatorname{conv}\{\tau_{L}:L\in\mathcal{L}\}$
where $(\operatorname{conv}{\mathcal{F}})(\alpha):=\sup\{f(\alpha):f\text{ is convex},f\preceq\mathcal{F}\}$ denotes the lower convex envelope.
This result follows immediately from Proposition˜10 and from noticing that the properties characterizing hockey-stick curves (see Proposition˜4) are all closed under pointwise suprema. Part (ii) follows from noticing that all of the properties of tradeoff curves (see Proposition˜7) are also closed under the lower convex envelope operation.
We rely on the supremum extensively in order to reason about natural privacy filters. In particular, the former characterization says that the supremum of PLDs is given by a pointwise supremum in the space of hockey-stick curves, which will be particularly useful for showing our main result. As far as we are aware, the supremum property of PLDs has not actually been documented in the literature.
3 Query-Restricted Adversaries
In this section we develop a general theory of the privacy characteristics of adversarial adaptive mechanism when the queries the adversary may issue are restricted. This will lead directly to a theory of natural filters as we will see in §4.
Definition 6.
Let $k$ be a positive integer. We define a privacy rule of length $k$ as a map $\Gamma$ that takes a sequence of PLDs $(L_{1},\dots,L_{k^{\prime}})$ ($0\leq k^{\prime}<k$) and outputs some collection of PLDs including Id. We define a $\Gamma$-adversary as an adaptively composed mechanism $M=M_{1}\otimes\dots\otimes M_{k}$ such that
$L_{i}:=\operatorname{PLD}(M_{i}(\cdot;y_{1},\dots,y_{i-1}))\in\Gamma(L_{1},\dots,L_{i-1})$
for every $(y_{1},\dots,y_{k-1})\in\mathcal{Y}_{1}\times\dots\times\mathcal{Y}_{k-1}$. We will write $\operatorname{Adv}(\Gamma)$ to denote the set of $\Gamma$-adversaries and we will denote by
$\operatorname{PLD}(\Gamma):=\sup\{\operatorname{PLD}(M):M\in\operatorname{Adv}(\Gamma)\}$
the worst-case privacy loss of arbitrary $\Gamma$-adversaries.
It turns out that the maximum privacy loss for an adversary respecting a rule $\Gamma$ occurs in some sense when the adversary first plays a move and then adversarially chooses a optimal strategy tuned to ensure that the overall strategy is optimal at a particular given point. Note that by $\lambda$ we mean the empty sequence.
Lemma 1.
Let $\Gamma$ be a rule of length $k$. If $k=1$, then $\operatorname{PLD}(\Gamma)=\sup{\Gamma(\lambda)}$ and, for $k>1$, we have
$\operatorname{PLD}(\Gamma)=\sup\{L\oplus\operatorname{PLD}(\Gamma_{L}):L\in\Gamma(\lambda)\}$
where $\Gamma_{L_{0}}(L_{1},\dots,L_{k^{\prime}}):=\Gamma(L_{0},L_{1},\dots L_{k^{\prime}})$ denotes the rule of length of k1k-1 that fixes the first move played against $\Gamma$ to $L_{0}$.
Proof.
The case where $k=1$ is immediate, so we assume $k>1$. We first show that
$\operatorname{PLD}(\Gamma)\preceq\sup\{L\oplus\operatorname{PLD}(\Gamma_{L}):L\in\Gamma(\lambda)\}.$
To that end, consider any $\Gamma$-adversary $M$. Decompose $M$ as the adaptive composition of $M_{1}(\cdot)$ with some $M^{\prime}(\cdot;y_{1})$ and let $L:=\operatorname{PLD}(M_{1})\in\Gamma(\lambda)$. Then, clearly, for every fixed $y_{1}$, $M^{\prime}(\cdot;y_{1})$ is a $\Gamma_{L}$-adversary and, in particular, $\operatorname{PLD}(M^{\prime}(\cdot;y_{1}))\preceq\operatorname{PLD}(\Gamma_{L})$. By Proposition˜3, we have that $\operatorname{PLD}(M)\preceq L\otimes\operatorname{PLD}(\Gamma_{L})$ and the desired inequality follows by taking suprema.
More surprising is the reverse inequality, which we prove constructively by passing to hockey-stick curves. Let $L\in\Gamma(\lambda)$, let $\kappa\in\mathbb{R}^{\times}$, and let $\varepsilon>0$. Choose any mechanism $M_{1}$ that has PLD $L$ and let $\ell:=\frac{dM_{1}(x_{1})}{dM_{1}(x_{2})}$ denote the corresponding likelihood function. Moreover, for any $\kappa\in\mathbb{R}^{\times}$ we can choose by Proposition˜11, a $\Gamma_{L}$-adversary $M^{\prime}_{\kappa}$ such that
$H_{\kappa}(M^{\prime}_{\kappa})\geq\sup_{M^{\prime\prime}\in\operatorname{Adv}(\Gamma_{L})}H_{\kappa}(M^{\prime\prime})-\varepsilon.$
Now consider the $\Gamma$-adversary
$M(x):=(Y_{1},Y^{\prime}),Y_{1}\sim M_{1}(x),Y^{\prime}\sim M^{\prime}_{\kappa/\ell(Y_{1})}(x).$
By Proposition˜6, we get
$h_{M}(\kappa)$ $=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(x_{1})}[h_{M^{\prime}_{\kappa/\ell(Y_{1})}}(\kappa/\ell(Y_{1}))]$
$\geq\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(x_{1})}[\sup_{M^{\prime\prime}\in\operatorname{Adv}(\Gamma_{L})}h_{M^{\prime\prime}}(\kappa/\ell(Y_{1}))-\varepsilon]$
$=(h_{M_{1}}\otimes\sup_{M^{\prime\prime}\in\operatorname{Adv}(\Gamma_{L})}h_{M^{\prime\prime}})(\kappa)-\varepsilon.$
Again, taking suprema, we get that
$\sup_{M\in\operatorname{Adv}(\Gamma)}h_{M}\succeq h_{M_{1}}\otimes\sup_{M^{\prime\prime}\in\operatorname{Adv}(\Gamma_{L})}h_{M^{\prime\prime}},$
and thus $\operatorname{PLD}(\Gamma)\succeq L\otimes\operatorname{PLD}(\Gamma_{L})$, which completes the proof since $L$ was arbitrary. ∎
Intuitively, this result says that a worst-case adversary plays a move with corresponding PLD $L$, observes the outcome $Y_{1}$, and adversarially chooses a followup in $\operatorname{Adv}(\Gamma_{L})$ to ensure that the overall process “traces out” the privacy curve of $\operatorname{Adv}(\Gamma_{L})$.
4 Characterizing the Natural Filter
Our goal is to study when the natural filter Algorithm˜1 with budget $B$ and queries in $\mathcal{L}$ comes at no additional privacy cost, i.e. the mechanism induced by the interaction between the filter and an adversarial analyst itself has privacy bounded by $B$. In this case, we say that the natural filter is free. We will write $L\ominus L^{\prime}:=\sup\{L^{\prime\prime}:L^{\prime}\oplus L^{\prime\prime}\preceq L\}$ so that $L^{\prime\prime}\preceq L\ominus L^{\prime}$ exactly when $L^{\prime}\oplus L^{\prime\prime}\preceq L$. We also write $\mathcal{L}_{\preceq L}:=\{L^{\prime}\in\mathcal{L}:L^{\prime}\preceq L\}$ to denote the queries in $\mathcal{L}$ dominated by $L$. For simplicity, we will assume that $\text{Id}\in\mathcal{L}$.
Algorithm 1 NaturalFilter
0: Privacy budget $L$, adversary (or analyst) $\mathcal{A}$, family of allowed PLDs $\mathcal{L}$, query capacity $k$
$i=0$
while $\mathcal{A}$ continues and $i<k$ do
  $i\leftarrow i+1$
  $\mathcal{A}$ gives mechanism $\mathcal{M}_{i}$ that has PLD $L_{i}\in\mathcal{L}$ $\triangleright$ can depend on previous results $Y_{<i}$
  if $L_{1}\oplus\dots\oplus L_{i}\preceq L$ then
   $\mathcal{A}$ receives $Y_{i}\sim\mathcal{M}_{i}$
  end if
end while
return $(Y_{1},\ldots,Y_{i})$
Now, notice that the natural filter Algorithm˜1 with budget $L$ and $k$ queries in $\mathcal{L}$ is free exactly when $\operatorname{PLD}(\operatorname{Filter}_{\mathcal{L},B,k})\preceq B$ where
$\operatorname{Filter}_{\mathcal{L},B,k}(L_{1},\dots,L_{k^{\prime}}):=\mathcal{L}_{\preceq B\ominus L_{1}\dots\ominus L_{k^{\prime}}}=\{L_{k^{\prime}+1}\in\mathcal{L}:L_{1}\oplus\dots\oplus L_{k^{\prime}}\oplus L_{k^{\prime}+1}\preceq B\}$
for $0\leq k^{\prime}<k$. In particular, by applying Lemma˜1 to the rule $\operatorname{Filter}_{\mathcal{L},B,k}$ we immediately get the following characterization.
Theorem 1.
Let $\mathcal{L}$ be a family of PLDs, let $B$ be a PLD, and let $k$ be a positive integer. For any $L_{1},\dots,L_{k^{\prime}}\in\mathcal{L}$ ($0\leq k^{\prime}<k$) such that $L_{1}\oplus\dots\oplus L_{k^{\prime}}\preceq B$, let
$V_{k-k^{\prime}}(L_{1},\dots,L_{k^{\prime}}):=\sup\{L_{k^{\prime}+1}\oplus V_{k-(k^{\prime}+1)}(L_{1},\dots,L_{k^{\prime}+1}):L_{k^{\prime}+1}\in\mathcal{L}_{B\ominus L_{1}\ominus\dots\ominus L_{k^{\prime}}}\}$
where $V_{0}(L_{1},\dots,L_{k}):=\text{Id}$. The natural filter algorithm˜1 for $k$ queries in $\mathcal{L}$ with budget $B$ is free if and only if $V_{k}(\lambda)\preceq B$.
Informally, $V_{k-k^{\prime}}(L_{1},\dots,L_{k^{\prime}})$ is the maximal remaining privacy cost of an adversary that has already played moves $L_{1},\dots,L_{k^{\prime}}$. Naturally, when the adversary has already selected all of its moves, the remaining privacy cost is Id, i.e. nothing.
It is also of interest to understand when a given family of queries $\mathcal{L}$ simultaneously admits free natural filters against an arbitrary budget $B$. We show that this is the case exactly when the supremum commutes with the convolution operator.
Theorem 2.
Fix a family of PLDs $\mathcal{L}$. The natural filter Algorithm˜1 is free for every budget $B$ simultaneously and query capacity $k$ if and only if $\mathcal{L}$ has the following property:
$\forall L\in\mathcal{L}\ \textrm{and}\ \mathcal{L}^{\prime}\subseteq\mathcal{L}^{<k}\quad L\oplus\sup{\mathcal{L}^{\prime}}=\sup\{L\oplus L^{\prime}:L^{\prime}\in\mathcal{L}^{\prime}\}.$
Proof.
First, assume that the natural filter is free for every budget $B$ and for any query capacity. Now, let $L\in\mathcal{L}$ and $\mathcal{L}^{\prime}\subseteq\mathcal{L}^{<k}$. For every $L^{\prime}\in\mathcal{L}^{\prime}$, we have $L\oplus L^{\prime}\preceq L\oplus L\oplus\sup{\mathcal{L}^{\prime}}$ by Proposition˜2 and hence $\sup\{L\oplus\mathcal{L}^{\prime}\}\preceq L\oplus\sup{\mathcal{L}^{\prime}}$. As for the reverse inequality, consider the rule
$\Gamma(\lambda):=\{L\}\quad\quad\Gamma(L):=\mathcal{L}^{\prime}$
of length $2$ and the budget $B:=\sup(L\oplus\mathcal{L}^{\prime})$. Now, for any $M\in\operatorname{Adv}(\Gamma)$, notice that we can transform $M$ into $M^{\prime}\in\operatorname{Adv}(\operatorname{Filter}_{\mathcal{L},B,k})$ such that $\operatorname{PLD}(M)=\operatorname{PLD}(M^{\prime})$. Indeed, let $M=M_{1}\oplus M_{2}\in\operatorname{Adv}(\Gamma)$ and fix $y_{1}\in\mathcal{Y}_{1}$ so that we have that $\operatorname{PLD}(M_{2}(\cdot;y_{1}))=L_{1}\oplus\dots\oplus L_{k-1}\in\mathcal{L}^{<k}$. For each $L_{i}$, we can find a mechanism $M^{i}$ with $\operatorname{PLD}(M^{i})=L_{i}$. Now, set $M^{\prime}_{2}(\cdot;y_{1}):=M^{1}\otimes\dots\otimes M^{k-1}$ so that $\operatorname{PLD}(M^{\prime}_{2}(\cdot;y_{1}))=L_{1}\oplus\dots\oplus L_{k^{\prime}}=\operatorname{PLD}(M_{2}(\cdot;y_{1}))$. Clearly, $M^{\prime}:=M_{1}\otimes M^{\prime}_{2}\in\operatorname{Adv}(\operatorname{Filter}_{\mathcal{L},B,k})$ and, by Proposition˜3, we have $\operatorname{PLD}(M)=\operatorname{PLD}(M^{\prime})$, so by Lemma˜1 and since the natural filter is free for $B$ by assumption, we have
$L\oplus\sup{\mathcal{L}^{\prime}}=\sup_{M\in\operatorname{Adv}(\Gamma)}\operatorname{PLD}(M)\preceq\sup_{M^{\prime}\in\operatorname{Adv}(\operatorname{Filter}_{\mathcal{L},B,k})}\operatorname{PLD}(M^{\prime})\preceq B=\sup\{L\oplus\mathcal{L}^{\prime}\}.$
On the other hand, assume that convolution commutes with supremum in $\mathcal{L}^{\infty}$. Let $B$ be any budget and let $k$ be a positive integer. With an eye toward Theorem˜1, we claim that the maximal remaining privacy cost of an adversary that has selected some if its moves already satisfies
$V_{k-k^{\prime}}(L_{1},\dots,L_{k^{\prime}})=\sup\{L_{k^{\prime}+1}\oplus\dots\oplus L_{k}:L_{k^{\prime}+1},\dots,L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
for $0\leq k^{\prime}<k$. Indeed,
$V_{1}(L_{1},\dots,L_{k-1})$ $=\sup\{L_{k}\oplus V_{0}(L_{1},\dots,L_{k}):L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
$=\sup\{L_{k}:L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
and, if the claim holds for $k^{\prime}$, then
$V$ $_{k-(k^{\prime}-1)}(L_{1},\dots,L_{k^{\prime}-1})$
$=\sup\{L_{k^{\prime}}\oplus V_{k-k^{\prime}}(L_{1},\dots,L_{k^{\prime}}):L_{k^{\prime}}\in\mathcal{L},\oplus_{i=1}^{k^{\prime}}L_{i}\preceq B\}$
$=\sup\{L_{k^{\prime}}\oplus\sup\{L_{k^{\prime}+1}\oplus\dots\oplus L_{k}:L_{k^{\prime}+1},\dots,L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}:L_{k^{\prime}}\in\mathcal{L},\oplus_{i=1}^{k^{\prime}}L_{i}\preceq B\}$
$=\sup\{\sup\{L_{k^{\prime}}\oplus L_{k^{\prime}+1}\oplus\dots\oplus L_{k}:L_{k^{\prime}+1},\dots,L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}:L_{k^{\prime}}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
$=\sup\{L_{k^{\prime}}\oplus\dots\oplus L_{k}:L_{k^{\prime}},\dots,L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
and the claim holds by induction. In particular,
$V_{k}(\lambda)=\sup\{L_{1}\oplus\dots\oplus L_{k}:L_{1},\dots,L_{k}\in\mathcal{L},L_{1}\oplus\dots\oplus L_{k}\preceq B\}\preceq B$
and so the natural filter is free for $k$ queries in $\mathcal{L}$ with budget $B$ by Theorem˜1. ∎
Our key result is that that this commutativity condition can essentially only be satisfied for family of queries $\mathcal{L}$ whose closure under composition $\mathcal{L}^{\infty}$ is well-ordered.
Theorem 3.
Fix a family of non-degenerate PLDs $\mathcal{L}$, i.e. every $L\in\mathcal{L}$ is supported at at least two points in $\mathbb{R}$. The natural filter Algorithm˜1 is free for every budget $B$ and every query capacity $k$ if and only if $\mathcal{L}^{\infty}$ is well-ordered with respect to $\preceq$.
We prove here that families of queries that are well-ordered under composition satisfy the commutativity condition. The proof of the other direction is significantly more involved, so we delay it until Section˜6.
Proof of Sufficient Condition for Theorem˜3.
Assume that $\mathcal{L}^{\infty}$ is well-ordered and let $L\in\mathcal{L},\mathcal{L}^{\prime}\subseteq\mathcal{L}^{<k}\subseteq\mathcal{L}^{\infty}$. For convenience, we will denote by $h_{L}$ the unique hockey-stick curve associated to a PLD $L$, namely $h_{L}(x)=\mathop{\mathbb{E}}_{Z\sim L}[(1-xe^{-Z})_{+}]$ (see Proposition˜5). Now, as in the proof of Theorem˜2, we already have $\sup\{L\oplus\mathcal{L}^{\prime}\}\preceq L\oplus\sup{\mathcal{L}^{\prime}}$ for free, so due to Proposition˜11, we just need to show that $h_{L\oplus\sup{\mathcal{L}^{\prime}}}\preceq h_{\sup\{L\oplus\mathcal{L}^{\prime}\}}$. To that end, let $\gamma>0$. By well-ordering of $\mathcal{L}^{\prime}$ and by Proposition˜11, we can find some $L^{\prime}\in\mathcal{L}^{\prime}$ so that $h_{\sup{\mathcal{L}^{\prime}}}\preceq h_{L^{\prime}}+\gamma$. In this case, we have
$h_{L\oplus\sup{\mathcal{L}^{\prime}}}(x)$ $=\mathop{\mathbb{E}}_{Z\sim L}[\mathop{\mathbb{E}}_{Z^{\prime}\sim\sup{\mathcal{L}^{\prime}}}[(1-xe^{-(Z+Z^{\prime})})_{+}]]$
$=\mathop{\mathbb{E}}_{Z\sim L}[h_{\sup{\mathcal{L}^{\prime}}}(xe^{-Z})]$
$\leq\mathop{\mathbb{E}}_{Z\sim L}[h_{L^{\prime}}(xe^{-Z})+\gamma]$
$=\mathop{\mathbb{E}}_{Z\sim L}[\mathop{\mathbb{E}}_{Z^{\prime}\sim L^{\prime}}[(1-xe^{-(Z+Z^{\prime})})_{+}]]+\gamma$
$=h_{L\oplus L^{\prime}}(x)+\gamma$
$\leq h_{\sup\{L\oplus L^{\prime}\}}(x)+\gamma$
for any $x\in\mathbb{R}^{\times}$, which yields the desired result as $\gamma$ was arbitrary. ∎
5 Notable implications
5.1 PLD and $f$-DP filters are NOT free in general
Theorems˜2 and 3 give a necessary and sufficient condition for free PLD filters. We now switch our view to $f$-DP using Propositions˜10 and 11, for which free filters correspond to Algorithm˜2 in Appendix, to illustrate these results and their implications. The condition in Theorem˜2, which is necessary and sufficient for free privacy filters, translates as:
$\forall f\in\mathcal{F},\ \text{and}\ \mathcal{F}^{\prime}\subseteq\mathcal{F}^{\infty},f\otimes\operatorname{conv}(\mathcal{F}^{\prime})=\operatorname{conv}(f\otimes\mathcal{F}^{\prime}),$ (1)
where $\otimes$ is $f$-DP composition, $f\otimes\mathcal{F}^{\prime}=\{f\otimes f^{\prime}:f^{\prime}\in\mathcal{F}^{\prime}\}$, and $\operatorname{conv}$ is the lower convex envelope of a set of tradeoff curves.
All examples also apply to PLD by Proposition˜10, but we visualize them using trade-off curves. For examples and visualizations, we work with piece-wise linear tradeoff curves, which can be composed exactly using the following result.
Proposition 12 (Exact composition of piecewise-linear tradeoff curves).
Consider two piecewise-linear tradeoff curves $f_{1},f_{2}$. Denote $\alpha_{i,n_{i}}\in[0,1]$ the lengths of intervals over which $f_{i}$ is linear, and $f^{\prime}_{i,n_{i}}\leq 0$ the slope of $f_{i}$ on those intervals. Denote $\delta_{i}\triangleq 1-f_{i}(0)$.
That is, $\sum_{j=1}^{n_{i}}\alpha_{i,j}=1$, and, noting $m_{i}(\alpha)=\arg\max_{n:\in\{1,n_{i}\}}\sum_{j=1}^{n}\alpha_{i,j}\leq\alpha$, we have:
$f_{i}(\alpha)=1-\delta_{i}+\sum_{j=1}^{m_{i}(\alpha)}f^{\prime}_{i,j}\alpha_{i,j}+\big(\alpha-\sum_{j=1}^{m_{i}(\alpha)}\alpha_{i,j}\big)f^{\prime}_{i,j+1}.$
Let $k\in[1,n_{1}n_{2}]$ index all combinations of $(j_{1,k},j_{2,k})\in[1,n_{1}]\times[1,n_{2}]$ in order of increasing values of $f^{\prime}_{1,j_{1,k}}f^{\prime}_{2,j_{2,k}}$. That is, $k$ indexes all products of a slope from $f_{1}$ and a slope for $f_{2}$ (which are now positive) sorted in decreasing order of magnitude. With ne minus sign, the first index is thus the steepest negative slope.
Then the composition $f\triangleq f_{1}\otimes f_{2}$ is also a piecewise linear tradeoff curve, with $\delta=\delta_{1}+\delta_{2}-\delta_{1}\delta_{2}$, and $\alpha_{k}=\alpha_{1,k}\alpha_{2,k}$ and slopes $f^{\prime}_{k}=-f^{\prime}_{1,j_{1,k}}f^{\prime}_{2,j_{2,k}}$. That is, denoting $m(\alpha)=\arg\max_{k:\in\{1,n_{1}n_{2}\}}\sum_{k=1}^{n}\alpha_{k}\leq\alpha$, $f=f_{1}\otimes f_{2}$ is such that:
$f(\alpha)=1-\delta_{1}-\delta_{2}+\delta_{1}\delta_{2}+\sum_{j=1}^{m(\alpha)}f^{\prime}_{k}\alpha_{k}+\big(\alpha-\sum_{j=1}^{m(\alpha)}\alpha_{k}\big)f^{\prime}_{k+1}.$
Proof.
See Section˜A.1. ∎
A first implication Theorems˜2 and 3 is that there are no free $f$-DP filters in general:
Corollary 1 ($f$-DP filters are NOT free).
Consider $\mathcal{T}$ the family of all tradeoff curves, and $\mathcal{F}=\{f_{\varepsilon,\delta},\ \varepsilon\geq 0,\ \delta\in[0,1]\}$ the set of tradeoff curves for all $(\varepsilon,\delta)$-DP mechanisms.
Neither $\mathcal{T}$ nor $\mathcal{F}$ have free $f$-DP filters. Precisely, Algorithm˜2 is not $f$-DP.
Proof.
$\mathcal{F}\subset\mathcal{T}$, and $\mathcal{F}$ has crossing tradeoff functions, which are thus not well ordered. By Theorem˜3, Algorithm˜1 is not PLD for budget $B$, and so by Propositions˜10 and 11 Algorithm˜1 is not $f$-DP. ∎
Figure˜1 directly illustrates a counter-example to Theorem˜2 (Equation˜1) visually. We show that there exists tradeoff curves in $\mathcal{F}$ (approximate DP mechanisms) such that $\alpha\ \textrm{s.t.}\ g_{1}\otimes\operatorname{conv}(g_{1},g_{2})(\alpha)<\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})(\alpha)$. That is, the exists $f=g1$ and $\mathcal{H}=\{g_{1},g_{2}\}$ such that $f\otimes\operatorname{conv}(\mathcal{H})\neq\operatorname{conv}(f\otimes\mathcal{H})$, violating (Equation˜1). By Propositions˜10 and 11, there exists $L$ and $\mathcal{L}^{\prime}$ such that $L\oplus\sup{\mathcal{L}^{\prime}}\neq\sup\{L\oplus L^{\prime}:L^{\prime}\in\mathcal{L}^{\prime}\}$, violating Theorem˜2. In fact, Figure˜1 illustrates the two cases to account for, to directly prove a special case of Theorem˜3 for $(\varepsilon,\delta)$-DP:
Proposition 13 ($(\varepsilon,\delta)$-DP curves that cross always violate Theorem˜2 (Equation˜1)).
Consider $f\in\mathcal{T}$, and any set of tradeoff curves $\mathcal{F}$ such that $f_{\varepsilon_{1},\delta_{1}}$ and $f_{\varepsilon_{2},\delta_{2}}$ cross, and $f_{\varepsilon_{1},\delta_{1}},f_{\varepsilon_{2},\delta_{2}}\in\mathcal{F}$.
$\mathcal{F}$ does not have free $f$-DP filters. Precisely, Algorithm˜2 is not $f$-DP.
Proof.
Direct corollary of Theorem˜3, self standing direct proof in Section˜A.2. ∎
(a) Case 1: $g_{1}\otimes g_{1}$ and $g_{1}\otimes g_{2}$ cross.
(b) Case 2: $g_{1}\otimes g_{1}\geq g_{1}\otimes g_{2}$
Figure 1: Counter examples to Theorem 2 (Equation˜1), showing that $f$-DP filters are NOT free in general (for the set of all tradeoff curves $\mathcal{T}$), or when choosing adaptively among the set of $(\varepsilon,\delta)$-DP mechanisms. This is true even if $f\in\mathcal{F}$.
Another interesting case is that of pure DP, or $\varepsilon$-DP, mechanisms. Indeed, this family is well ordered ($\varepsilon_{1}>\varepsilon_{2}\geq 0\Rightarrow f_{\varepsilon_{1}}\preceq f_{\varepsilon_{2}}$, illustrated on Figure˜2(a)), and known to admit privacy filters under basic composition [RRUV16]. However, this family of DP mechanisms is not closed under $f$-DP or PLD composition, and its self composition is not well ordered anymore, as illustrated on Figure˜2(b). As we can see, this is sufficient to violate Theorems˜2 and 3, and hence rule out free $f$-DP/PLD filters for the family of $\varepsilon$-DP under $f$-DP/PLD composition.
(a) $\varepsilon$-DP admits two-step filters (example)
(b) $\varepsilon$-DP general filter counter-example
Figure 2: $\varepsilon$-DP mechanisms are well ordered, but not closed under composition. If $\mathcal{F}=\{f_{\varepsilon}:\ \varepsilon\geq 0\}$, then $\mathcal{F}^{2}\subset\mathcal{F}^{\infty}$ is not well ordered (Figure˜2(a)), and by Theorem˜3 pure DP mechanisms do not admit an $f$-DP/PLD filter (Figure˜2(b)).
We have shown that $f$-DP/PLD filters are not free in general (when $\mathcal{F}$ is the set of all tradeoff curves), or for the set of $(\varepsilon,\delta)$-DP mechanisms. These results are surprising, as all previous filter results were that filters are free (RDP/zCDP [feldman2021individual, lecuyer2021practical], GDP [koskelaindividual, smith2022fullyadaptivecompositiongaussian]) or asymptotically free ($(\varepsilon,\delta)$-DP [rogers2016privacy, whitehouse2023fully]). One might hope that restricting $f$, the bound on privacy loss used as input to the $f$-DP filter, would be sufficient to ensure free filters. Two natural candidates come to mind. First, restricting $f\in\mathcal{F}$, a practical condition of choosing privacy bounds in the set of tradeoff curves used in the filter. Second, we could restrict $f$ to a GDP curve $G_{\mu}$ (the tradeoff curve of a Gaussian mechanism), or an approximate GDP curve $f_{0,\delta}\otimes G_{\mu}$ to support approximate DP $f_{\varepsilon,\delta}$ tradeoff curves in $\mathcal{F}$. These are good candidates, as $\{G_{\mu}:\mu\in\mathbb{R}^{+}\}$ is the only known set of tradeoff curves to have free $f$-DP filters [koskelaindividual, smith2022fullyadaptivecompositiongaussian].
Unfortunately, neither restriction is sufficient:
Corollary 2.
Consider the set of $(\varepsilon,\delta)$-DP tradeoff curves $\mathcal{F}=\{f_{\varepsilon,\delta}:\varepsilon\geq 0,\ \delta\in[0,1]\}$. $\mathcal{F}$ does not have free $f$-DP filters even for $f\in\mathcal{F}$.
Proof.
Consider two tradeoff curves that cross $g_{1}=f_{\varepsilon_{1},\delta_{1}}\in\mathcal{F}$ and $g_{2}=f_{\varepsilon_{2},\delta_{2}}\mathcal{F}$, that fall in case 1 in the proof of Proposition˜13 (this is possible, but the proof is identical for case 2, swapping the proper slope).
Pick $f=f_{\varepsilon,\delta}$ with $\varepsilon=\log\big(e^{2\varepsilon_{1}}-(1+e^{\varepsilon_{1}})\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}\big)$ (that is, $e^{-\varepsilon}$ is the first slope of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$) and $\delta=\delta_{1}+\delta_{2}-\delta_{1}\delta_{2}$ (that is, the $\delta$ of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$).
Consider an adversary playing Algorithm˜2 with a $g_{1}$ mechanism first, and then an adaptive choice between $g_{1},g_{2}$. By construction, $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})\geq f$, so $g_{1}\otimes g_{1}\geq f$ and $g_{1}\otimes g_{2}\geq f$: Algorithm˜2 always accepts the query.
From the proof of Theorem˜2 we know that a two-steps filter can play $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$. By construction, we have that $f(\alpha)=\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})(\alpha)$ on $\alpha\in(0,\alpha_{1}^{\star 2}]$ From Figure˜1(a) (and the proof of Proposition˜13) we know that on $\alpha\in(0,\alpha_{1}^{\star 2}]$, we have $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})(\alpha)<\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})(\alpha)=f(\alpha)$. Algorithm˜2 is thus not $f$-DP, concluding the proof. ∎
(a) GDP $G_{\mu}$ filter bound
(b) Approximate GDP $f_{0,\delta}\otimes G_{\mu}$ filter bound
Figure 3: Counter example to free $f$-DP filters when $f$ is GDP (a) or approximate GDP (b).
Corollary 3.
The set of all tradeoff curves $\mathcal{T}$ does not have free $f$-DP filters even for $f\in\{G_{\mu}:\mu\geq 0\}$.
The set of $(\varepsilon,\delta)$-DP tradeoff curves $\mathcal{F}=\{f_{\varepsilon,\delta}:\varepsilon\geq 0,\ \delta\in[0,1]\}$ does not have free $f$-DP filters even for $f\in\{f_{0,\delta}\otimes G_{\mu}:\mu\geq 0,\delta\in[0,1]\}$.
Proof.
Following the reasoning of Corollary˜2, we show that we can construct a set of two tradeoff curves $g_{1},g_{2}\in\mathcal{F}$, and a bound $f$ in the correct family, such that $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})\geq f$ but there exists values of $\alpha$ s.t. $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})(\alpha)<f$. Figure˜3(a) shows such a construction for $f\in\{G_{\mu}:\mu\geq 0\}$ and Figure˜3(b) for $f\in\{f_{0,\delta}\otimes G_{\mu}:\mu\geq 0,\delta\in[0,1]\}$. All trade-off curves are (compositions of) piecewise linear functions, or Gaussian trade-offs. They can thus be computed exactly in closed-form. ∎
5.2 Families of mechanisms with free filters
Theorem˜3, combined with Proposition˜10, show that a family of DP mechanisms that is closed under composition, and with well ordered tradeoff curves, admits free $f$-DP (and PLD by Proposition˜10) filters.
Corollary 4.
The set of all Gaussian tradeoff curves $G=\{G_{\mu}:\mu\geq 0\}$ has free filters. That is Algorithm˜2 is $f$-DP.
Proof.
For any $\mu_{1},\mu_{2}\geq 0$, $G_{\mu_{1}}\otimes G_{\mu_{2}}=G_{\sqrt{\mu_{1}^{2}+\mu_{2}^{2}}}\in G$. If $\mu_{1}>\mu_{2}$, then $G_{\mu_{1}}\preceq G_{\mu_{2}}$. Gaussian mechanisms are closed under composition and well ordered: by Theorem˜3, they have free $f$-DP/PLD filters, as already shown in [koskelaindividual, smith2022fullyadaptivecompositiongaussian]. ∎
Corollary 5.
The set of all $0,\delta$-DP tradeoff curves $\mathcal{H}=\{f_{0,\delta}:\delta\in[0,1]\}$ has free filters. That is Algorithm˜2 if $f$-DP.
Proof.
For any $\delta_{1},\delta_{2}\geq 0$, $f_{0,\delta_{1}}\otimes f_{0,\delta_{2}}=f_{0,\delta_{1}+\delta_{2}-\delta_{1}\delta_{2}}\in\mathcal{H}$. If $\delta_{1}>\delta_{2}$, then $f_{0,\delta_{1}}\preceq f_{0,\delta_{2}}$. Pure $\delta$ mechanisms are closed under composition and well ordered: by Theorem˜3, they have free $f$-DP/PLD filters. ∎
6 A Proof of Theorem˜3
In this section, we give a proof of Theorem˜3 that goes through the topology of hockey-stick curves.
As we have seen, Theorem˜3 can be proved by showing that, for any family $\mathcal{F}$ that is not well-ordered, we can find $g_{1},g_{2}\in\mathcal{F}^{\infty}$ for which
$g_{1}\otimes\operatorname{conv}\{g_{1},g_{2}\}\prec\operatorname{conv}\{g_{1}\otimes g_{1},g_{1}\otimes g_{2}\}$
Equivalently, in the language of PLDs, we would like to show that if $L_{1}$ and $L_{2}$ are not well-ordered, then $L_{1}\oplus\sup\{L_{1},L_{2}\}\succ\sup\{L_{1}\oplus L_{1},L_{1}\oplus L_{2}\}$.
Our first step is to show that gaps between PLDs can be amplified in some sense by composition. For a distribution $A$ in the reals, we denote by $\operatorname{supp}(A):=\{t\in\mathbb{R}:\forall\text{open }U\ni t,A(U)>0\}$ the support of $A$, namely the points with locally positive probability mass. It turns out that any PLD whose support can be shifted to hit gaps between other PLDs causes the gaps to align when composed.
Proposition 14 (PLD Gap Amplification).
Let $A,B,B_{1},B_{2}$ be PLDs so that $B_{i}\prec B$ with $h_{B_{i}}(e^{\varepsilon_{i}+u})<h_{B}(e^{\varepsilon_{i}+u})$ for some $\varepsilon_{1},\varepsilon_{2}\in\operatorname{supp}(A)$ and some $u\in\mathbb{R}$. Then
$A\oplus B\succ\sup\{A\oplus B_{1},A\oplus B_{2}\}.$
Proof.
First, it is clear that $A\oplus B\succeq\sup\{A\oplus B_{1},A\oplus B_{2}\}$, so we just need to find a single gap.
To that end, by continuity, we can find a constant $\rho>0$ s.t $\inf_{\varepsilon\in(\varepsilon_{i}+u\pm\rho)}h_{B}(e^{\varepsilon})-h_{B_{i}}(e^{\varepsilon})>0$. Now, set $\varepsilon^{\star}:=\varepsilon_{1}+\varepsilon_{2}+u$ and notice that
$h_{A\oplus B}(e^{\varepsilon^{\star}})-h_{A\oplus B_{i}}(e^{\varepsilon^{\star}})$ $=\mathbb{E}_{Z\sim A\oplus B}[(1-e^{\varepsilon^{\star}-Z})_{+}]-\mathbb{E}_{Z\sim A\oplus B_{i}}[(1-e^{\varepsilon^{\star}-Z})_{+}]$
$=\mathbb{E}_{Z_{1}\sim A}[\mathbb{E}_{Z_{2}\sim B}[(1-e^{(\varepsilon^{\star}-Z_{1})-Z_{2}})_{+}]-\mathbb{E}_{Z_{2}\sim B_{i}}[(1-e^{(\varepsilon^{\star}-Z_{1})-Z_{2}})_{+}]]$
$=\mathbb{E}_{Z\sim A}[h_{B}(e^{\varepsilon^{\star}-Z})-h_{B_{i}}(e^{\varepsilon^{\star}-Z})]$
$\geq\mathbb{P}_{Z\sim A}(Z\in(\varepsilon_{2-i}\pm\rho))\cdot\inf_{\varepsilon\in(\varepsilon_{2-i}\pm\rho)}h_{B}(e^{\varepsilon_{1}+\varepsilon_{2}+u-\varepsilon})-h_{B_{i}}(e^{\varepsilon_{1}+\varepsilon_{2}+u-\varepsilon})$
$=\mathbb{P}_{Z\sim A}(Z\in(\varepsilon_{2-i}\pm\rho))\cdot\inf_{\varepsilon\in(\varepsilon_{i}+u\pm\rho)}h_{B}(e^{\varepsilon})-h_{B_{i}}(e^{\varepsilon})$
$>0.$
In particular, $h_{\sup\{A\oplus B_{1},A\oplus B_{2}\}}(e^{\varepsilon^{\star}})=\max\{h_{A\oplus B_{1}}(e^{\varepsilon^{\star}}),h_{A\oplus B_{2}}(e^{\varepsilon^{\star}})\}<h_{A\oplus B}(e^{\varepsilon^{\star}})$. ∎
In order to apply gap amplification to PLDs that are not well-ordered i.e. $L_{1},L_{2}\prec\sup\{L_{1},L_{2}\}$, we would like to reduce the problem to geometric and topological properties of convex curves over $\mathbb{R}^{\times}:=(0,\infty)$. To that end, we also notice that the support of a PLD is determined by the topology of its hockey-stick curve. This result follows immediately from Proposition˜5
Proposition 15.
For a PLD $L$ with hockey-stick curve $h_{L}$, denote by $\operatorname{supp}(h_{L}):=\{x\in\mathbb{R}^{\times}:\forall\text{open }U\ni x,h|_{U}\text{ is not affine}\}$ the points at which $h_{L}$ “bends”. Then for any $e^{\varepsilon}\in\operatorname{supp}(h_{L})$, we have $\varepsilon\in\operatorname{supp}(L)$.
Our problem is now purely geometric. Given decreasing, convex, unordered curves $h_{1},h_{2}:\mathbb{R}^{\times}\to[0,1]$ that meet in the limit $x\to 0$, we would like to show that one of them can be scaled along the x-axis so that its elbows hit a region where $h_{1}<h_{2}$ and another region where $h_{1}>h_{2}$.
We proceed by analysis of these regions’ topology within the ambient space $\mathbb{R}^{\times}$. Given $S\subseteq\mathbb{R}^{\times}$, let $\partial S$ denote its boundary, i.e. those points whose neighbourhoods meet both $S$ and $S^{C}:=\mathbb{R}^{\times}\setminus S$, let $\mathrm{int}(S):=S\setminus\partial S$ denote its interior, and let $\overline{S}:=S\cup\partial S$ denote its closure. Now, let
$X(h_{1},h_{2}):=\{x\in\mathbb{R}^{\times}:h_{1}(x)<h_{2}(x)\},$
let $X(h_{2},h_{1})$ be defined similarly, and let $E(h_{1},h_{2}):=(X(h_{1},h_{2})\cup X(h_{2},h_{1}))^{C}$ be the points where $h_{1}$ and $h_{2}$ agree. Note that, convex curves over $\mathbb{R}^{\times}$ are also continuous, so $X(h_{1},h_{2})$ and $X(h_{2},h_{1})$ are both open and $E(h_{1},h_{2})$ is closed. Finally, let
$C(h_{1},h_{2}):=\{x\in\mathbb{R}^{\times}:(xr^{-1},x)\subseteq X(h_{1},h_{2}),(x,xr)\subseteq X(h_{2},h_{1})\text{ or vice-versa for some }r>1\}$
denote the simple crossing points of $h_{1}$ and $h_{2}$. We would like to show that either $\operatorname{supp}(h_{1})$ or $\operatorname{supp}(h_{2})$ “links” $X(h_{1},h_{2})$ with $X(h_{2},h_{1})$ in the following sense.
Definition 7.
For subsets $S,T_{1},T_{2}$ of $\mathbb{R}^{\times}$, we say that $S$ (multiplicatively) links $T_{1}$ and $T_{2}$ if there is $c>0$ so that $c\cdot S\cap T_{1}\neq\varnothing$ and $c\cdot S\cap T_{2}\neq\varnothing$. Furthermore, for $S,T\subseteq\mathbb{R}^{\times}$, we say that $S$ divides $T$, written $S\mid T$, if $S$ does not link $T$ and $T^{C}$.
In some cases, it will be easier to directly link a set with the boundary of another set. The following lemma shows that this will suffice for our purposes.
Lemma 2.
Let $S,U_{1},U_{2}\subseteq\mathbb{R}^{\times}$ with $U_{1}$ and $U_{2}$ open. The following are equivalent.
  1. (i)
    $S$ links $U_{1}$ and $U_{2}$
  2. (ii)
    $S$ links $U_{1}$ and $\overline{U_{2}}$
  3. (iii)
    $S$ links $\overline{U_{1}}$ and $U_{2}$
Proof.
Clearly, (i) implies (ii) and (iii). Moreover, (ii) and (iii) are symmetric, so we just need to show (ii) $\implies$ (i). To that end, assume that there is $x_{1},x_{2}\in S$ and $c>0$ such that $cx_{1}\in U_{1}$ and $cx_{2}\in\overline{U_{2}}$. Now, if $cx_{2}\in U_{2}$, we are done, so assume $cx_{2}\in\partial U_{2}$. That is, any arbitrarily small neighbourhood of $cx_{2}$ meets $U_{2}$, so, since $U_{1}\ni cx_{1}$ is open, we can just choose $c^{\prime}>0$ sufficiently close to $c$ so that $c^{\prime}x_{1}$ remains in $U_{1}$ and $c^{\prime}x_{2}\in U_{2}$. ∎
Critically, non-trivial closed and divisible subsets of $\mathbb{R}^{\times}$ have a highly regular multiplicative structure enclosed byt their divisors.
Proposition 16.
Let $S\subseteq\mathbb{R}^{\times}$ contain more than one point and let $\varnothing\subsetneq T\subsetneq\mathbb{R}^{\times}$ be closed. Then, if $S\mid T$, there is $a>1$ and $B\subseteq[1,a)$ such that $T=a^{\mathbb{Z}}B=\{a^{k}b:k\in\mathbb{Z},b\in B\}$ and such that $a\leq s^{\prime}/s$ for any $s<s^{\prime}\in S$.
Proof.
We first claim that $(s^{\prime}/s)^{\mathbb{Z}}T\subseteq T$ for any $s,s^{\prime}\in S$. Indeed, let $t\in T$ and $s,s^{\prime}\in S$. Then $t/s\cdot S\ni t/s\cdot s=t$ meets $T$, so, since $S\mid T$, we must have $T\supseteq t/s\cdot S\ni t/s\cdot s^{\prime}=(s^{\prime}/s)\cdot t$. By swapping $s$ and $s^{\prime}$, we get that $(s^{\prime}/s)^{-1}\cdot t\in T$ as well. Repeating inductively, we get that $(s^{\prime}/s)^{k}\cdot t\in T$ for arbitrary $k\in\mathbb{Z}$ as desired.
Now, let $a:=\inf\{s^{\prime}/s:s<s^{\prime}\in S\}$. Then there is a sequence $a_{n}:=s^{\prime}_{n}/s_{n}\to a$ for some $s_{n}<s^{\prime}_{n}\in S$. By the preceding claim, we have $T=\overline{T}\supseteq\overline{\bigcup_{n\in\mathbb{N}}a_{n}^{\mathbb{Z}}\ T}$. We claim that $a>1$. Indeed, if $a=1$, then $a_{n}^{\mathbb{Z}}T$ forms a cover of $\mathbb{R}^{\times}$ with maximum multiplicative distance $a_{n}\to 1$, so $\bigcup_{n=1}^{\infty}a_{n}^{\mathbb{Z}}T$ must be dense in $\mathbb{R}^{\times}$, which forces $T=\mathbb{R}^{\times}$, whereas we assumed otherwise. Therefore $a>1$ and
$T\supseteq\overline{\bigcup_{n\in\mathbb{N}}a_{n}^{\mathbb{Z}}T}=\overline{\bigcup_{k\in\mathbb{Z}}\bigcup_{t\in T}\{a_{n}^{k}t:n\in\mathbb{N}\}}\supseteq\bigcup_{k\in\mathbb{Z}}\bigcup_{t\in T}\overline{\{a_{n}^{k}t:n\in\mathbb{N}\}}\supseteq\bigcup_{k\in\mathbb{Z}}\bigcup_{t\in T}\{a^{k}t\}=a^{\mathbb{Z}}T.$
Finally, let $B:=T\cap[1,a)$. For any $t\in T$, we can choose $k\in\mathbb{Z}$ so that $a^{k}t\in[1,a)$ and, furthermore, $a^{k}t\in a^{\mathbb{Z}}T\subseteq T$, so in fact $b:=a^{k}t\in B$ and thus $t=a^{-k}a^{k}t=a^{-k}b\in a^{\mathbb{Z}}B$. That is, $T\subseteq a^{\mathbb{Z}}B\subseteq a^{\mathbb{Z}}T\subseteq T$. ∎
We also require a slight variation on the usual characterization of convexity: a concave curve that passes above and then below a convex curve must remain below it.
Lemma 3.
Let $f,g:(a,b)\to\mathbb{R}$ so that $f$ is convex and $g$ is concave. Then, if $f(x_{1})\geq g(x_{1})$ and $f(x_{2})\leq g(x_{2})$ for some $x_{1}<x_{2}\in(a,b)$, then we have $f(x)\geq g(x)$ for all $x\in(a,x_{1})$.
Proof.
Suppose we could find $x\in(a,x_{1})$ with $f(x)<g(x)$. Then, by convexity of $f$ and concavity of $g$ over $x,x_{2}]$, we have
$f(x_{1})\leq\frac{x_{2}-x_{1}}{x_{2}-x}\underbrace{f(x)}_{<g(x)}+\frac{x_{1}-x}{x_{2}-x}\underbrace{f(x_{2})}_{\leq g(x_{2})}<\frac{x_{2}-x_{1}}{x-x_{2}}g(x)+\frac{x_{1}-x}{x-x_{2}}g(x_{2})\leq g(x_{1}),$
which is a contradiction. ∎
Now, our first important observation is that $h_{i}$ is supported at the boundary of its region of dominance, excluding simple crossings. This is essentially because convex and affine curves can only cross at points.
Proposition 17.
Let $h_{1},h_{2}:\mathbb{R}^{\times}\to\mathbb{R}$ be convex. Then
  1. (i)
    $\partial X(h_{1},h_{2})\setminus C(h_{1},h_{2})\subseteq\operatorname{supp}(h_{2})$; and
  2. (ii)
    $\partial X(h_{2},h_{1})\setminus C(h_{1},h_{2})\subseteq\operatorname{supp}(h_{1})$.
Proof.
We just prove (i) as (ii) follows by symmetry. In this case it is equivalent to show that $\partial X(h_{1},h_{2})\setminus\operatorname{supp}(h_{2})\subseteq C(h_{1},h_{2})$.
To that end, consider any $x\in\partial X(h_{1},h_{2})\setminus\operatorname{supp}(h_{2})$. Then there must be some interval $(xr^{-1},xr)$ around $x$ on which $h_{2}$ is affine. Moreover, since $x\in\partial{X(h_{1},h_{2})}$, there must be some $x^{\prime}\in(xr^{-1},xr)$ for which $h_{1}(x^{\prime})<h_{2}(x^{\prime})$. On the other hand, both $X(h_{1},h_{2})$ and $X(h_{2},h_{1})$ are disjoint open subsets of a connected space, so neither meets $\partial{X(h_{1},h_{2})}$. In particular, $h_{1}(x)=h_{2}(x)$ and $x\neq x^{\prime}$. Without loss of generality, $xr^{-1}<x^{\prime}<x$.
Now, we claim that $(x^{\prime},x)\subseteq X(h_{1},h_{2})$ and $(x,xr)\subseteq X(h_{2},h_{1})$. Indeed, if there were $x^{\prime\prime}\in(x^{\prime},x)$ with $h_{1}(x^{\prime\prime})\geq h_{2}(x^{\prime\prime})$, then by applying Lemma 3 with $x_{1}=x^{\prime\prime}$ and $x_{2}=x$, we get $h_{1}(x^{\prime})\geq h_{2}(x^{\prime})$, which is a contradiction. Likewise, if we could find $x^{\prime\prime}\in(x,xr)$ with $h_{1}(x^{\prime\prime})\leq h_{2}(x^{\prime\prime})$, then applying Lemma 3 to $x_{1}=x$ and $x_{2}=x^{\prime\prime}$ yields once again $h_{1}(x^{\prime})\geq h_{2}(x^{\prime})$, which is impossible. Overall, shrinking $r$ a little as needed, we have that $x\in C(h_{1},h_{2})$. ∎
The other key observation is that the support of $h_{i}$ meets every maximal interval in which it is dominated. Recall that every open set can be decomposed into connected components (disjoint open intervals in this case). For convenience, we will write $h(0):=\lim_{x\to 0}h(x)$ when this limit exists.
Proposition 18.
Let $h_{1},h_{2}:\mathbb{R}^{\times}\to\mathbb{R}$ be convex, bounded, and decreasing such that $h_{1}(0)=h_{2}(0)$. Then for any connected component $I\neq\mathbb{R}^{\times}$ of $X(h_{1},h_{2})$, we have $I\cap\operatorname{supp}(h_{1})\neq\varnothing$. Similarly for each connected component $I\neq\mathbb{R}^{\times}$ of $X(h_{2},h_{1})$, we have $I\cap\operatorname{supp}(h_{2})\neq\varnothing$.
Proof.
The two claims are symmetric so we just prove the first one. Let $I=(a,b)\neq\mathbb{R}^{\times}$ be a connected component of $X(h_{1},h_{2})$ and assume that $I$ does not meet $\operatorname{supp}(h_{1})$, i.e. $h_{1}$ is affine over the interval $\overline{I}$.
First, we claim that $b<\infty$. Indeed, suppose that we had $b=\infty$. Then $a>0$ as $I\neq\mathbb{R}^{\times}$. Now, since $h_{1}$ and $h_{2}$ are convex and, in particular continuous over $\mathbb{R}^{\times}$, if it were the case that $h_{1}(a)<h_{2}(a)$, then we could extend the interval $(a,\infty)$ to the left to obtain a larger interval in $X(h_{1},h_{2})$, whereas we assumed that $I$ was a connected component. Thus $h_{2}(a)\leq h_{1}(a)$. But $h_{1}$ is affine and bounded on an unbounded interval $a,\infty)$, so the only possibility is that $h_{1}$ is constant on $a,\infty)$. In particular since $h_{2}$ is decreasing, $h_{2}$ must be dominated by $h_{1}$ over all of $I=(a,\infty)\subseteq X(h_{1},h_{2})$, which is a contradiction.
Next, we claim that $h_{2}$ is dominated by $h_{1}$ at the endpoints $\{a,b\}$. Indeed, if $a=0$, then $h_{1}(a)=h_{2}(a)$ by assumption and, if $a>0$, then by the same interval extension argument as before, we have $h_{2}(a)\leq h_{1}(a)$. Likewise, since $b<\infty$ and since the interval $I$ is maximal, we must have $h_{2}(b)\leq h_{1}(b)$ as well. Overall, $h_{2}$ is dominated at the endpoints of $I$ by an affine curve $h_{1}$, so by convexity $h_{2}$ must be dominated by $h_{1}$ over all of $I\subseteq X(h_{1},h_{2})$, which is impossible. ∎
We now can prove our main result of the section, which implies Theorem˜3. We note the small caveat that the queries must be non-degenerate, namely their PLDs must be supported at at least two finite points. This is not serious practical limitation.
Theorem 4.
Let $h_{1},h_{2}:\mathbb{R}^{\times}\to\mathbb{R}$ be non-degenerate hockey-stick curves that are not well-ordered. Then one of $\operatorname{supp}(h_{1}),\operatorname{supp}(h_{2})$ links $X(h_{1},h_{2})$ and $X(h_{2},h_{1})$.
Proof.
We begin with the easy case: Assume that one of $\partial X(h_{1},h_{2}),\partial X(h_{2},h_{1})$ is not fully contained in $C(h_{1},h_{2})$. Without loss of generality, suppose $\partial X(h_{1},h_{2})\nsubseteq C(h_{1},h_{2})$. By Proposition 17, $\operatorname{supp}(h_{2})$ meets $\partial X(h_{1},h_{2})\subseteq\overline{X(h_{1},h_{2})}$. Furthermore, since $h_{1}$ and $h_{2}$ are not well-ordered, $X(h_{2},h_{1})\neq\varnothing,\mathbb{R}^{\times}$ and thus $\operatorname{supp}(h_{2})$ meets $X(h_{2},h_{1})$ as well by Proposition 18. By Lemma 2, $\operatorname{supp}(h_{2})$ must link $X(h_{1},h_{2})$ and $X(h_{2},h_{1})$.
On the other hand, assume that $\partial X(h_{1},h_{2})$ and $\partial X(h_{2},h_{1})$ are both contained in $C(h_{1},h_{2})$ and assume toward a contradiction that neither $\operatorname{supp}(h_{1})$ nor $\operatorname{supp}(h_{2})$ links $X(h_{1},h_{2})$ and $X(h_{2},h_{1})$. We will derive a contradiction in two steps: First, we establish the structure of $E(h_{1},h_{2})$, $\operatorname{supp}(h_{1})$, $\operatorname{supp}(h_{2})$ and second, we derive a recurrence for the values of $h_{1}$ and $h_{2}$ at the points where they meet. By analyzing the characteristic polynomial of this recurrence, we can show that $h_{1}$ and $h_{2}$ must be either unbounded or constant, contradicting our assumptions.
Now, our first major goal is to argue that $E(h_{1},h_{2})$ can be expressed of the form $a^{\mathbb{Z}}\{b_{1},b_{2}\}$ for some $1\leq b_{1}<b_{2}<a$ and, exchanging the roles of $h_{1}$ and $h_{2}$ as needed, that $\operatorname{supp}(h_{1})=a^{\mathbb{Z}}s_{1}$ and $\operatorname{supp}(h_{2})=a^{\mathbb{Z}}s_{2}$ for some $s_{1}\in(b_{1},b_{2})$ and $s_{2}\in(b_{2},ab_{1})$.
To this end, we first argue that $E(h_{1},h_{2})$ lives in both $\overline{X(h_{1},h_{2})}$ and $\overline{X(h_{2},h_{1})}$. Indeed, suppose we could find some $x\in\partial\textrm{int}(E(h_{1},h_{2}))$. In this case,
$x$ $\in\partial E(h_{1},h_{2})$
$=\partial(X(h_{1},h_{2})\cup X(h_{2},h_{1}))^{C}$
$=\partial(X(h_{1},h_{2})\cup X(h_{2},h_{1}))$
$\subseteq\partial X(h_{1},h_{2})\cup\partial X(h_{2},h_{1})$
$\subseteq C(h_{1},h_{2}),$
so, by construction of $C(h_{1},h_{2})$, there is some punctured neighbourhood $(xr^{-1},xr)\setminus\{x\}$ of $x$ lying outside of $E(h_{1},h_{2})\supseteq\textrm{int}(E(h_{1},h_{2}))$, which would contradict $x\in\partial\textrm{int}(E(h_{1},h_{2}))$. Thus $\textrm{int}(E(h_{1},h_{2}))$ has empty boundary, so it must be both closed and open in $\mathbb{R}^{\times}$. But $h_{1}$ and $h_{2}$ cross, so we cannot have $\textrm{int}(E(h_{1},h_{2}))=\mathbb{R}^{\times}$. By connectivity of $\mathbb{R}^{\times}$, the only possibility is that $\textrm{int}(E(h_{1},h_{2}))=\varnothing$. In particular, by construction of the simple crossing points, we have
$E(h_{1},h_{2})=\partial E(h_{1},h_{2})\subseteq C(h_{1},h_{2})\subseteq\partial X(h_{1},h_{2})\cap\partial X(h_{2},h_{1})\subseteq\overline{X(h_{1},h_{2})}\cap\overline{X(h_{2},h_{1})}.$
We can now deduce the structure of $E(h_{1},h_{2})$. First, notice that $\operatorname{supp}(h_{i})$ divides $E(h_{1},h_{2})$. If not, then $\operatorname{supp}(h_{i})$ must link $E(h_{1},h_{2})$ with $\mathbb{R}^{\times}\setminus E(h_{1},h_{2})=X(h_{1},h_{2})\cup X(h_{2},h_{1})$, so it must either link $E(h_{1},h_{2})\subseteq\overline{X(h_{2},h_{1})}$ with $X(h_{1},h_{2})$ or it must link $E(h_{1},h_{2})\subseteq\overline{X(h_{1},h_{2})}$ with $X(h_{2},h_{1})$. In either case, by Lemma 2, $\operatorname{supp}(h_{i})$ links $X(h_{1},h_{2})$ with $X(h_{2},h_{1})$, whereas we assumed otherwise. Therefore, $\operatorname{supp}(h_{i})$ divides $E(h_{1},h_{2})$, so, by Proposition 16, there must be $a>1$ and $B\subseteq[1,a)$ such that $E(h_{1},h_{2})=a^{\mathbb{Z}}B$ and such that the elements of $\operatorname{supp}(h_{1})$ and of $\operatorname{supp}(h_{2})$ are multiplicatively separated by at least a factor of $a$. We claim that $B$ contains no more than two points. Indeed, suppose we could find $b_{1}<b_{2}<b_{3}\in B$. By construction of $C(h_{1},h_{2})\ni b_{1},b_{2},b_{3}$, $(b_{1},b_{2})\cup(b_{2},b_{3})\cup(b_{3},ab_{1})$ contains at least three connected components of $X(h_{1},h_{2})\cup X(h_{2},h_{1})$, so in particular, it must contain at least two connected components of $X(h_{1},h_{2})$ or of $X(h_{2},h_{1})$. By Proposition 18, $(b_{1},ab_{1})$ contains at least two points from $\operatorname{supp}(h_{1})$ or from $\operatorname{supp}(h_{2})$, which is impossible because these points were multiplicatively separated by $a$. Thus $|B|\leq 2$. On the other hand, if $B=\{b\}$, then we have $E(h_{1},h_{2})=a^{\mathbb{Z}}\{b\}=a^{2\mathbb{Z}}\{b\}\cup a^{2\mathbb{Z}+1}\{b\}=(a^{2})^{\mathbb{Z}}\{b,ab\}$, so in all cases we can write $E(h_{1},h_{2})=a^{\mathbb{Z}}\{b_{1},b_{2}\}$ for some $1\leq b_{1}<b_{2}<a$.
Finally, we determine the structure of $\operatorname{supp}(h_{1})$ and $\operatorname{supp}(h_{2})$. For $k\in\mathbb{Z}$, consider the alternating intervals $I_{k}:=(a^{k}b_{1},a^{k}b_{2})$ and $J_{k}:=(a^{k}b_{2},a^{k+1}b_{1})$. Then, for every $k\in\mathbb{Z}$, $I_{k}$ does not meet $E(h_{1},h_{2})$, so, by the intermediate value theorem, either $I_{k}\subseteq X(h_{1},h_{2})$ or $I_{k}\subseteq X(h_{2},h_{1})$ and, similarly, either $J_{k}\subseteq X(h_{1},h_{2})$ or $J_{k}\subseteq X(h_{2},h_{1})$. Now, assume without loss of generality that $I_{0}\subseteq X(h_{1},h_{2})$. By induction and by construction of the simple crossings $C(h_{1},h_{2})\supseteq E(h_{1},h_{2})$, this forces $I_{k}\subseteq X(h_{1},h_{2})$ and $J_{k}\subseteq X(h_{2},h_{1})$ for every $k\in\mathbb{Z}$. By Proposition 18, we can find $s_{1}\in(b_{1},b_{2})\cap\operatorname{supp}(h_{1})$ and $s_{2}\in(b_{2},ab_{1})\cap\operatorname{supp}(h_{2})$. In particular, $b_{1}/s_{1}\cdot\operatorname{supp}(h_{1})\ni b_{1}$ meets $E(h_{1},h_{2})=a^{\mathbb{Z}}\{b_{1},b_{2}\}\ni b_{1}$, so, since $\operatorname{supp}(h_{1})\mid E(h_{1},h_{2})$, we must have that $b_{1}/s_{1}\cdot\operatorname{supp}(h_{1})\subseteq a^{\mathbb{Z}}\{b_{1},b_{2}\}$, i.e. $\operatorname{supp}(h_{1})\subseteq a^{\mathbb{Z}}\{s_{1},b_{2}/b_{1}\cdot s_{1}\}$. We now eliminate $b_{2}/b_{1}\cdot s_{1}$ from the residue set. To that end, suppose we could find $s^{\prime}_{1}:=a^{k}b_{2}/b_{1}\cdot s_{1}\in\operatorname{supp}(h_{1})$. Then $\frac{b_{1}^{2}}{s_{1}b_{2}}\cdot s^{\prime}_{1}=a^{k}b_{1}\in E(h_{1},h_{1})$, so, since $\operatorname{supp}(h_{1})\mid E(h_{1},h_{2})$, we must have $\frac{b_{1}^{2}}{b_{2}}=\frac{b_{1}^{2}}{s_{1}b_{2}}\cdot s_{1}\in E(h_{1},h_{2})=a^{\mathbb{Z}}\{b_{1},b_{2}\}$ as well. But, $b_{1}/b_{2}\in(a^{-1},1)$, so $\frac{b_{1}^{2}}{b_{2}}=a^{k}b_{1}$ is impossible, whereas $\frac{b_{1}^{2}}{b_{2}}=a^{k}b_{2}$ forces $b_{1}/b_{2}=a^{-1/2}$. Moreover, $s_{1}\in(b_{1},b_{2})$, so we have
$s^{\prime}_{1}=a^{k}b_{2}/b_{1}\cdot s_{1}\in(a^{k}b_{2}/b_{1}\cdot b_{1},a^{k}b_{2}/b_{1}\cdot b_{2})=(a^{k}b_{2},a^{k}b_{1}\cdot(b_{2}/b_{1})^{2})=(a^{k}b_{2},a^{k+1}b_{1})=J_{k},$
which contradicts our assumption that $\operatorname{supp}(h_{1})$ fails to link $X(h_{1},h_{2})\supseteq I_{0}$ with $X(h_{2},h_{1})\supseteq J_{k}$. Therefore $\operatorname{supp}(h_{1})\subseteq a^{\mathbb{Z}}s_{1}$ which, noting that $a^{\mathbb{Z}}s_{1}\cap I_{k}=a^{k}s_{1}$, forces $\operatorname{supp}(h_{1})=a^{\mathbb{Z}}s_{1}$ due to Proposition 18. By a symmetric argument, $\operatorname{supp}(h_{2})=a^{\mathbb{Z}}s_{2}$ as well.
Now that we understand the structure of the crossings and of the support, we can calculate the trajectory of $h_{1}$ and $h_{2}$ over the crossing points. To this end, consider $y_{k}:=h_{1}(a^{k}b_{1})=h_{2}(a^{k}b_{1})$ and $y^{\prime}_{k}:=h_{1}(a^{k}b_{2})=h_{2}(a^{k}b_{2})$. Now, for any $k\in\mathbb{Z}$, the interval $a^{k}b_{1},a^{k+1}b_{2}]=\overline{I_{k}}\cup\overline{J_{k}}\cup\overline{I_{k+1}}$ meets $\operatorname{supp}(h_{2})$ at $a^{k}s_{2}$ and nowhere else. Thus $h_{2}$ is affine over $a^{k}b_{1},a^{k}s_{2}]$ and $a^{k}s_{2},a^{k+1}b_{2}]$ separately. By extrapolation in the former interval, we get
$h_{2}(a^{k}s_{2})=\frac{a^{k}b_{2}-a^{k}s_{2}}{a^{k}b_{2}-a^{k}b_{1}}h_{1}(a^{k}b_{1})+\frac{a^{k}s_{2}-a^{k}b_{1}}{a^{k}b_{2}-a^{k}b_{1}}h_{1}(a^{k}b_{2})=\frac{b_{2}-s_{2}}{b_{2}-b_{1}}y_{k}+\frac{s_{2}-b_{1}}{b_{2}-b_{1}}y^{\prime}_{k}$
and, furthermore, extrapolation over the latter interval yields
$y^{\prime}_{k+1}$ $=h_{2}(a^{k+1}b_{2})$
$=\frac{a^{k+1}b_{1}-a^{k+1}b_{2}}{a^{k+1}b_{1}-a^{k}s_{2}}h_{1}(a^{k}s_{2})+\frac{a^{k+1}b_{2}-a^{k}s_{2}}{a^{k+1}b_{1}-a^{k}s_{2}}h_{2}(a^{k+1}b_{1})$
$=\frac{a(b_{1}-b_{2})}{ab_{1}-s_{2}}\left(\frac{b_{2}-s_{2}}{b_{2}-b_{1}}y_{k}+\frac{s_{2}-b_{1}}{b_{2}-b_{1}}y^{\prime}_{k}\right)+\frac{ab_{2}-s_{2}}{ab_{1}-s_{2}}y_{k+1}$
$=\underbrace{\frac{a(s_{2}-b_{2})}{ab_{1}-s_{2}}}_{=:U^{\prime}}y_{k}+\underbrace{-\frac{a(s_{2}-b_{1})}{ab_{1}-s_{2}}}_{=:V^{\prime}}y^{\prime}_{k}+\underbrace{\frac{ab_{2}-s_{2}}{ab_{1}-s_{2}}}_{=:W^{\prime}}y_{k+1}.$
By an analogous argument, we have that
$y_{k+2}=\underbrace{\frac{a(s_{1}-b_{1})}{b_{2}-s_{1}}}_{=:U}y^{\prime}_{k}+\underbrace{-\frac{as_{1}-b_{2}}{b_{2}-s_{1}}}_{=:V}y_{k+1}+\underbrace{\frac{ab_{1}-s_{1}}{b_{2}-s_{1}}}_{=:W}y^{\prime}_{k+1}$
as well, which together yields the system
$\left\{\begin{aligned} y^{\prime}_{k+1}&=U^{\prime}y_{k}+V^{\prime}y^{\prime}_{k}+W^{\prime}y_{k+1}\\ y_{k+2}&=Uy^{\prime}_{k}+Vy_{k+1}+Wy^{\prime}_{k+1}\\ y^{\prime}_{k+2}&=U^{\prime}y_{k+1}+V^{\prime}y^{\prime}_{k+1}+W^{\prime}y_{k+2}\\ y_{k+3}&=Uy^{\prime}_{k+1}+Vy_{k+2}+Wy^{\prime}_{k+2}\end{aligned}\right$.
for $k\in\mathbb{Z}$. By isolating $y^{\prime}_{k+1}$ in terms of $y_{k},y_{k+1},y_{k+2}$ from the first two equations, substituting into the equation for $y^{\prime}_{k+2}$, and finally substituting this into the equation for $y_{k+3}$, we obtain the symmetrized recurrence
$y_{k+3}=UU^{\prime}y_{k}+(UW^{\prime}+U^{\prime}W-VV^{\prime})y_{k+1}+(V+V^{\prime}+WW^{\prime})y_{k+2}$
for $k\in\mathbb{Z}$. Thus $(y_{k})_{k\in\mathbb{Z}}$ satisfies a linear homogeneous recurrence with characteristic polynomial
$f(t)=t^{3}-(V+V^{\prime}+WW^{\prime})t^{2}-(UW^{\prime}+U^{\prime}W-VV^{\prime})t-UU^{\prime},$
which after applying some computational effort factors as
$f(t)=(t-1)(t-a)\left(t-a\left(\frac{s_{1}-b_{1}}{b_{2}-s_{1}}\right)\left(\frac{s_{2}-b_{2}}{ab_{1}-s_{2}}\right)\right).$
In particular, every root of $f$ lies in $\mathbb{R}^{\times}$. On the other hand, a homogeneous linear recurrence over the integers admits a bounded non-constant solution if and only its characteristic polynomial has a non-unit root lying on the complex unit circle (see e.g. Section 2.3 of [elaydi2005introduction]). Therefore either $y_{k}=h_{1}(a^{k}b_{1})=h_{2}(a^{k}b_{1})$ is unbounded, which is impossible because hockey-stick curves are bounded, or it must be constant, which is also impossible because hockey-stick curves are monotone, so this would force $h_{1}$ and $h_{2}$ to be constant and thus well-ordered. ∎
Theorem˜3 now follows immediately by combining Propositions 14 and 15 with Theorem 4 by taking $B:=\sup\{B_{1},B_{2}\}$.
Appendix A $f$-DP Results, Proofs and Special Cases
We present additional results focusing on $f$-DP, which we use to instantiate our main results Theorems˜1, 2 and 3 on special cases, that may help with intuition. For completeness, Algorithm˜2 states the privacy filter algorithm for $f$-DP, and is a direct analogy to Algorithm˜1.
Algorithm 2 NaturalFilter ($f$-DP)
0: Privacy bound $f$, adversary (or analyst) $\mathcal{A}$, family of tradeoff curves $\mathcal{F}$, query capacity $k$
$i=0$
while $\mathcal{A}$ continues and $i<k$ do
  $i\leftarrow i+1$
  $\mathcal{A}$ gives mechanism $\mathcal{M}_{i}$ that is $f_{i}$-DP, with $f_{i}\in\mathcal{F}$ $\triangleright$ can depend on previous results $Y_{<i}$
  if $f_{1}\otimes\ldots\otimes f_{i}\geq f$ then
   $\mathcal{A}$ receives $Y_{i}\sim\mathcal{M}_{i}$
  end if
end while
return $(Y_{1},\ldots,Y_{i})$
A.1 Proof of Proposition˜12
Proof.
For each $i\in\{1,2\}$, let $f_{i}$ be a tradeoff curve with $n_{i}$ linear segments and $f_{i}(0)=1-\delta_{i}$. Each linear segment $j\in[n_{i}]$ has a slope $f^{\prime}_{i,j}\leq 0$ over an interval $\alpha_{i,j},\alpha_{i,j+1})$. The width of interval $j$ is denoted by $w_{i,j}$. We also assume $f^{\prime}_{i,1}\geq-f^{\prime}_{i,2}\geq\cdots\geq-f^{\prime}_{i,n_{i}}$. Note that we have $\alpha_{i,1}=0$, $\alpha_{i,n_{i}+1}=1$, $f_{i}(1)=0$, and $\sum_{j=1}^{n_{i}}w_{i,j}=1$.
For each $i\in\{1,2\}$, we construct a probability distribution $Q_{i}$ supported by $[0,1]$ with density function as follows:
$q_{i}(x)=\begin{cases}-f^{\prime}_{i,1},&\text{if }x\in[\alpha_{i,1},\alpha_{i,2}),\\ -f^{\prime}_{i,2},&\text{if }x\in[\alpha_{i,2},\alpha_{i,3}),\\ \vdots&\\ -f^{\prime}_{i,j},&\text{if }x\in[\alpha_{i,j},\alpha_{i,j+1}),\\ \vdots&\\ -f^{\prime}_{i,n_{i}},&\text{if }x\in[\alpha_{i,n_{i}},\alpha_{i,n_{i}+1}),\\ (1-f_{i}(0))\cdot\delta(x-1),&\text{if }x=\alpha_{i,n_{i}+1}.\par\end{cases}$ (2)
Let $P$ be the uniform distribution over $[0,1]$. Now, we show that $T(P,Q_{i})=f_{i}$.
When $\alpha\in[\alpha_{i,1},\alpha_{i,2})$, the optimal rejection rule is
$\phi(x)=\begin{cases}\alpha/w_{i,1},&\text{if }x\in[\alpha_{i,1},\alpha_{i,2}),\\ 1,&\text{if }x=1,\\ 0,&\text{otherwise}.\end{cases}$
Note that the type I error is $\mathbb{E}_{P}[\phi]=\alpha=c\cdot w_{i,1}$. The type II error is
$\begin{array}[]{l}1-\mathbb{E}_{Q_{i}}[\phi]=1-c\cdot w_{i,1}\cdot(-f^{\prime}_{i,1})-(1-f_{i}(0))=f_{i}(0)+\alpha f^{\prime}_{i,1}=f_{i}(\alpha)=T(P,Q_{i})(\alpha).\end{array}$
When $\alpha\in[\alpha_{i,2},\alpha_{i,3})$, the optimal rejection rule is
$\phi(x)=\begin{cases}1,&\text{if }x\in[\alpha_{i,1},\alpha_{i,2})\bigcup\{1\},\\ c,&\text{if }x\in[\alpha_{i,2},\alpha_{i,3}),\\ 0,&\text{otherwise}.\end{cases}$
Note that the type I error is $\mathbb{E}_{P}[\phi]=\alpha=w_{i,1}+c\cdot w_{i,2}$. The type II error is
$\begin{array}[]{l}1-\mathbb{E}_{Q_{i}}[\phi]=f_{i}(0)-w_{i,1}\cdot(-f^{\prime}_{i,1})-c\cdot w_{i,2}\cdot(-f^{\prime}_{i,2})=f_{i}(0)+w_{i,1}\cdot f^{\prime}_{i,1}+(\alpha-w_{i,1})f^{\prime}_{i,2}=f_{i}(\alpha)=T(P,Q_{i})(\alpha).\end{array}$
More generally, when $\alpha\in[\alpha_{i,j},\alpha_{i,j+1})$, the optimal rejection rule is
$\phi(x)=\begin{cases}1,&\text{if }x\in[\alpha_{i,1},\alpha_{i,j-1})\bigcup\{1\},\\ c,&\text{if }x\in[\alpha_{i,j},\alpha_{i,j+1}),\\ 0,&\text{otherwise}.\end{cases}$
Note that the type I error is $\mathbb{E}_{P}[\phi]=\alpha=\sum_{s=1}^{j-1}w_{i,s}+c\cdot w_{i,j}$. The type II error is
$\begin{array}[]{lll}1-\mathbb{E}_{Q_{i}}[\phi]&=&f_{i}(0)-\sum_{s=1}^{j-1}w_{i,s}\cdot(-f^{\prime}_{i,s})-c\cdot w_{i,j}\cdot(-f^{\prime}_{i,j})\\ &=&f_{i}(0)+\sum_{s=1}^{j-1}w_{i,s}\cdot f^{\prime}_{i,s}+(\alpha-\sum_{s=1}^{j-1}w_{i,s})f^{\prime}_{i,j}=f_{i}(\alpha)=T(P,Q_{i})(\alpha).\end{array}$
Now, we construct a testing problem P×P $\times P\text{ vs }Q_{1}\times Q_{2}$. We first compute $(-f^{\prime}_{1,k_{1}})\cdot(-f^{\prime}_{2,k_{2}})$ for all $k_{1}\in[n_{1}]$ and $k_{2}\in[n_{2}]$. Then, we sort all these $n_{1}\cdot n_{2}$ values in a non-increasing order, using $k\in[n_{1}\cdot n_{2}]$ as the index. For a fixed $k$, we know its respective $k_{1},k_{2}$ and then $w_{1,k_{1}},w_{2,k_{2}}$ and $f^{\prime}_{1,k_{1}},f^{\prime}_{2,k_{2}}$. Let $\tilde{w}_{k}:=w_{1,k_{1}}\cdot w_{2,k_{2}}$ and $\tilde{f}^{\prime}_{k}:=f^{\prime}_{1,k_{1}}\cdot f^{\prime}_{2,k_{2}}$.
  1. 1.
    When $\alpha\in\left[0,\ \tilde{w}_{1}\right)$, we have the optimal rejection rule is
    $\phi(x)=\begin{cases}1,&\text{if }(x_{1},x_{2})\in[0,1)\times\{1\}\bigcup\{1\}\times[0,1)\bigcup\{1\}\times\{1\}\\ c,&\text{if }(x_{1},x_{2})\in[\alpha_{1,1},\alpha_{1,2})\times[\alpha_{2,1},\alpha_{2,2})\\ 0,&\text{otherwise}.\end{cases}$
    The type I error is $\mathbb{E}_{P\times P}[\phi]=\alpha=c\cdot\tilde{w}_{1}$. The type II error is
    $\mathbb{E}_{Q_{1}\times Q_{2}}[\phi]=\delta_{1}+\delta_{2}+\delta_{1}\cdot\delta_{2}+c\cdot\tilde{w}_{1}\cdot\tilde{f}^{\prime}_{1}=\delta_{1}-\delta_{2}-\delta_{1}\delta_{2}-\alpha\cdot\tilde{f}^{\prime}_{1}.$
    Therefore, we have
    $\begin{array}[]{ll}&T(P\times P,Q_{1}\times Q_{2})(\alpha)=1-\mathbb{E}_{Q_{1}\times Q_{2}}[\phi]=1-\delta_{1}-\delta_{2}-\delta_{1}\delta_{2}-\alpha\cdot\tilde{f}^{\prime}_{1}.\end{array}$
  2. 2.
    When $\alpha\in\left[\tilde{w}_{1},\ \tilde{w}_{1}+\tilde{w}_{2}\right)$, we have
    $\begin{array}[]{l}\alpha=\tilde{w}_{1}+c\cdot\tilde{w}_{2}\quad\Rightarrow c\cdot\tilde{w}_{2}=\alpha-\tilde{w}_{1},\\ T(P\times P,Q_{1}\times Q_{2})(\alpha)=1-\delta_{1}-\delta_{2}-\delta_{1}\delta_{2}-\tilde{w}_{1}\cdot\tilde{f}^{\prime}_{1}-c\cdot\tilde{w}_{2}\cdot\tilde{f}^{\prime}_{2}\\ =1-\delta_{1}-\delta_{2}-\delta_{1}\delta_{2}-\tilde{w}_{1}\cdot\tilde{f}^{\prime}_{1}-(\alpha-\tilde{w}_{1})\cdot\tilde{f}^{\prime}_{2}.\end{array}$
  3. 3.
    When $\alpha\in\left[\tilde{w}_{1}+\tilde{w}_{2},\ \tilde{w}_{1}+\tilde{w}_{2}+\tilde{w}_{3}\right)$, we have
    $\begin{array}[]{l}\alpha=\sum_{s=1}^{2}\tilde{w}_{s}+c\cdot\tilde{w}_{3}\quad\Rightarrow c\cdot\tilde{w}_{3}=\alpha-\sum_{s=1}^{2}\tilde{w}_{s},\\ T(P\times P,Q_{1}\times Q_{2})(\alpha)=1-\delta_{1}-\delta_{2}-\delta_{1}\delta_{2}-\sum_{s=1}^{2}\tilde{w}_{s}\cdot\tilde{f}^{\prime}_{s}-c\cdot\tilde{w}_{3}\cdot\tilde{f}^{\prime}_{3}\\ =1-\delta_{1}-\delta_{2}-\delta_{1}\delta_{2}-\sum_{s=1}^{2}\tilde{w}_{s}\cdot\tilde{f}^{\prime}_{s}-(\alpha-\sum_{s=1}^{2}\tilde{w}_{s})\cdot\tilde{f}^{\prime}_{3}.\end{array}$
  4. 4.
    When $\alpha\in\left[\sum_{s=1}^{j}\tilde{w}_{s},\ \sum_{s=1}^{j+1}\tilde{w}_{s}\right)$, we have
    $\begin{array}[]{l}\alpha=\sum_{s=1}^{j}\tilde{w}_{s}+c\cdot\tilde{w}_{j+1}\quad\Rightarrow c\cdot\tilde{w}_{j+1}=\alpha-\sum_{s=1}^{j}\tilde{w}_{s},\\ T(P\times P,Q_{1}\times Q_{2})(\alpha)=1-\delta_{1}-\delta_{2}-\delta_{1}\delta_{2}-\sum_{s=1}^{j}\tilde{w}_{s}\cdot\tilde{f}^{\prime}_{s}-c\cdot\tilde{w}_{j+1}\cdot\tilde{f}^{\prime}_{j+1}\\ =1-\delta_{1}-\delta_{2}-\delta_{1}\delta_{2}-\sum_{s=1}^{j}\tilde{w}_{s}\cdot\tilde{f}^{\prime}_{s}-(\alpha-\sum_{s=1}^{j}\tilde{w}_{s})\cdot\tilde{f}^{\prime}_{j+1}.\end{array}$
A.2 A Simple Derivation of Theorem˜3 for Approximate DP
Definition 8 (Crossing tradeoff curves).
We say that two tradeoff curves $f,g\in\mathcal{T}$ cross if there exists $\alpha_{1},\alpha_{2}\in[0,1]$ s.t.:
$f(\alpha_{1})>g(\alpha_{1})\ \textrm{and}\ f(\alpha_{2})<g(\alpha_{2}).$
By continuity, there has to be $\alpha\in(\alpha_{1},\alpha_{2})$ s.t. $f(\alpha)=g(\alpha)$. For any two $(\varepsilon,\delta)$-DP with tradeoff curves that cross, there cannot be free $f$-DP filters.
We can now prove Proposition˜13, a special case of Theorem˜3 for $(\varepsilon,\delta)$-DP.
Proof.
Consider $g_{1}\triangleq f_{\varepsilon_{1},\delta_{1}}$ and $g_{2}\triangleq f_{\varepsilon_{2},\delta_{2}}$ that cross. By symmetry w.r.t. the $y=x$ axis, they have to cross at an $\alpha<\min\{\alpha_{1}^{\star},\alpha_{2}^{\star}\}$, where $\alpha_{i}^{\star}\ \textrm{s.t.}\ \alpha_{i}^{\star}=g_{i}(\alpha_{i}^{\star})$ (otherwise they would never cross). This means that their linear segments cross: one function starts strictly lower, and then ends-up strictly higher by $\min\{\alpha_{1}^{\star},\alpha_{2}^{\star}\}$ (again, otherwise they would never cross). Without loss of generality, we reorder functions so that $g_{2}$ starts strictly lower.
We now know that $\delta_{2}>\delta_{1}\geq 0$ ($g_{2}$ starts lower); $\varepsilon_{1}>\varepsilon_{2}\geq 0$ ($g_{1}$ has a steeper slope to catch-up to $g_{2}$ so that they can cross); and $\alpha_{2}^{\star}>\alpha_{1}^{\star}$. The last inequality holds because $g_{1}$ is strictly lower by the time is crosses $y=x$ at $\alpha_{1}^{\star}$, so at that point $g_{2}(\alpha_{1}^{\star})>\alpha_{1}^{\star}$, and by continuity of $g_{2}$ we have that $\alpha_{2}^{\star}>\alpha_{1}^{\star}$. Since we are dealing with approximate DP curves, we also know that $\alpha_{i}^{\star}=\frac{1-\delta_{i}}{1+e^{\varepsilon_{i}}}$.
Now, in Theorem˜2, pick $\mathcal{H}=\{g_{1},g_{2}\}\subseteq\mathcal{F}$ and $g=g_{1}\in\mathcal{H}\subseteq\mathcal{F}$. We make a few useful observations about the first segment of $g_{1}\otimes g_{1}$, $g_{1}\otimes g_{2}$, and $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$, by applying Proposition˜12:
The first segment of $g_{1}\otimes g_{1}$ starts at $(1-\delta_{1})^{2}$, with slope $e^{2\varepsilon_{1}}$ for $\alpha\in[0,\alpha_{1}^{\star 2}]$. The second slope is $e^{\varepsilon_{1}}e^{-\varepsilon_{1}}=-1$.
The first segment of $g_{1}\otimes g_{2}$ starts at $(1-\delta_{1})(1-\delta_{2})$, with slope $e^{\varepsilon_{1}+\varepsilon_{2}}$ for $\alpha\in[0,\alpha_{1}^{\star}\alpha_{2}^{\star}]$. The second slope is $e^{\varepsilon_{1}-\varepsilon_{2}}<-1$.
The first segment of $\operatorname{conv}(g_{1},g_{2})$ starts at $(1-\delta_{2})$ (the lowest point at $\alpha=0$), and since $g_{1}$ and $g_{2}$ cross, it meets $g_{1}$ at $\alpha=\alpha_{1}^{\star}$, where its value is $g_{1}(\alpha_{1}^{\star})=1-\delta_{1}-e^{\varepsilon_{1}}\alpha_{1}^{\star}$. As a result, the first slope of $\operatorname{conv}(g_{1},g_{2})$ is $\frac{1-\delta_{1}-e^{\varepsilon_{1}}\alpha_{1}^{\star}-(1-\delta_{2})}{\alpha_{1}^{\star}}=\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}-e^{\varepsilon_{1}}$.
The first segment of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$ starts at $(1-\delta_{1})(1-\delta_{2})$, with first slope $\frac{\delta_{1}-\delta_{2}}{\alpha_{1}^{\star}}e^{\varepsilon_{1}}-e^{2\varepsilon_{1}}$ for $\alpha\in[0,\alpha_{1}^{\star 2}]$.
Let us study the first segment of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$ and compare it to that of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$. We first reason about the possible shapes of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$. We saw that $g_{1}\otimes g_{1}(0)=(1-\delta_{1})^{2}>(1-\delta_{1})(1-\delta_{2})=g_{1}\otimes g_{2}(0)$, so $g_{1}\otimes g_{2}$ starts lower. Since $e^{2\varepsilon_{1}}<-e^{\varepsilon_{1}+\varepsilon_{2}}<-e^{\varepsilon_{1}-\varepsilon_{2}}$, $g_{1}\otimes g_{1}<-1$, we know that $g_{1}\otimes g_{1}$ starts with a steepest slope, which stops first (at $\alpha_{1}^{\star 2}<\alpha_{1}^{\star}\alpha_{2}^{\star}$) and then gets the smallest second slope (until at least $\alpha\geq\alpha_{g_{1}\otimes g_{2}}^{\star}$). As a result, either that first slope is steep enough such that $g_{1}\otimes g_{1}(\alpha_{1}^{\star 2})<g_{1}\otimes g_{2}(\alpha_{1}^{\star 2})$ and the composed curves cross, or they will never cross and $g_{1}\otimes g_{2}\leq g_{1}\otimes g_{1}$. We treat each case in turn.
Case 1: $g_{1}\otimes g_{1}(\alpha_{1}^{\star 2})<g_{1}\otimes g_{2}(\alpha_{1}^{\star 2})$ and the composed curves cross (Figure˜1(a)). In that case, we know that $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})(\alpha_{1}^{\star 2})=g_{1}\otimes g_{1}(\alpha_{1}^{\star 2})=(1-\delta_{1})^{2}-e^{2\varepsilon_{1}}\alpha_{1}^{\star 2}$. As a result, the slope of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$ in $\alpha\in[0,\alpha_{1}^{\star 2}]$ is $\frac{(1-\delta_{1})^{2}-e^{2\varepsilon_{1}}\alpha_{1}^{\star 2}-(1-\delta_{1})(1-\delta_{2})}{\alpha_{1}^{\star 2}}=\frac{(1-\delta_{1})(\delta_{2}-\delta_{1})}{\alpha_{1}^{\star 2}}-e^{2\varepsilon_{1}}=(1+e^{\varepsilon_{1}})\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}-e^{2\varepsilon_{1}}$ (the last step uses $\alpha_{i}^{\star}=\frac{1-\delta_{i}}{1+e^{\varepsilon_{i}}}$).
Remember that $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})(0)=(1-\delta_{1})(1-\delta_{2})=\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$, and the first slope of with first slope of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})=\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}e^{\varepsilon_{1}}-e^{2\varepsilon_{1}}<(1+e^{\varepsilon_{1}})\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}-e^{2\varepsilon_{1}}$ since $\delta_{2}-\delta_{1}>0$.
We have shown that in this case, $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})>\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$ at least on $(0,\alpha_{1}^{\star 2}]$, so by Theorem˜2, and there are no $f$-DP filters.
Case 2: $g_{1}\otimes g_{1}(\alpha_{1}^{\star 2})\geq g_{1}\otimes g_{2}(\alpha_{1}^{\star 2})$ and $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})=g_{1}\otimes g_{2}$ (Figure˜1(b)). Intuitively, we know that the first slope of $\operatorname{conv}(g_{1},g_{2})$ is steeper than that of $g_{2}$, since $g_{2}$ starts lower and crosses with $g_{1}$. Formally, the first slope of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$ is $\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}e^{\varepsilon_{1}}-e^{2\varepsilon_{1}}=-e^{\varepsilon_{1}}\big(e^{\varepsilon_{1}}-\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}\big)$. We know that $1-\delta_{1}-e^{\varepsilon_{1}}\alpha_{1}^{\star}=g_{1}(\alpha_{1}^{\star})<g_{2}(\alpha_{1}^{\star 2})=1-\delta_{2}-e^{\varepsilon_{2}}\alpha_{1}^{\star}$, so $\delta_{2}-\delta_{1}<\alpha_{1}^{\star}(e^{\varepsilon_{1}}-e^{\varepsilon_{2}})$ and $0<e^{\varepsilon_{2}}<e^{\varepsilon_{1}}-\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}$. That is, $e^{\varepsilon_{1}}\big(e^{\varepsilon_{1}}-\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}\big)<-e^{\varepsilon_{1}}e^{\varepsilon_{2}}$.
In this case as well, the first slope of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$ is strictly more steep than that of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$, so $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})>\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$ at least on $(0,\alpha_{1}^{\star 2}]$. By Theorem˜2, and there are no free $f$-DP filters. ∎

### v2

@@
 Abstract
-We study natural privacy filters, which enable the exact composition of differentially private (DP) mechanisms with adaptively chosen privacy characteristics.
-Earlier privacy filters consider only simple privacy parameters such as Rényi-DP or Gaussian DP parameters.
-Natural filters account for the entire privacy profile of every query, promising greater utility for a given privacy budget.
+We study privacy filters, which enable privacy accounting for differentially private (DP) mechanisms with adaptively chosen privacy characteristics.
+We develop a general theory that characterizes the worst-case privacy loss of an interaction involving an analyst that respects some restrictions on what queries they may issue.
+We apply this theory to develop residue filters, which unifies existing privacy filters.
+We develop the Gaussian DP (GDP) residue filter, which strictly improves upon the naïve GDP filter.
+We also show that residue filters capture the natural filter, which promises greater utility by leveraging exact privacy accounting techniques.
+Earlier privacy filters consider only simple privacy parameters such as Rényi-DP or GDP parameters.
+Natural filters account for the entire privacy profile of every query, promising more efficient use of a given privacy budget.
 We show that, contrary to other forms of DP, natural privacy filters are not free in general.
-Indeed, we show that only families of privacy mechanisms that are well-ordered when composed admit free natural privacy filters.
+We present a characterization of when a family of private queries admits free natural filters for a given budget.
+In particular, only families of privacy mechanisms that are totally-ordered when composed admit free natural privacy filters with respect to an arbitrary privacy budget.
+Finally, we show that, while the natural approximate-DP filter can fail in the presence of adaptive adversary, it cannot fail too badly: the output remains approximate-DP with parameters at most poly-logarithmically worse than the intended privacy parameters.
 1 Introduction
-Differential Privacy (DP) [dwork2006calibrating] is a rigorous notion of data privacy, used to protect against a host of different privacy attacks, including membership inference [homer2008resolving, dwork2015robust, carlini2022membership] and data reconstruction [dinur2003revealing, dwork2007price, carlini2021extracting, balle2022reconstructing, dick2023confidence, nasr2023scalable].
-As such, there is a sustained push to integrate and deploy DP in a large number of applications [opendp-registry], from machine learning (ML) and artificial intelligence (AI) models [kairouz2021practical, sinha2025vaultgemma], to analytics collection [rogers2020linkedin, apple-emoji, ding2017collecting], to data releases [adeleye2023publishing, abowd20222020], as well as in general libraries and systems supporting those applications [gaboardi2020programming, wilson2020differentially, jax-privacy2022github, tholoniat2024cookie, ghazi2025differential].
+Differential Privacy (DP) Dwork et al.
+(2006b) is a rigorous notion of data privacy that protects against a host of attacks including membership inference Homer et al.
+(2008); Dwork et al.
+(2015); Carlini et al.
+(2022) and data reconstruction Dinur and Nissim (2003); Dwork et al.
+(2007); Carlini et al.
+(2021); Balle et al.
+(2022); Dick et al.
+(2023); Nasr et al.
+(2023).
+As such, there is a sustained push to integrate and deploy DP in a large number of applications OpenDP (2025), from machine learning (ML) and artificial intelligence (AI) models Kairouz et al.
+(2021); Sinha et al.
+(2025), to analytics collection Rogers et al.
+(2020); Apple (2017); Ding et al.
+(2017), to data releases Adeleye et al.
+(2023); Abowd et al.
+(2022), as well as in general libraries and systems supporting those applications Gaboardi et al.
+(2020); Wilson et al.
+(2020); Balle et al.
+(2025); Tholoniat et al.
+(2024); Ghazi et al.
+(2025).
 This practical DP push also motivates important theoretical progress.
-A prominent direction of theoretical improvement comes from alternative definitions of DP, that enable tighter privacy loss accounting under composition and improve the privacy-accuracy tradeoff.
-The most successful definitions include $(\varepsilon,\delta)$-DP [dwork2006our] (known as approximate DP), the Rényi divergence family of DP definitions such as zCDP [dwork2016concentrated, bun2016concentrated] and RDP [mironov2017renyi], and privacy definitions that generalize approximate DP by considering all $(\varepsilon,\delta)$-DP values compatible with a given mechanism, such as the privacy loss distribution (PLD) [sommer2019privacy] and $f$-DP [dong2022gaussian].
-The latter definitions, PLD and $f$-DP, are theoretically appealing as they tightly capture resilience to membership inference attacks framed as hypothesis tests [3, dong2022gaussian].
-However, composition under these definitions is more complex and can be computationally difficult, with analyses of specific algorithms that rely on numerical approaches to handle composition in practice [koskela2020computing, gopi2021numerical].
-Two other theoretical improvements enable new forms of composition that are necessary requirements for practical, real world DP usage.
-First, concurrent composition [vadhan2021concurrent] supports composition of DP mechanisms running over multiple rounds of interaction with the user—such as the sparse vector technique [dwork2009complexity]—that are running in parallel.
-This is required in practical systems which typically have to handle queries from multiple parties simultaneously, and where it is not possible to enforce a sequential ordering of the queries.
-Even individual users may want to run DP computations concurrently with long-running, multi-step mechanism (e.g., [kostopoulou2023turbo]).
-Second, adaptive composition, the main topic of this paper, enables the composition of DP mechanisms with adaptively chosen privacy parameters (as opposed to regular composition, in which these parameters have to be known in advance).
-Adaptive composition is formalized through the notion of privacy filters [RRUV16], which accept DP mechanisms as long as their composition remains within a bound fixed in advance (more details in §2).
-Such adaptive composition is necessary to support long running applications, in which analysts can adaptively query a system with mechanisms and DP parameters of their choice [lecuyer2019privacy, kostopoulou2023turbo, kuchler2024cohere].
-Furthermore, filters enable crucial privacy accounting savings through advanced versions of parallel composition [lecuyer2019privacy], as well as individual (personalized) DP [ebadi2015differential], which enables data dependent privacy accounting in ML model training [feldman2021individual, yu2022individual] and measurement systems [tholoniat2024cookie, ghazi2025differential].
-A key question in DP composition is whether these various improvements combine with each other.
-Ideally these more flexible forms of composition come for free: that is, that can we apply composition results under concurrent mechanisms or adaptive budgets using privacy filters, without a penalty in the final privacy loss after composition.
-Flexible composition for free is important not only because it is efficient privacy-wise, but also because it enables reusing all existing composition results, including tight analyses for specific mechanisms and empirical computational results.
-As it turns out, concurrent composition is indeed free for all DP definitions [vadhan2023concurrent, lyu2022composition].
-Moreover, concurrent and adaptive filter composition reduces to adaptive filters [haney2023concurrent], and thus has the same composition cost.
-Therefore, the full expressivity of both concurrency and adaptivity hinges on free privacy filter composition for different notions of DP.
-Filters, however, are not as well understood.
-We know that adaptive composition under privacy filters is free for $\varepsilon$-DP [RRUV16] and Rényi-based DP definitions [FZ21, lecuyer2021practical] (zCDP, RDP).
-Filters are also free for Gaussian mechanisms under $f$-DP and PLD accounting [ST22, KTH22].
-However, for $(\varepsilon,\delta)$-DP, the best known result for adaptive composition with filters is only asymptotically free, with worse constants than non-adaptive composition [WRRW23].
-Even less is known about filters under $f$-DP and PLD.
-In particular, we consider natural filters, which leverage lossless composition of the privacy characteristics of individual queries.
-Natural filters can be realized via PLDs, privacy profiles, or $f$-DP accounting.
-In this paper, we make three important contributions towards fully understanding natural privacy filters:
+A prominent direction of theoretical improvement comes from alternative quantifications of privacy, especially those that offer tighter or even lossless accounting of composition.
+The classical $(\varepsilon,\delta)$-DP Dwork et al.
+(2006a) (known as approximate DP) was first refined by forms of privacy derived from Rényi divergence including zCDP Dwork and Rothblum (2016); Bun and Steinke (2016) and RDP Mironov (2017).
+Later, exact forms of privacy accounting were developed including privacy profiles, privacy loss distributions (PLDs) Sommer et al.
+(2019), and $f$-DP Dong et al.
+(2019).
+These exact forms of accounting are practically and theoretically appealing as they tightly capture composition as well as resilience to membership inference attacks framed as hypothesis tests Wasserman and Zhou (2010); Dong et al.
+(2019).
+Exact composition, while more complex, can be computed efficiently to arbitrary precision by numerical methods Koskela et al.
+(2020); Gopi et al.
+(2021).
+Theoretical improvements enabling new forms of composition are also necessary for practical DP usage.
+Concurrent composition Vadhan and Wang (2021) supports composition of DP mechanisms running over multiple rounds of parallel interactions with the user—such as the sparse vector technique Dwork et al.
+(2009).
+This is required in systems that handle queries from multiple parties simultaneously where it is not possible to enforce a sequential ordering of the queries.
+Individual users may also run DP computations concurrently with a long-running, multi-step mechanism (see e.g.
+Kostopoulou et al.
+(2023)).
+A natural question is whether developments in privacy accounting are compatible with new forms of composition.
+Ideally, more flexible forms of composition come for free, i.e.
+existing results for standard composition can be applied to flexibly composed mechanisms without additional privacy costs.
+Free flexible composition also enables reuse of existing results including tight analyses for specific mechanisms and empirical computational results.
+As it turns out, concurrent composition is indeed free for all common forms of privacy Vadhan and Zhang (2023); Lyu (2022).
+Specifically, Vadhan and Zhang Vadhan and Zhang (2023) show that for both $f$-DP (including PLD and approximate DP) and RDP (including zCDP) notions of privacy, concurrent composition can be seen as a post-processing of standard composition and thus has the same cost.
+Another form of flexible composition and the subject of this work is fully adaptive composition, which concerns the composition of DP mechanisms with interactively chosen privacy costs.
+Standard composition on the other hand requires the privacy costs to be fixed in advance.
+This is insufficient for long-running applications in which analysts interactively re-assess the privacy budget allocated to queries in order to meet operational needs Lécuyer et al.
+(2019); Kostopoulou et al.
+(2023); Küchler et al.
+(2024).
+Fully adaptive analysis can also lead to substantial privacy loss savings through advanced versions of parallel composition Lécuyer et al.
+(2019) and individual (personalized) DP Ebadi et al.
+(2015), which enables data dependent privacy accounting in ML model training Feldman and Zrnic (2021); Yu et al.
+(2022) and measurement systems Tholoniat et al.
+(2024); Ghazi et al.
+(2025).
+Fully adaptive composition is formalized in terms of privacy filters Rogers et al.
+(2016).
+A privacy filter is an algorithm that monitors an interaction between an analyst and a data curator and permits queries as long as a privacy budget condition is maintained.
+The stopping condition should be as permissive as possible such that resulting interaction provably satisfies a meaningful privacy guarantee.
+As with concurrent composition, a key research question is whether advanced privacy accounting techniques can be used to construct free privacy filters for fully adaptive composition.
+Haney et al.
+Haney et al.
+(2023) show that concurrent, adaptive composition reduces to adaptive composition.
+Therefore, the full expressivity of both concurrency and adaptivity hinges on the existence of free privacy filters for different notions of DP.
+However, privacy filters are not as well understood as concurrent composition.
+We know that adaptive composition under privacy filters is free for $\varepsilon$-DP Rogers et al.
+(2016) and Rényi-based DP Feldman and Zrnic (2021); Lécuyer (2021) (zCDP, RDP).
+The GDP filter, which applies exact ($f$-DP, PLD) accounting to the Gaussian mechanism is also free Smith and Thakurta (2022); Koskela et al.
+(2022).
+This filter was also recently extended to support approximate GDP budgets Tran et al.
+(2026).
+However, for $(\varepsilon,\delta)$-DP filters, the best known filter by Whitehouse et al.
+Whitehouse et al.
+(2023) extends the RDP filter from Feldman and Zrnic (2021) to approximate RDP and zCDP, enabling a filter based on lossy translation from $(\varepsilon,\delta)$-DP to approximate zCDP Bun and Steinke (2016).
+This yields an $(\varepsilon,\delta)$-DP privacy filter that is asymptotically free but with worse constants than non-adaptive composition due to translation loss.
+Even less is known about filters under exact forms of privacy.
+We study the natural filter in which the realized privacy costs of each query are composed by exact accounting and the realized total cost is compared to a fixed budget.
+We make four important contributions towards fully understanding privacy filters:
   • First, in §3 we present a general theory of adaptive mechanisms that issue adversarial queries with a constrained per-query privacy cost.
-  • In §4 we apply this theory to the natural filter and, surprisingly, we show that adaptive composition is NOT free under the PLD nor $f$-DP notions of privacy (see Corollary˜1).
-  • Finally, we show in §6 that filters are only free for very restricted families of mechanisms: those that are closed under composition, and admit totally ordered tradeoff functions.
+  • In §4, we show that existing privacy filters are subsumed by residue filters, which perform exact accounting under a residual update condition.
+We develop a new GDP residue filter that offers significant privacy savings compared to the standard GDP filter for some settings.
+  • In §5 we apply our theory of constrained adversaries to calculate the exact privacy cost of an interaction with the natural filter.
+We show that the natural filter is universally free for any budget only for queries that are closed under composition and totally-ordered, in which case the natural filter is captured by a residue filter.
+  • In §7, we consider the widely used natural $(\varepsilon,\delta)$-DP filter, i.e.
+the filter based on exact composition and $(\varepsilon,\delta)$-DP budget.
+We show that, while this filter can fail, it does not fail too badly: the interaction still satisfies $(\chi\cdot\varepsilon,\chi\cdot\delta)$-DP where $\chi$ is only poly-logarithmic in $1/\varepsilon,1/\delta$ and the number of rounds of adaptivity.
 Our work also includes some minor contributions including that PLDs can be assigned a natural ordering satisfying a completeness property enabling one to take a supremum of PLDs (see Proposition˜11).
 As far as we are aware this property has not been documented in the privacy accounting literature.
-We alo document in §5 some families of queries and budgets that do and do not admit free filters.
+We also record in §6 a number of families of queries and budgets that do and do not admit free natural filters.
+Our key result showing the necessity of total-ordering for universally free natural filters is given in §8 and involves a privacy gap amplification technique that may be of general interest.
+We note that a recent preprint independent and simultaneous to our work also studies natural filters Tran et al.
+(2026).
+That work states that a total-ordering condition is necessary and sufficient for free natural filters when the queries have unbounded support.
+Our work contains an analogous result for general queries.
+In particular, our result also applies to discretized finite-support queries that arise in practical privacy accounting.
+Our work also offers a fine-grained analysis of the cost of interacting with the natural filter for particular budgets.
 2 Preliminaries
-Differential privacy is the study of randomized algorithms (called mechanisms) $M:\mathcal{D}\to\mathcal{Y}$ where $\mathcal{D}$ denotes some space of datasets containing sensitive personal information.
+Differential privacy is the study of randomized algorithms (called mechanisms) $M:\mathcal{D}\to\mathcal{Y}$ where $\mathcal{D}$ denotes some space of datasets containing sensitive information.
 Differential privacy requires that an adversary should not gain much statistical advantage for inferring whether the mechanism was run on dataset $D_{1}\in\mathcal{D}$ or a second dataset $D_{2}\in\mathcal{D}$ when the datasets are close to each other in some sense.
 That is, the distributions of $M(D_{1})$ and $M(D_{2})$ should be difficult to distinguish from one another in the hypothesis testing sense and this should hold uniformly for any close datasets.
-The measure of closeness for datasets is generally defined by a neighbouring relation but this turns out not to be relevant to our work.
+The measure of closeness for datasets is determined by a neighbouring relation but this turns out not to be relevant to our work.
 In fact, our conclusions hold for an arbitrary pair of datasets.
@@
 $M(D_{2})$.
+In general, the privacy characteristics of a mechanism may not be determined by a single pair.
+For the purposes of constructing filters, this can be handled by considering multiple dataset pairs that characterize some regime of the worst-case privacy characteristics of a mechanism.
+One can also smooth over the choices of datasets by considering dominating pairs, though this is lossy in general.
+See Zhu et al.
+(2022) for more discussion.
 In our work we reason about the composition of adaptive mechanisms, which are permitted to inspect outcomes of any previously run mechanism.
@@
 $(M_{1}\otimes M_{2})(D):=(Y_{1},Y_{2}),Y_{1}\sim M_{1}(D),Y_{2}\sim M_{2}(D;Y_{1}).$
-It is natural in general to view the privacy characteristics of any private mechanism as a hypothesis testing problem [3, DRS22].
-As such it is desirable to furnish hypothesis testing problem $(P,Q)$ with some kind of ordering so that we can compare the privacy guarantees offered by various mechanisms.
-We also require an ordering so that we can make sense of privacy filters.
+It is natural in general to view the privacy characteristics of any private mechanism as a hypothesis testing problem Wasserman and Zhou (2010); Dong et al.
+(2019).
+As such it is desirable to furnish hypothesis testing problems $(P,Q)$ with some kind of ordering so that we can compare the privacy guarantees offered by various mechanisms.
+We also require an ordering so that we can make sense of natural privacy filters.
 Fortunately testing pairs can be equipped with a very natural order, namely the Blackwell order.
 Definition 2.
-For a pair of distributions $(P,Q)$ defined on a common probability space $\Omega$ and a second pair $(P^{\prime},Q^{\prime})$ defined on $\Omega^{\prime}$, we say that $(P^{\prime},Q^{\prime})$ dominates $(P,Q)$ in the Blackwell order, written $(P,Q)\preceq(P^{\prime},Q^{\prime})$, if we can find a Markov kernel $\phi$ from $\Omega^{\prime}$ to $\Omega$ such that $P=\phi P^{\prime}$ and $Q=\phi Q^{\prime}$.
+For a pair of distributions $(P,Q)$ defined on a common probability space $\Omega$ and a second pair $(P^{\prime},Q^{\prime})$ defined on $\Omega^{\prime}$, we say that $(P^{\prime},Q^{\prime})$ dominates $(P,Q)$ in the Blackwell order, written $(P,Q)\preceq(P^{\prime},Q^{\prime})$, if we can find a Markov kernel $\kappa$ from $\Omega^{\prime}$ to $\Omega$ such that $P=\kappa P^{\prime}$ and $Q=\kappa Q^{\prime}$.
 We write $(P,Q)\prec(P^{\prime},Q^{\prime})$ if $(P,Q)\preceq(P^{\prime},Q^{\prime})$ but not vice-versa.
+If $(P,Q)\preceq(Q,P)\preceq(P,Q)$ the pair is called symmetric.
 Informally, the Blackwell order says that more information is available for distinguishing $P^{\prime}$ and $Q^{\prime}$ compared to $P$ vs $Q$.
-For mechanisms, we can interpret this as $(M(D_{1}),M(D_{2}))\preceq(M^{\prime}(D_{1}),M^{\prime}(D_{2}))$ when $M$ can be reconstructed from $M^{\prime}$ by introducing additional randomness through the kernel $\phi$.
+For mechanisms, we can interpret this as $(M(D_{1}),M(D_{2}))\preceq(M^{\prime}(D_{1}),M^{\prime}(D_{2}))$ when $M$ can be reconstructed from $M^{\prime}$ by introducing additional randomness through the kernel $\kappa$.
 Note that the Blackwell order is a partial order.
-Instead of a carrying around a pair of distributions to represent the privacy characteristics of a mechanism, we can consolidate them into the privacy loss distribution (PLD) [2].
-As we will see the PLD possesses very desirable composition properties and is a practical basis for privacy loss computations.
+Instead of a carrying around a pair of distributions to represent the privacy characteristics of a mechanism, we can consolidate them into a privacy loss distribution (PLD) Dwork and Rothblum (2016).
+As we will see, the PLD possesses very desirable composition properties and is a practical basis for privacy loss computations.
 To formalize the PLD, we require likelihood ratios.
-Recall that, for a distribution $P$ absolutely continuous with respect to a distribution $Q$ ($P\ll Q$) on a probability domain $\Omega$, their likelihood ratio is the Radon-Nikodym derivative $\ell(\omega):=\frac{dP}{dQ}(\omega)$.
+Recall that, for a distribution $P$ absolutely continuous with respect to a distribution $Q$ ($P\ll Q$) on a probability domain $\Omega$, their likelihood ratio is the Radon-Nikodym derivative $\frac{dP}{dQ}$.
 However, in some cases, we would like to consider private mechanisms that violate absolute continuity and still construct the PLD.
-This can be achieved by taking any measure $\nu$ on $\Omega$ such that $P,Q\ll\nu$ (e.g.
-$\nu:=P+Q$) and setting $\ell(\omega):=\frac{dP}{d\nu}(\omega)/\frac{dQ}{d\nu}(\omega)$ ($\cdot/0:=\infty$).
-This construction of the likelihood ratio is unique only outside of a null-measure set [^1] More carefully, assume $P,Q\ll\nu,\nu^{\prime}$ and set $\mu:=\nu+\nu^{\prime}$ so that $P,Q\ll\nu,\nu^{\prime}\ll\mu$.
-By Radon–Nikodym and the chain rule for measures, we have, for $\omega\sim P$, that $\frac{dP}{d\nu}(\omega)/\frac{dQ}{d\nu}(\omega)=\frac{dP}{d\mu}(\omega)/\frac{dQ}{d\mu}(\omega)=\frac{dP}{d\nu^{\prime}}(\omega)/\frac{dQ}{d\nu^{\prime}}(\omega)$ with probability $1$., which is enough to construct a unique PLD.
-For this reason, we denote by $\frac{dP}{dQ}$ the likelihood ratio of $P$ with respect to $Q$ even when $P\not\ll Q$.
+In this case, we consider the Lebesgue decomposition of $P=P_{\ll}+P_{\bot}$ into absolutely continuous and singular components.
+For points in the support of the singular measure $P_{\bot}$, we assign likelihood $\infty$ and, for points outside, we may safely assign the likelihood $dP_{\ll}/dQ$.
+For simplicity, we denote this extended likelihood ratio as $dP/dQ$.
 Definition 3.
 Let $P$ and $Q$ be distributions on $\Omega$.
-The PLD of $(P,Q)$, denoted $\operatorname{PLD}(P\parallel Q)$, is the distribution over $\mathbb{R}\cup\{\infty\}$ of $\log(\frac{dP}{dQ}(\omega)),\omega\sim P$.
+The PLD of $(P,Q)$, denoted $\operatorname{PLD}(P\parallel Q)$, is the distribution over $\overline{\mathbb{R}}:=\mathbb{R}\cup\{\pm\infty\}$ of $\log(\frac{dP}{dQ}(\omega)),\omega\sim P$.
 For shorthand, we denote by $\operatorname{PLD}(M):=\operatorname{PLD}(M(D_{1})\parallel M(D_{2}))$ the PLD of a mechanism $M$. Note that a distribution can be easily recognized as a PLD as follows.
 Proposition 1.
-A distribution $L$ on $\mathbb{R}\cup\{\infty\}$ is the PLD of some $(P,Q$) if and only if it satisfies $\mathop{\mathbb{E}}_{Z\sim L}[e^{-Z}]\leq 1$.
+A distribution $L$ on $\overline{\mathbb{R}}$ is the PLD of some $(P,Q$) if and only if it satisfies $\mathop{\mathbb{E}}_{Z\sim L}[e^{-Z}]\leq 1$.
 In this case, one such pair is $(L,L^{\prime})$ where $L^{\prime}$ is the Esscher tilt of $L$, namely $dL^{\prime}(z):=e^{-z}dL(z)$ for $z\in\mathbb{R}$ and $L^{\prime}(\{-\infty\}):=1-\mathop{\mathbb{E}}_{Z\sim L}[e^{-Z}]$.
@@
 It is straightforward to show that this order is well-defined regardless of the underlying representations $(P,Q)$ and $(P^{\prime},Q^{\prime})$.
-PLDs also form a monoid under convolution that is monotone in the Blackwell order.
+When equipped with the Blackwell order, PLDs also form a partially-ordered commutative monoid under convolution.
 Proposition 2.
-For any PLDs $L_{1},L_{2},L_{3},L^{\prime}_{1},L^{\prime}_{2}$,
+For any PLDs $L_{1},L_{2},L_{3},L^{\prime}_{1}$,
   1. 1.
@@
   2. 2.
-    The identically zero distribution Id is a PLD such that $L_{1}\oplus\text{Id}=\text{Id}\oplus L_{1}=L_{1}$;
+    The identically zero distribution Id and the identically $\infty$ distribution, also denoted $\infty$, are PLDs such that $\text{Id}\preceq L_{1}\preceq\infty$, $L_{1}\oplus\text{Id}=L_{1}$, and $L_{1}\oplus\infty=\infty$;
   3. 3.
-    $(L_{1}\oplus L_{2})\oplus L_{3}=L_{1}\oplus(L_{2}\oplus L_{3})$; and
+    $(L_{1}\oplus L_{2})\oplus L_{3}=L_{1}\oplus(L_{2}\oplus L_{3})$;
   4. 4.
-    If $L_{1}\preceq L^{\prime}_{1}$ and $L_{2}\preceq L^{\prime}_{2}$, then $L_{1}\oplus L_{2}\preceq L^{\prime}_{1}\oplus L^{\prime}_{2}$.
+    $L_{1}\oplus L_{2}=L_{2}\oplus L_{1}$; and
+  5. 5.
+    If $L_{2}\preceq L^{\prime}_{2}$, then $L_{1}\oplus L_{2}\preceq L_{1}\oplus L^{\prime}_{2}$.
 A highly prized property of PLDs is that the privacy characteristics of an adaptive mechanism corresponds to convolution of PLDs, provided that each component fixes a privacy budget in advance.
-This is important because the convolution of a long sequence of distributions can be computed very efficiently by applying the fast Fourier transform [koskela2020computing].
-Note that we will require more than this to understand fully adaptive composition where the privacy bound of each component is itself adaptive.
+This is important because the convolution of a long sequence of distributions can be computed very efficiently by applying the fast Fourier transform Koskela et al.
+(2020).
+Note that we will require more than this to understand fully adaptive composition.
 Proposition 3.
-Let $M_{1}:\mathcal{D}\to\mathcal{Y}_{1}$ and $M_{2}:\mathcal{D}\times\mathcal{Y}_{1}\to\mathcal{Y}_{2}$ be adaptive mechanisms such that $\operatorname{PLD}(M_{1})\preceq L_{1}$ and $\operatorname{PLD}(M_{2}(\cdot;y_{1}))\preceq L_{2}$ for every $y_{1}\in\mathcal{Y}_{1}$.
-Then $\operatorname{PLD}(M_{1}\otimes M_{2})\preceq L_{1}\oplus L_{2}$.
-Moreover, if $\operatorname{PLD}(M_{1})\preceq\operatorname{PLD}(M^{\prime}_{1})$ and $\operatorname{PLD}(M_{2}(\cdot;y_{1}))\preceq\operatorname{PLD}(M^{\prime}_{2}(\cdot;y_{1}))$ for every $y_{1}\in\mathcal{Y}_{1}$, then $\operatorname{PLD}(M_{1}\otimes M_{2})\preceq\operatorname{PLD}(M^{\prime}_{1}\otimes M^{\prime}_{2})$.
-Closely related to the PLD is the hockey-stick curve: a convex reparameterization of the privacy profile of a mechanism generalizing the classical notion of $(\varepsilon,\delta)$-DP.
+Let $M_{1}:\mathcal{D}\to\mathcal{Y}_{1}$, $M^{\prime}_{1}:\mathcal{D}\to\mathcal{Y}^{\prime}_{1}$, $M_{2}:\mathcal{D}\times\mathcal{Y}_{1}\to\mathcal{Y}_{2}$, and $M^{\prime}_{2}:\mathcal{D}\times\mathcal{Y}^{\prime}_{1}\to\mathcal{Y}^{\prime}_{2}$ be adaptive mechanisms such that $\operatorname{PLD}(M_{1})\preceq\operatorname{PLD}(M^{\prime}_{1})$ and $\operatorname{PLD}(M_{2}(\cdot;y_{1}))\preceq\operatorname{PLD}(M^{\prime}_{2}(\cdot;y_{1}))$ for every $y_{1}\in\mathcal{Y}_{1}$.
+Then $\operatorname{PLD}(M_{1}\otimes M_{2})\preceq\operatorname{PLD}(M^{\prime}_{1}\otimes M^{\prime}_{2})$.
+Moreover, if $\operatorname{PLD}(M_{1})=L_{1}$ and $\operatorname{PLD}(M_{2}(\cdot;y_{1}))=L_{2}$ for every $y_{1}\in\mathcal{Y}_{1}$, then $\operatorname{PLD}(M_{1}\otimes M_{2})=L_{1}\oplus L_{2}$.
+We also define a few important classes of PLDs that will feature later in our discussion.
+Definition 4.
+We denote by $\mathcal{L}_{\mathrm{all}}$ the set of all PLDs and by $\mathcal{L}_{\mathrm{sym}}:=\{\operatorname{PLD}(P\parallel Q):(P,Q)\text{ symmetric}\}$ the set of all symmetric PLDs.
+Definition 5.
+For $\varepsilon\geq 0$, $\delta\in[0,1]$, let
+$R_{\varepsilon,\delta}:=\operatorname{PLD}(\delta 1_{\bot}+(1-\delta)\textrm{Bern}(1/(1+e^{\varepsilon}))\parallel\delta 1_{\top}+(1-\delta)\textrm{Bern}(1/(1+e^{-\varepsilon})))$
+We say that $L\in\mathcal{L}_{\mathrm{all}}$ satisfies $(\varepsilon,\delta)$-DP if $L\preceq R_{\varepsilon,\delta}$ or $\varepsilon$-DP if $L\preceq R_{\varepsilon}:=R_{\varepsilon,0}$.
+We denote $\mathcal{L}_{\mathrm{pure}}:=\{R_{\varepsilon}:\varepsilon\geq 0\}$ and $\mathcal{L}_{\mathrm{approx}}:=\{R_{\varepsilon,\delta}:\varepsilon\geq 0,\delta\in[0,1]\}$.
+Definition 6.
+For $\mu\geq 0$, we define $G_{\mu}:=\operatorname{PLD}(N(0,1)\parallel N(\mu,1))$ and say that $L\in\mathcal{L}_{\mathrm{all}}$ satisfies $\mu$-GDP when $L\preceq G_{\mu}$.
+We denote $\mathcal{L}_{\mathrm{gdp}}:=\{G_{\mu}:\mu\geq 0\}$.
+Definition 7.
+Lastly, we say that $L\in\mathcal{L}_{\mathrm{all}}$ satisfies $(\alpha,\rho$)-RDP if $\mathop{\mathbb{E}}_{Z\sim L}[e^{(\alpha-1)Z}]\leq e^{(\alpha-1)\rho}$.
+All three forms of privacy offer simple composition guarantees: For pure DP, $R_{\varepsilon_{1}}\oplus R_{\varepsilon_{2}}\preceq R_{\varepsilon_{1}+\varepsilon_{2}}$ Dwork et al.
+(2006b), for GDP $G_{\mu_{1}}\oplus G_{\mu_{2}}=G_{\sqrt{\mu_{1}^{2}+\mu_{2}^{2}}}$ Dong et al.
+(2019), and given $L_{1}$ satisfying $(\alpha,\rho_{1})$-RDP and $L_{2}$ satisfying $(\alpha,\rho_{2})$-RDP, $L_{1}\oplus L_{2}$ satisfies $(\alpha,\rho_{1}+\rho_{2})$-RDP Mironov (2017).
+More discussion of symmetry can be found in Appendix˜C.
+Closely related to the PLD is the hockey-stick curve: a convex reparameterization of the privacy profile, which is in turn a generalization of $(\varepsilon,\delta)$-DP.
 Definition 8.
@@
 $H_{x}(P\parallel Q):=\sup_{E\subseteq\Omega}P(E)-xQ(E).$
-Again, for convenience, we write $H_{x}(M):=H_{x}(M(D_{1})\parallel M(D_{2}))$ and sometimes $h_{M}(x):=H_{x}(M)$ for the hockey-stick curve of a mechanism $M$.
-We will rely on hockey-stick curves in order to reason about adversarial behaviour of adaptive mechanisms.
-The hockey-stick curve is very closely related to the classical $(\varepsilon,\delta)$ form of differential privacy.
-In particular, by construction of the hockey-stick divergence a mechanism $M$ satisfies $(\varepsilon,\delta)$-DP exactly when $H_{e^{\varepsilon}}(M)\leq\delta$.
+For convenience, we write $H_{x}(M):=H_{x}(M(D_{1})\parallel M(D_{2}))$ and sometimes $h_{M}(x):=H_{x}(M)$ for the hockey-stick curve of a mechanism $M$.
+The hockey-stick curve is very closely related to the classical $(\varepsilon,\delta)$ form of differential privacy: a mechanism $M$ satisfies $(\varepsilon,\delta)$-DP exactly when $H_{e^{\varepsilon}}(M)\leq\delta$.
 In general, hockey-stick curves are characterized by a few simple conditions.
-Proposition 4 ([4] Lemma 9).
-A curve $h:\mathbb{R}^{\times}\to[0,1]$ is a hockey-stick curve of some pair $(P,Q)$ if and only if $h$ is convex and decreasing such that $\lim_{x\to 0}h(x)=1$ and $h(x)\geq 1-x$ for all $x\in\mathbb{R}^{\times}$.
-Hockey-stick curves are also closely related to PLDs and can be fully recovered from the PLD via the following formula.
+Proposition 4 (Zhu et al. (2022) Lemma 9).
+A curve $h:\mathbb{R}^{\times}\to[0,1]$ is the hockey-stick curve of some $(P,Q)$ iff $h$ is convex and decreasing with $\lim_{x\to 0}h(x)=1$ and $h(x)\geq 1-x$ for all $x\in\mathbb{R}^{\times}$.
+Hockey-stick curves are also closely related to PLDs and can be fully recovered from the PLD via the following formula. The proof is in Appendix˜A. Note that we denote $(t)_{+}:=\max\{0,t\}$.
 Proposition 5.
-For any mechanism $M$ and $x\in\mathbb{R}^{\times}$,
-$h_{M}(x)=\mathop{\mathbb{E}}_{Z\sim\operatorname{PLD}(M)}[(1-xe^{-Z})_{+}]$
-where $(t)_{+}:=\max\{0,t\}$.
-Proof.
-For any pair $(P,Q)$ defined on $\Omega$ with likelihood ratio $\ell$, we have that
-$H_{x}(P\parallel Q)$ $=\sup_{E\subseteq\Omega}P(E)-xQ(E)$
-$=\sup_{E\subseteq\Omega}\int_{E}\frac{d(P-xQ)}{dP}(\omega)\,dP(\omega)$
-$=\sup_{E\subseteq\Omega}\int_{E}1-x/\frac{dP}{dQ}(\omega)\,dP(\omega)$
-$=\int_{\Omega}\left(1-x/\frac{dP}{dQ}(\omega)\right)_{+}\,dP(\omega)$
-$=\int_{\Omega}(1-xe^{-\log(\ell(\omega))})_{+}\,dP(\omega)$
-$=\mathop{\mathbb{E}}_{Z\sim\operatorname{PLD}(P\parallel Q)}[(1-xe^{-Z})_{+}].$
+For any pair $(P,Q)$ with likelihood $\ell$ and $x\in\mathbb{R}^{\times}$,
+$H_{x}(P\parallel Q)=\mathop{\mathbb{E}}_{Y\sim P}[(1-x/\ell(Y))_{+}]=\mathop{\mathbb{E}}_{Z\sim\operatorname{PLD}(P\parallel Q)}[(1-xe^{-Z})_{+}].$
 For convenience, we will sometimes write $h_{L}$ to denote the unique hockey-stick curve associated to a PLD $L$, namely $h_{L}(x):=\mathop{\mathbb{E}}_{Z\sim L}[(1-xe^{-Z})_{+}]$.
 Moreover, like PLDs, hockey-stick curves also capture adaptive composition in a natural way.
+The proof is also in Appendix˜A.
 Proposition 6.
@@
 $h_{M_{1}\otimes M_{2}}(x)=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[h_{M_{2}(\cdot;Y_{1})}(x/\ell^{1}(Y_{1}))].$
-Proof.
-Let $\ell$ denote the likelihood ratio for $M_{1}\otimes M_{2}$ and let $\ell^{2}_{y_{1}}$ denote the likelihood ratio for $M_{2}(\cdot;y_{1})$.
-By Bayes’ rule, we have $\ell(y_{1},y_{2})=\ell^{1}(y_{1})\cdot\ell^{2}_{y_{1}}(y_{2})$, so we have
-$\operatorname{PLD}(M_{1}\otimes M_{2})\equiv\log(\ell^{1}(Y_{1}))+\log(\ell^{2}_{Y_{1}}(Y_{2})),Y_{1}\sim M_{1}(D_{1}),Y_{2}\sim M_{2}(D_{1};Y_{1}).$
-By Proposition˜5 and the law of total expectation, it follows that
-$h_{M_{1}\otimes M_{2}}(x)$ $=\mathop{\mathbb{E}}_{Z\sim\operatorname{PLD}(M_{1}\otimes M_{2})}[(1-xe^{-Z})_{+}]$
-$=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[\mathop{\mathbb{E}}_{Y_{2}\sim M_{2}(D_{1};Y_{1})}[(1-xe^{-(\log(\ell^{1}(Y_{1}))+\log(\ell^{2}_{Y_{1}}(Y_{2})))})_{+}]]$
-$=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[\mathop{\mathbb{E}}_{Y_{2}\sim M_{2}(D_{1};Y_{1})}[(1-x/\ell^{1}(Y_{1})\cdot e^{-\log(\ell^{2}_{Y_{1}}(Y_{2}))})_{+}]]$
-$=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[\mathop{\mathbb{E}}_{Z_{2}\sim\operatorname{PLD}(M_{2}(\cdot;Y_{1})}[(1-x/\ell^{1}(Y_{1})\cdot e^{-Z_{2}})_{+}]]$
-$=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[h_{M_{2}(\cdot;Y_{1})}(x/\ell^{1}(Y_{1}))].$
-Our last characterization of privacy is given the Type I/Type II error tradeoff curve, also known as the $f$-DP framework of privacy [1].
+Our last characterization of privacy is given the Type I/Type II error tradeoff curve, also known as the $f$-DP framework of privacy Dong et al. (2019).
 Definition 9.
-Let $(P,Q)$ be any distributions on $\Omega$.
-For any hypothesis test $\phi:\Omega\to\{P,Q\}$ for distinguishing $P$ and $Q$, we call $\alpha_{\phi}:=\mathbb{P}_{\omega\sim P}[\phi(\omega)=Q]$ its Type I error and $\beta_{\phi}:=\mathbb{P}_{\omega\sim Q}[\phi(\omega)=P]$ its Type II error.
+Let $(P,Q)$ be distributions on $\Omega$.
+For a randomized hypothesis test $\phi:\Omega\to\Delta(\{P,Q\})$, we call $\alpha_{\phi}:=\mathbb{P}_{\omega\sim P,\mu\sim\phi(\omega)}[\mu=Q]$ its Type I error and $\beta_{\phi}:=\mathbb{P}_{\omega\sim Q,\mu\sim\phi(\omega)}[\mu=P]$ its Type II error.
 The tradeoff curve for $(P,Q)$ is defined as
 $T_{\alpha}(P\parallel Q):=\inf_{\phi}\{\beta_{\phi}:\alpha_{\phi}\leq\alpha\}.$
-For convenience, we sometimes write the tradeoff curve of a mechanism $M$ with respect to the fixed dataset pair $(D_{1},D_{2})$ as $T_{\alpha}(M):=T_{\alpha}(M(D_{1})\parallel M(D_{2}))$ or sometimes as $\tau_{M}(\alpha):=T_{\alpha}(M)$ where appropriate.
 In general, we can characterize valid tradeoff curves as follows.
-Proposition 7 ([1] Proposition 2.2).
-A curve $\tau:[0,1]\to[0,1]$ is a tradeoff curve for some testing problem $(P,Q)$ exactly when $\tau$ is continuous, convex, decreasing, and $\tau(\alpha)\leq 1-\alpha$ for all $\alpha\in[0,1]$.
-Like hockey-stick curves, it is enough to understand the PLD of a testing problem in order to obtain its tradeoff curve. The following formula can be derived by the well-known Neyman–Pearson lemma.
+Proposition 7 (Dong et al. (2019) Proposition 2.2).
+A curve $\tau:[0,1]\to[0,1]$ is the tradeoff curve for some $(P,Q)$ iff $\tau$ is continuous, convex, decreasing, and $\tau(\alpha)\leq 1-\alpha$ for all $\alpha\in[0,1]$.
+As with hockey-stick curves, the PLD determines the tradeoff curve.
+The well-known Neyman–Pearson lemma states that the optimal testing rule thresholds the PLD.
+For convenience, given $z^{*}\in\overline{\mathbb{R}}$, we call $\psi:\overline{\mathbb{R}}\to[0,1]$ a $z^{*}$-threshold when $\psi(z)=0$ for $z<z^{*}$ and $\psi(z)=1$ for $z>z^{*}$.
 Proposition 8.
-Let $(P,Q)$ be a pair of distributions with tradeoff curve $\tau(\alpha):=T_{\alpha}(P\parallel Q)$ and let $F$ denote the CDF of $\operatorname{PLD}(P\parallel Q)$. Then
-$\tau(\alpha)=\mathbb{E}_{Z\sim\operatorname{PLD}(P\parallel Q)}[1(F(Z)>\alpha)\cdot e^{-Z}]$
-It is common to represent a pair of distributions by either their hockey-stick curve or their tradeoff function.
+Let $(P,Q)$ be a pair of distributions and let $F$ be the CDF of $\operatorname{PLD}(P\parallel Q)$.
+Then, for every $\alpha\in[0,1]$, setting $z_{\alpha}:=\min\{z\in\overline{\mathbb{R}}:F(z)\geq\alpha\}$, there is a $z_{\alpha}$-threshold $\psi_{\alpha}$ such that
+$\tau(\alpha)=\mathop{\mathbb{E}}_{Z\sim\operatorname{PLD}(P\parallel Q)}[\psi_{\alpha}(Z)e^{-Z}]$
+and such that $\mathop{\mathbb{E}}_{Z\sim\operatorname{PLD}(P\parallel Q)}[\psi_{\alpha}(Z)]=1-\alpha$.
+Moreover, if $\alpha=F(z^{*})$, then, for $Z\sim\operatorname{PLD}(P\parallel Q)$, $\psi_{\alpha}(Z)=1(Z>z^{*})$ almost surely.
+For convenience, we denote the tradeoff curve corresponding to a PLD $L$ as $\tau_{L}$.
+It is common to represent the privacy characteristics of a mechanism by either its hockey-stick curve or its tradeoff curve.
 We show that there is a natural link between these representations via inversion and convex conjugacy.
@@
 Note that we can extend tradeoff functions to the real-line by setting them to $\infty$ outside their support $[0,1]$.
-Note that a special case of the following formula appears in [1] (Prop 2.12) for symmetric tradeoff functions but this generalization is new to the best of our knowledge.
+Note that a special case of the following formula appears in Dong et al.
+(2019) (Prop 2.12) for symmetric tradeoff functions but this generalization is new to the best of our knowledge.
+The proof is in Appendix˜A
 Proposition 9.
@@
 $h(x)=1+(\tau^{-1})^{*}(-x)$
-Proof.
-First, let $L$ denote the PLD of $(P,Q)$ and let $F$ be its CDF. By Proposition˜8,
-$\tau(\alpha)=\mathbb{E}_{Z\sim L}[1(Z>F^{-1}(\alpha))\cdot e^{-Z}].$
-Therefore, by Proposition˜5
-$h(x)$ $=\mathbb{E}_{Z\sim L}[(1-xe^{-Z})_{+}]$
-$=\mathbb{E}_{Z\sim L}[1(Z>\log(x))\cdot\underbrace{(1-xe^{-Z})}_{>0\iff Z>\log(x)}]$
-$=\sup_{\varepsilon\in\mathbb{R}}\mathbb{E}_{Z\sim L}[1(Z>\varepsilon)\cdot(1-xe^{-Z})]$
-$=\sup_{\alpha\in[0,1]}\mathbb{E}_{Z\sim L}[1(Z>F^{-1}(\alpha))\cdot(1-xe^{-Z})]$
-$=\sup_{\alpha\in[0,1]}1-\alpha-x\tau(\alpha)$
-$=\sup_{\alpha^{\prime}\in[0,1]}1-\tau^{-1}(\alpha^{\prime})-x\alpha^{\prime}$ ($\alpha^{\prime}=\tau^{-1}(\alpha)$)
-$=1+(\tau^{-1})^{*}(-x).$
 It follows as a natural consequence that the Blackwell order on pairs, the ordering induced by the tradeoff function, as well as the hockey-stick induced ordering are all equivalent.
@@
 Let $(P,Q)$ and $(P^{\prime},Q^{\prime})$ be pairs of distributions with tradeoff functions $\tau$ and $\tau^{\prime}$ respectively as well as hockey-stick curves $h$ and $h^{\prime}$ respectively.
-The following are equivalent
-  1. (i)
-    $(P,Q)\preceq(P^{\prime},Q^{\prime})$
-  2. (ii)
-    $\tau\succeq\tau^{\prime}$
-  3. (iii)
-    $h\preceq h^{\prime}$
-The equivalence of (i) and (ii) is given by a celebrated theorem of Blackwell [Bla51, Theorem 10].
+Then
+$(P,Q)\preceq(P^{\prime},Q^{\prime})\iff\tau\succeq\tau^{\prime}\iff h\preceq h^{\prime}.$
+The equivalence of (i) and (ii) is given by a celebrated theorem of Blackwell (Blackwell, 1951, Theorem 10).
 As for the equivalence of (ii) and (iii), this now follows by applying Fenchel–Moreau duality to Proposition˜9 and recalling that convex conjugation is an order-reversing operation.
@@
 Proposition 11.
-For any non-empty family of PLDs $\mathcal{L}$ dominated by at least one PLD, there exists a unique PLD $\sup\mathcal{L}$ that dominates $\mathcal{L}$ but is dominated by every upper bound for $\mathcal{L}$.
-In this case,
+For any non-empty family of PLDs $\mathcal{L}$, there exists a unique PLD $\sup\mathcal{L}$ that dominates $\mathcal{L}$ but is dominated by every upper bound for $\mathcal{L}$. In this case,
   1. (i)
@@
 Part (ii) follows from noticing that all of the properties of tradeoff curves (see Proposition˜7) are also closed under the lower convex envelope operation.
-We rely on the supremum extensively in order to reason about natural privacy filters.
-In particular, the former characterization says that the supremum of PLDs is given by a pointwise supremum in the space of hockey-stick curves, which will be particularly useful for showing our main result.
-As far as we are aware, the supremum property of PLDs has not actually been documented in the literature.
+Notice that it may be the case that $\sup{\mathcal{L}}=\infty$.
+Also notice that we may construct infima of privacy losses in an analogous manner, namely by taking lower convex envelopes of hockey-stick curves or by taking pointwise suprema of tradeoff curves.
+We will not require infima for our work but we do rely on the supremum extensively in order to reason about natural privacy filters.
+As far as we are aware, the completeness property of PLDs has not actually been documented in the literature.
 3 Query-Restricted Adversaries
-In this section we develop a general theory of the privacy characteristics of adversarial adaptive mechanism when the queries the adversary may issue are restricted.
+In this section we develop a general theory of the privacy characteristics of an adversarial adaptive mechanism when the queries the adversary may issue are restricted.
 This will lead directly to a theory of natural filters as we will see in §5.
+For simplicity, we will assume throughout that an adversary may always “pass” its turn by issuing a query with no privacy loss.
+We will consider adversaries that issue finitely many queries as well as adversaries that issue a countable sequence of queries.
+We will abuse notation somewhat by writing such things as $M_{1}\otimes\dots\otimes M_{\infty}$ to denote the countable extension of the prefixes $M_{1}\otimes\dots\otimes M_{t}$ as $t\to\infty$.
 Definition 10.
-Let $k$ be a positive integer.
-We define a privacy rule of length $k$ as a map $\Gamma$ that takes a sequence of PLDs $(L_{1},\dots,L_{k^{\prime}})$ ($0\leq k^{\prime}<k$) and outputs some collection of PLDs including Id.
+Let $k\in\mathbb{Z}^{+}\cup\{\infty\}$.
+We define a privacy rule of length $k$ as a map $\Gamma$ that takes a sequence of PLDs $(L_{1},\dots,L_{t})$ ($0\leq t<k$) and outputs some collection of PLDs including Id.
 We define a $\Gamma$-adversary as an adaptively composed mechanism $M=M_{1}\otimes\dots\otimes M_{k}$ such that
-$L_{i}:=\operatorname{PLD}(M_{i}(\cdot;y_{1},\dots,y_{i-1}))\in\Gamma(L_{1},\dots,L_{i-1})$
+$L_{t}:=\operatorname{PLD}(M_{t}(\cdot;y_{1},\dots,y_{t-1}))\in\Gamma(L_{1},\dots,L_{t-1})$
 for every $(y_{1},\dots,y_{k-1})\in\mathcal{Y}_{1}\times\dots\times\mathcal{Y}_{k-1}$. We will write $\operatorname{Adv}(\Gamma)$ to denote the set of $\Gamma$-adversaries and we will denote by
@@
 the worst-case privacy loss of arbitrary $\Gamma$-adversaries.
-It turns out that the maximum privacy loss for an adversary respecting a rule $\Gamma$ occurs in some sense when the adversary first plays a move and then adversarially chooses a optimal strategy tuned to ensure that the overall strategy is optimal at a particular given point.
+Crucially, notice that that no single adversary fully realizes the privacy cost of the rule $\Gamma$.
+Recalling the $f$-DP interpretation of the supremum (see Proposition˜11), the privacy cost of the rule $\Gamma$ is in fact realized by a family of adversaries that minimize the false negative rate for every given confidence level.
+As the following core lemma reveals, it turns out that the maximum privacy cost of the rule $\Gamma$ for a given confidence level occurs when the adversary first plays the best move for this confidence level, observes the outcome, then depending on the likelihood of having observed this outcome recursively chooses its remaining strategy tuned to maximize the overall privacy cost.
 Note that by $\lambda$ we mean the empty sequence.
 Lemma 1.
-Let $\Gamma$ be a rule of length $k$. If $k=1$, then $\operatorname{PLD}(\Gamma)=\sup{\Gamma(\lambda)}$ and, for $k>1$, we have
-$\operatorname{PLD}(\Gamma)=\sup\{L\oplus\operatorname{PLD}(\Gamma_{L}):L\in\Gamma(\lambda)\}$
-where $\Gamma_{L_{0}}(L_{1},\dots,L_{k^{\prime}}):=\Gamma(L_{0},L_{1},\dots L_{k^{\prime}})$ denotes the rule of length of k1k-1 that fixes the first move played against $\Gamma$ to $L_{0}$.
+Let $\Gamma$ be a rule of length $k\in\mathbb{Z}^{+}\cup\{\infty\}$. If $k=1$, then $\operatorname{PLD}(\Gamma)=\sup{\Gamma(\lambda)}$ and, for $k>1$, we have
+$\operatorname{PLD}(\Gamma)=\sup\{L\oplus\operatorname{PLD}(\Gamma|_{L}):L\in\Gamma(\lambda)\}$
+where $\Gamma|_{L_{0}}(L_{1},\dots,L_{t}):=\Gamma(L_{0},L_{1},\dots L_{t})$ denotes the rule of length of k1k-1 that fixes the first move played against $\Gamma$ to $L_{0}$.
+We prove the result by passing to hockey-stick curves because the pointwise supremum is easier to analyze than the lower convex envelope.
+We will essentially show that a worst-case adversary first plays an optimal move with corresponding PLD $L$, observes the outcome $Y_{1}$, and adversarially chooses a followup in $\operatorname{Adv}(\Gamma|_{L})$ to ensure that the overall process “traces out” the hockey-stick curve of $L\oplus\operatorname{Adv}(\Gamma|_{L})$.
 Proof.
 The case where $k=1$ is immediate, so we assume $k>1$. We first show that
-$\operatorname{PLD}(\Gamma)\preceq\sup\{L\oplus\operatorname{PLD}(\Gamma_{L}):L\in\Gamma(\lambda)\}.$
-To that end, consider any $\Gamma$-adversary $M$.
-Decompose $M$ as the adaptive composition of $M_{1}(\cdot)$ with some $M^{\prime}(\cdot;y_{1})$ and let $L:=\operatorname{PLD}(M_{1})\in\Gamma(\lambda)$.
-Then, clearly, for every fixed $y_{1}$, $M^{\prime}(\cdot;y_{1})$ is a $\Gamma_{L}$-adversary and, in particular, $\operatorname{PLD}(M^{\prime}(\cdot;y_{1}))\preceq\operatorname{PLD}(\Gamma_{L})$.
-By Proposition˜3, we have that $\operatorname{PLD}(M)\preceq L\otimes\operatorname{PLD}(\Gamma_{L})$ and the desired inequality follows by taking suprema.
-More surprising is the reverse inequality, which we prove constructively by passing to hockey-stick curves.
-Let $L\in\Gamma(\lambda)$, let $\kappa\in\mathbb{R}^{\times}$, and let $\varepsilon>0$.
-Choose any mechanism $M_{1}$ that has PLD $L$ and let $\ell:=\frac{dM_{1}(x_{1})}{dM_{1}(x_{2})}$ denote the corresponding likelihood function.
-Moreover, for any $\kappa\in\mathbb{R}^{\times}$ we can choose by Proposition˜11, a $\Gamma_{L}$-adversary $M^{\prime}_{\kappa}$ such that
-$H_{\kappa}(M^{\prime}_{\kappa})\geq\sup_{M^{\prime\prime}\in\operatorname{Adv}(\Gamma_{L})}H_{\kappa}(M^{\prime\prime})-\varepsilon.$
-Now consider the $\Gamma$-adversary
-$M(x):=(Y_{1},Y^{\prime}),Y_{1}\sim M_{1}(x),Y^{\prime}\sim M^{\prime}_{\kappa/\ell(Y_{1})}(x).$
+$\operatorname{PLD}(\Gamma)\preceq\sup\{L\oplus\operatorname{PLD}(\Gamma|_{L}):L\in\Gamma(\lambda)\}.$
+To that end, consider any $\Gamma$-adversary $M=M_{1}\otimes M_{2}$ and let $L:=\operatorname{PLD}(M_{1})\in\Gamma(\lambda)$.
+Then, clearly, for every fixed $y_{1}$, $M_{2}(\cdot;y_{1})$ is a $\Gamma|_{L}$-adversary and, in particular, $\operatorname{PLD}(M_{2}(\cdot;y_{1}))\preceq\operatorname{PLD}(\Gamma|_{L})$.
+By Proposition˜3, $\operatorname{PLD}(M)\preceq L\oplus\operatorname{PLD}(\Gamma|_{L})$ and thus the desired inequality follows by taking suprema.
+More surprising is the reverse inequality, which we prove constructively using hockey-stick curves.
+Let $L\in\Gamma(\lambda)$, $x\in\mathbb{R}^{\times}$, and $\gamma>0$.
+Choose any mechanism $M_{1}$ with PLD $L$ and let $\ell_{1}:=\frac{dM_{1}(D_{1})}{dM_{1}(D_{2})}$ denote its likelihood function.
+Moreover, for any $x^{\prime}\in\mathbb{R}^{\times}$ we can choose by Proposition˜11, a $\Gamma|_{L}$-adversary $M^{x^{\prime}}$ such that $h_{M^{x^{\prime}}}(x^{\prime})\geq h_{M^{\star}}(x^{\prime})-\gamma$ where $\operatorname{PLD}(M^{\star})=\operatorname{PLD}(\Gamma|_{L})$.
+Now consider the $\Gamma$-adversary $M:=M_{1}\otimes M_{2}$ where $M_{2}(D;y_{1}):=M^{x/\ell_{1}(y_{1})}(D)$.
 By Proposition˜6, we get
-$h_{M}(\kappa)$ $=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(x_{1})}[h_{M^{\prime}_{\kappa/\ell(Y_{1})}}(\kappa/\ell(Y_{1}))]$
-$\geq\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(x_{1})}[\sup_{M^{\prime\prime}\in\operatorname{Adv}(\Gamma_{L})}h_{M^{\prime\prime}}(\kappa/\ell(Y_{1}))-\varepsilon]$
-$=(h_{M_{1}}\otimes\sup_{M^{\prime\prime}\in\operatorname{Adv}(\Gamma_{L})}h_{M^{\prime\prime}})(\kappa)-\varepsilon.$
-Again, taking suprema, we get that
-$\sup_{M\in\operatorname{Adv}(\Gamma)}h_{M}\succeq h_{M_{1}}\otimes\sup_{M^{\prime\prime}\in\operatorname{Adv}(\Gamma_{L})}h_{M^{\prime\prime}},$
-and thus $\operatorname{PLD}(\Gamma)\succeq L\otimes\operatorname{PLD}(\Gamma_{L})$, which completes the proof since $L$ was arbitrary. ∎
-Intuitively, this result says that a worst-case adversary plays a move with corresponding PLD $L$, observes the outcome $Y_{1}$, and adversarially chooses a followup in $\operatorname{Adv}(\Gamma_{L})$ to ensure that the overall process “traces out” the privacy curve of $\operatorname{Adv}(\Gamma_{L})$.
+$h_{M}(x)$ $=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[h_{M_{x/\ell_{1}(Y_{1})}}(x/\ell_{1}(Y_{1}))]$
+$\geq\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[h_{M^{\star}}(x/\ell_{1}(Y_{1}))-\gamma]$
+$=h_{M_{1}\otimes M^{\star}}(x)-\gamma.$
+Taking suprema, we get $\sup_{M\in\operatorname{Adv}(\Gamma)}h_{M}\succeq h_{M_{1}\otimes M^{\star}}$ and thus $\operatorname{PLD}(\Gamma)\succeq L\oplus\operatorname{PLD}(\Gamma|_{L})$ by Propositions˜11 and 3, which completes the proof since $L$ was arbitrary.
+∎
+A final tool we will require for analyzing adversaries respecting a privacy rule is that privacy rules of infinite length can be understood by passing to limits of finite privacy rules.
+Lemma 2.
+Let $\Gamma$ be a privacy rule of infinite length.
+For $k\in\mathbb{Z}^{+}$, let $\Gamma^{k}$ be the privacy rule of length $k$ denoting the restriction of $\Gamma$ to query sequences of length less than $k$.
+Then
+$\operatorname{PLD}(\Gamma)=\sup\{\operatorname{PLD}(\Gamma^{k}):k\in\mathbb{Z}^{+}\}.$
+Proof.
+It is immediate that $\operatorname{PLD}(\Gamma^{k})\preceq\operatorname{PLD}(\Gamma)$ for all $k\in\mathbb{Z}^{+}$ because any mechanism $M\in\operatorname{Adv}(\Gamma^{k})$ can be extended to a mechanism in $\operatorname{Adv}(\Gamma)$ with the same privacy loss by simply repeatedly issuing the Id query after the first $k$ rounds.
+As for the reverse inequality, we pass again to the hockey-stick curve.
+To this end, let $M\in\operatorname{Adv}(\Gamma)$, let $\ell$ denote the likelihood ratio for $M$, let $M^{k}$ denote the restriction of $M$ to the first $k$ outputs, and let $\ell^{k}$ denote the likelihood ratio of $M^{k}$.
+Denoting $y^{k}:=(y_{1},\dots,y_{k})$ for $y=(y_{1},y_{2},\dots)$, it is clear that $\ell(y)=\lim_{k\to\infty}\ell^{k}(y^{k})$.
+By Proposition˜5 and the dominated convergence theorem, we have
+$h_{M}(x)$ $=\mathop{\mathbb{E}}_{Y\sim M(D)}[(1-x/\ell(Y))_{+}]$
+$=\mathop{\mathbb{E}}_{Y\sim M(D)}[\lim_{k\to\infty}(1-x/\ell^{k}(Y^{k}))_{+}]$
+$=\lim_{k\to\infty}\mathop{\mathbb{E}}_{Y\sim M(D)}[(1-x/\ell^{k}(Y^{k}))_{+}]$
+$=\lim_{k\to\infty}\mathop{\mathbb{E}}_{Y\sim M^{k}(D)}[(1-x/\ell^{k}(Y))_{+}]$
+$=\lim_{k\to\infty}h_{M^{k}}(x)$
+$\leq\sup\{h_{M^{k}}(x):k\in\mathbb{Z}^{+}\}.$
+Since clearly $M^{k}\in\operatorname{Adv}(\Gamma^{k})$, the desired inequality follows from Proposition˜11. ∎
+4 Residue Filters
+In this section, we develop a new class of privacy filters capable of leveraging exact privacy accounting: the residue filters.
+We show that residue filters capture all of the filters available in the literature including the pure DP filter, the Rényi filter, the GDP filter, as well as the natural filter.
+Algorithm 1 UpdateFilter
+0: Analyst $\mathcal{A}$, update rule $U:\mathcal{B}\times\mathcal{L}\to\mathcal{B}$, budget $B\in\mathcal{B}$, dataset $D$
+for $i\leftarrow 1,2,\dots$ do
+  $\mathcal{A}(Y_{1},\dots,Y_{i-1})$ gives mechanism $M_{i}$ that has PLD $L\in\mathcal{L}$
+  if $L\preceq B$ then
+   yield $Y_{i}\sim M_{i}(D)$
+   $B\leftarrow U(B,L)$
+  else
+   break
+A key piece of structure that appears to enable existing privacy filters is the notion of a remaining privacy budget.
+This same filter structure can be outfitted naturally with exact accounting.
+For this, we require some simple ingredients: Let $\mathcal{B}$ be a set of PLDs representing legal budgets, let $\mathcal{L}$ be a family of allowed queries, and let $U:\mathcal{B}\times\mathcal{L}\to\mathcal{B}$ be a map taking a legal budget and query and producing an updated budget.
+For simplicity, we assume $\text{Id}\in\mathcal{B},\mathcal{L}$.
+Algorithm˜1 formalizes the interaction between analyst and the update filter with initial budget $B$.
+Now, notice that Algorithm˜1 can be captured succinctly by the recursive privacy rule
+$\operatorname{Update}_{U,B}(\lambda)$ $:=\{L\in\mathcal{L}:L\preceq B\}$
+$\operatorname{Update}_{U,B}|_{L_{1}}$ $:=\textrm{Update}_{U,U(B,L_{1})}.$
+Our key result is that privacy filters derived from budget updates are valid as long as the updates satisfy a natural residue condition.
+Informally, it is enough that every budget encountered during the interaction dominates the privacy cost of the next budget plus the cost of the next query.
+In this case, we say that the updated budget is a residue of the prior budget with respect to the query.
+Theorem 1.
+Fix an update rule $U:\mathcal{B}\times\mathcal{L}\to\mathcal{B}$ and suppose that, for every budget $B\in\mathcal{B}$ and every $L\in\mathcal{L}$ with $L\preceq B$, we have
+$U(B,L)\oplus L\preceq B.$ ($\star$)
+Then, for any initial budget $B\in\mathcal{B}$, the update filter Algorithm˜1 is free, i.e.
+$\operatorname{PLD}(\operatorname{Update}_{U,B})\preceq B$.
+In this case, we call the filter a residue filter.
+Proof.
+By Lemma˜2, it suffices to prove that the finite rule $\operatorname{Update}_{U,B}^{k}$ is free for all $B\in\mathcal{B}$ and $k\in\mathbb{Z}^{+}$ under the residue condition.
+We proceed by induction.
+For $k=1$, we have
+$\operatorname{PLD}(\operatorname{Update}_{U,B}^{k})=\sup\{L\in\mathcal{L}:L\preceq B\}\preceq B$
+for any $B\in\mathcal{B}$. Moreover, if this is true for $k\in\mathbb{Z}^{+}$, then by Lemma˜1 and by assumption,
+$\operatorname{PLD}(\operatorname{Update}_{U,B}^{k+1})$ $=\sup\{L\oplus\operatorname{PLD}(\operatorname{Update}_{U,B}^{k+1}|_{L}):L\in\mathcal{L},L\preceq B\}$
+$=\sup\{L\oplus\operatorname{PLD}(\operatorname{Update}_{U,U(B,L)}^{k}):L\in\mathcal{L},L\preceq B\}$
+$\preceq\sup\{L\oplus U(B,L):L\in\mathcal{L},L\preceq B\}\preceq B$
+for any $B\in\mathcal{B}$ as well. ∎
+It turns out that residue filters already capture the existing privacy filters in addition to the natural filter, as we will see.
+Remark 2.
+The GDP filter Koskela et al.
+(2022), which permits a sequence of queries that are $\mu_{1},\dots,\mu_{t}$-GDP as long as $\mu_{1}^{2}+\dots+\mu_{t}^{2}\leq\mu^{2}$ for a fixed budget $\mu$, is exactly the residue filter with initial budget $B=G_{\mu}$ and update rule $U(G_{\mu},L):=G_{\sqrt{\mu^{2}-\nu^{2}}}$ where $\nu\geq 0$ is minimal such that $L\preceq G_{\nu}$.
+Because
+$U(G_{\mu},L)+L\preceq G_{\sqrt{\mu^{2}-\nu^{2}}}+G_{\nu}=G_{\mu},$
+Theorem˜1 immediately yields that the GDP filter is free.
+Remark 3.
+The pure DP filter Rogers et al.
+(2016), which permits a sequence of queries that are $\varepsilon_{1},\dots,\varepsilon_{t}$-DP as long as $\varepsilon_{1}+\dots+\varepsilon_{t}\leq\varepsilon$ for a fixed budget $\varepsilon$, is exactly the residue filter with initial budget $B=R_{\varepsilon}$ and update rule $U(R_{\varepsilon},L):=R_{\varepsilon-\varepsilon_{L}}$ where $\varepsilon_{L}\geq 0$ is minimal such that $L\preceq R_{\varepsilon_{L}}$.
+Because
+$U(R_{\varepsilon},L)+L\preceq R_{\varepsilon-\varepsilon_{L}}+R_{\varepsilon_{L}}\preceq R_{\varepsilon},$
+Theorem˜1 immediately yields that the pure DP filter is free.
+Unlike the GDP and pure DP forms of differential privacy, Rényi DP Feldman and Zrnic (2021) cannot be certified exactly by domination by a canonical PLD, so the Rényi filter cannot directly be expressed as a residue filter.
+Nonetheless:
+Remark 4.
+The Rényi filter Feldman and Zrnic (2021), which permits a sequence of queries that are $(\alpha,\rho_{1}),\dots,(\alpha,\rho_{t})$-RDP as long as $\rho_{1}+\dots+\rho_{t}\leq\rho$ for a fixed budget $\rho$ and order $\alpha$, can be expressed as a budget maintaining filter satisfying a residue condition.
+The maintained budget is initially $\rho$ and, upon receiving an $(\alpha,\rho_{i})$-RDP query, the filter accepts the query as long as $\rho_{i}\leq\rho$ and updates the budget to $\rho-\rho_{i}$.
+By linearity properties of RDP, this update satisfies an analogous residue condition to ($\star$1) and hence the same techniques can be used to show that this filter is free.
+In particular, derivative filters including the zCDP and the advanced composition filter seen in e.g. Whitehouse et al. (2023) are also a form of residue filter.
+Perhaps more notable is that the natural filter, whenever it is free, can also be understood as a residue filter.
+We will discuss this in detail in §5.
+The key takeaway is that residue filters either directly or in essence subsume previous privacy filters as well as the natural filter itself.
+This is strongly suggestive evidence that filters based on update rules satisfying a residue-like condition should be viewed as a standard technique for constructing new privacy filters.
+We strengthen this claim by constructing a new filter for GDP budgets that strictly improves upon the GDP filter.
+Theorem 5.
+Consider the update rule
+$U(G_{\mu},L):=G_{\mu^{\prime}},\quad\mu^{\prime}\geq 0\text{ maximal such that }G_{\mu^{\prime}}\oplus L\preceq G_{\mu}$
+for GDP budgets $G_{\mu}$ and arbitrary queries $L\preceq G_{\mu}$.
+This rule satisfies the residue condition ($\star$1) and hence the update filter derived from it is a residue filter.
+We call this filter the GDP residue filter.
+(a) Nearly Gaussian query: gain of $\mu\approx 0.036$ GDP budget.
+(b) Nearly Gaussian query, low budget: gain of $\mu\approx 0.078$ GDP budget.
+(c) Non-Gaussian query: gain of $\mu\approx 0.147$ GDP budget.
+Figure 1: A naïve GDP filter directly approximates the query by GDP and determines a new budget $\mu_{naive}$.
+When composing $G_{\mu_{naive}}$ with the query (purple), a gap with the original budget shows that budget was wasted.
+The GDP residue filter instead searches for the largest valid residual GDP budget $\mu_{res}$.
+Composing $G_{\mu_{res}}$ with the query almost matches the budget: very little is wasted.
+It is immediate that this update rule satisfies the residue condition [^1] One can show carefully that the maximum exists by noticing that $\mathcal{L}_{\mathrm{gdp}}$ is totally-ordered and applying Lemma 3, which we develop in the next section..
+The update can be computed efficiently by binary search over $\mu^{\prime}\in[0,\mu]$.
+It is also clear that the GDP residue filter cannot produce an updated budget smaller than the standard GDP filter.
+What is perhaps less obvious is that the GDP residue filter is different from the GDP filter.
+Indeed, intuitively the GDP filter may significantly overestimate the cost of a query $L\preceq G_{\nu}$ if $L$ is very hard to approximate by $G_{\nu}$.
+On the other hand, even if $L$ itself is badly approximated by $G_{\nu}$, the residue class $\{B^{\prime}:B^{\prime}\oplus L\preceq G_{\mu}\}$ that remains when removing $L$ from $G_{\mu}$ may admit a much tighter approximation.
+Figure˜1 illustrates using $f$-DP curves the savings of our residual filter in three different privacy regimes.
+Each plot shows an initial GDP budget, a pure DP query, the updated budget under the naïve GDP filter update ($\mu_{naive}$), and the updated budget under our GDP residue filter $\mu_{res}$.
+Generally, the residue filter always outperforms the naïve filter: $\mu_{res}>\mu_{naive}$.
+When the query is well-approximated by GDP and the budget is much larger than the cost of the query, the gains are fairly small, though still meaningful (fig.˜1(a)).
+When the query is well-approximated by GDP but the budget is small, perhaps due to depletion from previous queries, the gains are much more significant (fig.˜1(b)).
+The gains are predictably most significant for queries that are poorly approximated by GDP (fig.˜1(c)).
+In each case, the residual GDP budget, when composed with the query, is very close to the starting budget, showing that most of the budget has in fact been put to use.
 5 Characterizing the Natural Filter
-Our goal is to study when the natural filter Algorithm˜1 with budget $B$ and queries in $\mathcal{L}$ comes at no additional privacy cost, i.e.
+We now turn toward the natural filter Algorithm˜2 with some budget $B$ and queries restricted to some set $\mathcal{L}$.
+Our goal is to understand when this filter comes at no additional privacy cost, i.e.
 the mechanism induced by the interaction between the filter and an adversarial analyst itself has privacy bounded by $B$.
 In this case, we say that the natural filter is free.
-We will write $L\ominus L^{\prime}:=\sup\{L^{\prime\prime}:L^{\prime}\oplus L^{\prime\prime}\preceq L\}$ so that $L^{\prime\prime}\preceq L\ominus L^{\prime}$ exactly when $L^{\prime}\oplus L^{\prime\prime}\preceq L$.
-We also write $\mathcal{L}_{\preceq L}:=\{L^{\prime}\in\mathcal{L}:L^{\prime}\preceq L\}$ to denote the queries in $\mathcal{L}$ dominated by $L$.
 For simplicity, we will assume that $\text{Id}\in\mathcal{L}$.
 Algorithm 2 NaturalFilter
-0: Privacy budget $L$, adversary (or analyst) $\mathcal{A}$, family of allowed PLDs $\mathcal{L}$, query capacity $k$
-$i=0$
-while $\mathcal{A}$ continues and $i<k$ do
-  $i\leftarrow i+1$
-  $\mathcal{A}$ gives mechanism $\mathcal{M}_{i}$ that has PLD $L_{i}\in\mathcal{L}$ $\triangleright$ can depend on previous results $Y_{<i}$
-  if $L_{1}\oplus\dots\oplus L_{i}\preceq L$ then
-   $\mathcal{A}$ receives $Y_{i}\sim\mathcal{M}_{i}$
-  end if
-end while
-return $(Y_{1},\ldots,Y_{i})$
-Now, notice that the natural filter Algorithm˜1 with budget $L$ and $k$ queries in $\mathcal{L}$ is free exactly when $\operatorname{PLD}(\operatorname{Filter}_{\mathcal{L},B,k})\preceq B$ where
-$\operatorname{Filter}_{\mathcal{L},B,k}(L_{1},\dots,L_{k^{\prime}}):=\mathcal{L}_{\preceq B\ominus L_{1}\dots\ominus L_{k^{\prime}}}=\{L_{k^{\prime}+1}\in\mathcal{L}:L_{1}\oplus\dots\oplus L_{k^{\prime}}\oplus L_{k^{\prime}+1}\preceq B\}$
-for $0\leq k^{\prime}<k$. In particular, by applying Lemma˜1 to the rule $\operatorname{Filter}_{\mathcal{L},B,k}$ we immediately get the following characterization.
+0: Budget $B$, analyst $\mathcal{A}$, legal query set $\mathcal{L}$, query capacity $k\in\mathbb{Z}^{+}\cup\{\infty\}$, dataset $D$
+for $t\leftarrow 1,\dots,k$ do
+  $\mathcal{A}(Y_{1},\dots,Y_{t-1})$ gives mechanism $M_{t}$ that has PLD $L_{t}\in\mathcal{L}$
+  if $L_{1}\oplus\dots\oplus L_{t}\preceq B$ then
+   yield $Y_{t}\sim M_{t}(D)$
+  else
+   break
+Now, notice that the natural filter Algorithm˜2 with budget $B$, query set $\mathcal{L}$, and capacity $k$ is free exactly when $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L},B}^{k})\preceq B$ where
+$\operatorname{Nat}_{\mathcal{L},B}(L_{1},\dots,L_{t}):=\{L_{t+1}\in\mathcal{L}:L_{1}\oplus\dots\oplus L_{t+1}\preceq B\}.$
+In particular, by applying Lemma˜1 to the rule $\operatorname{Nat}_{\mathcal{L},B}^{k}$, we immediately get the following characterization of the cost of interacting with the natural filter.
 Theorem 6.
-Let $\mathcal{L}$ be a family of PLDs, let $B$ be a PLD, and let $k$ be a positive integer.
-For any $L_{1},\dots,L_{k^{\prime}}\in\mathcal{L}$ ($0\leq k^{\prime}<k$) such that $L_{1}\oplus\dots\oplus L_{k^{\prime}}\preceq B$, let
-$V_{k-k^{\prime}}(L_{1},\dots,L_{k^{\prime}}):=\sup\{L_{k^{\prime}+1}\oplus V_{k-(k^{\prime}+1)}(L_{1},\dots,L_{k^{\prime}+1}):L_{k^{\prime}+1}\in\mathcal{L}_{B\ominus L_{1}\ominus\dots\ominus L_{k^{\prime}}}\}$
-where $V_{0}(L_{1},\dots,L_{k}):=\text{Id}$. The natural filter algorithm˜1 for $k$ queries in $\mathcal{L}$ with budget $B$ is free if and only if $V_{k}(\lambda)\preceq B$.
-Informally, $V_{k-k^{\prime}}(L_{1},\dots,L_{k^{\prime}})$ is the maximal remaining privacy cost of an adversary that has already played moves $L_{1},\dots,L_{k^{\prime}}$.
-Naturally, when the adversary has already selected all of its moves, the remaining privacy cost is Id, i.e.
-nothing.
-It is also of interest to understand when a given family of queries $\mathcal{L}$ simultaneously admits free natural filters against an arbitrary budget $B$.
-We show that this is the case exactly when the supremum commutes with the convolution operator.
+Let $\mathcal{L}\subseteq\mathcal{L}_{\mathrm{all}}$, let $B\in\mathcal{L}_{\mathrm{all}}$, and consider, for $L^{\prime}\preceq B$, the PLD operator
+$V_{k+1}(L^{\prime}):=\sup\{L\oplus V_{k}(L^{\prime}\oplus L):L\in\mathcal{L},L^{\prime}\oplus L\preceq B\}$
+where $V_{1}(L^{\prime}):=\sup\{L\in\mathcal{L}:L^{\prime}\oplus L\preceq B\}$.
+Then, for any $k\in\mathbb{Z}^{+}$ and legal $L_{1},\dots,L_{t}\in\mathcal{L}$, we have $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L},B}^{k}\mid_{L_{1},\dots,L_{t}})=V_{k-t}(L_{1}\oplus\dots\oplus L_{t})$ and, in particular, $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L},B}^{k})=V_{k}(\text{Id})$.
+Informally, $V_{k-t}(L_{1}\oplus\dots\oplus L_{t})$ is the maximal remaining privacy cost of an adversary that has already played moves $L_{1},\dots,L_{t}$.
+Naturally, when only one move is left to select, the effect of adaptivity vanishes and the adversary simply selects the valid move of maximal cost.
+It is also of interest to understand when a given family of queries $\mathcal{L}$ admits free natural filters against an arbitrary budget $B$.
+When a filter is free against all budgets, we say that the filter is universally free.
+In the case of the natural filter, we show that it is universally free exactly when it can be realized as a residue filter.
+For convenience, given $k\in\mathbb{Z}^{+}\cup\{\infty\}$ we let $\mathcal{L}^{<k}:=\{L_{1}\oplus\dots\oplus L_{t}:L_{1},\dots,L_{t}\in\mathcal{L},0\leq t<k\}$ denote the closure of $\mathcal{L}$ under fewer than $k$ convolutions and we let $L\oplus\mathcal{L}^{\prime}:=\{L\oplus L^{\prime}:L^{\prime}\in\mathcal{L}^{\prime}\}$ for a set of PLDs $\mathcal{L}^{\prime}$.
 Theorem 7.
-Fix a family of PLDs $\mathcal{L}$. The natural filter Algorithm˜1 is free for every budget $B$ simultaneously and query capacity $k$ if and only if $\mathcal{L}$ has the following property:
-$\forall L\in\mathcal{L}\ \textrm{and}\ \mathcal{L}^{\prime}\subseteq\mathcal{L}^{<k}\quad L\oplus\sup{\mathcal{L}^{\prime}}=\sup\{L\oplus L^{\prime}:L^{\prime}\in\mathcal{L}^{\prime}\}.$
+Fix $\mathcal{L}\subseteq\mathcal{L}_{\mathrm{all}}$, $k\in\mathbb{Z}^{+}\cup\{\infty\}$, and consider the update rule $U:\mathcal{L}_{\mathrm{all}}\times\mathcal{L}\to\mathcal{L}_{\mathrm{all}}$ given by $U(B,L):=\sup\{L^{\prime}\in\mathcal{L}^{<k}:L\oplus L^{\prime}\preceq B\}$.
+The following are equivalent:
+  1. (i)
+    For all $B\in\mathcal{L}_{\mathrm{all}}$, $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L},B}^{k})\preceq B$;
+  2. (ii)
+    For all $L\in\mathcal{L}$ and $\mathcal{L}^{\prime}\subseteq\mathcal{L}^{<k}$, $L\oplus\sup{\mathcal{L}^{\prime}}=\sup(L\oplus\mathcal{L}^{\prime})$;
+  3. (iii)
+    For all $B\in\mathcal{L}_{\mathrm{all}}$ and $L\in\mathcal{L}$, $U(B,L)\oplus L\preceq B$; and
+  4. (iv)
+    For all $B\in\mathcal{L}_{\mathrm{all}}$, $\operatorname{PLD}(\operatorname{Update}_{U,B}^{k})\preceq B$.
+In any of these cases, the natural filter $\operatorname{Nat}_{\mathcal{L},B}^{k}$ and the residue filter $\operatorname{Update}_{U,B}^{k}$ are identical.
 Proof.
-First, assume that the natural filter is free for every budget $B$ and for any query capacity.
-Now, let $L\in\mathcal{L}$ and $\mathcal{L}^{\prime}\subseteq\mathcal{L}^{<k}$.
-For every $L^{\prime}\in\mathcal{L}^{\prime}$, we have $L\oplus L^{\prime}\preceq L\oplus L\oplus\sup{\mathcal{L}^{\prime}}$ by Proposition˜2 and hence $\sup\{L\oplus\mathcal{L}^{\prime}\}\preceq L\oplus\sup{\mathcal{L}^{\prime}}$.
-As for the reverse inequality, consider the rule
-$\Gamma(\lambda):=\{L\}\quad\quad\Gamma(L):=\mathcal{L}^{\prime}$
-of length $2$ and the budget $B:=\sup(L\oplus\mathcal{L}^{\prime})$.
-Now, for any $M\in\operatorname{Adv}(\Gamma)$, notice that we can transform $M$ into $M^{\prime}\in\operatorname{Adv}(\operatorname{Filter}_{\mathcal{L},B,k})$ such that $\operatorname{PLD}(M)=\operatorname{PLD}(M^{\prime})$.
-Indeed, let $M=M_{1}\oplus M_{2}\in\operatorname{Adv}(\Gamma)$ and fix $y_{1}\in\mathcal{Y}_{1}$ so that we have that $\operatorname{PLD}(M_{2}(\cdot;y_{1}))=L_{1}\oplus\dots\oplus L_{k-1}\in\mathcal{L}^{<k}$.
-For each $L_{i}$, we can find a mechanism $M^{i}$ with $\operatorname{PLD}(M^{i})=L_{i}$.
-Now, set $M^{\prime}_{2}(\cdot;y_{1}):=M^{1}\otimes\dots\otimes M^{k-1}$ so that $\operatorname{PLD}(M^{\prime}_{2}(\cdot;y_{1}))=L_{1}\oplus\dots\oplus L_{k^{\prime}}=\operatorname{PLD}(M_{2}(\cdot;y_{1}))$.
-Clearly, $M^{\prime}:=M_{1}\otimes M^{\prime}_{2}\in\operatorname{Adv}(\operatorname{Filter}_{\mathcal{L},B,k})$ and, by Proposition˜3, we have $\operatorname{PLD}(M)=\operatorname{PLD}(M^{\prime})$, so by Lemma˜1 and since the natural filter is free for $B$ by assumption, we have
-$L\oplus\sup{\mathcal{L}^{\prime}}=\sup_{M\in\operatorname{Adv}(\Gamma)}\operatorname{PLD}(M)\preceq\sup_{M^{\prime}\in\operatorname{Adv}(\operatorname{Filter}_{\mathcal{L},B,k})}\operatorname{PLD}(M^{\prime})\preceq B=\sup\{L\oplus\mathcal{L}^{\prime}\}.$
-On the other hand, assume that convolution commutes with supremum in $\mathcal{L}^{\infty}$.
-Let $B$ be any budget and let $k$ be a positive integer.
-With an eye toward Theorem˜1, we claim that the maximal remaining privacy cost of an adversary that has selected some if its moves already satisfies
-$V_{k-k^{\prime}}(L_{1},\dots,L_{k^{\prime}})=\sup\{L_{k^{\prime}+1}\oplus\dots\oplus L_{k}:L_{k^{\prime}+1},\dots,L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
-for $0\leq k^{\prime}<k$. Indeed,
-$V_{1}(L_{1},\dots,L_{k-1})$ $=\sup\{L_{k}\oplus V_{0}(L_{1},\dots,L_{k}):L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
-$=\sup\{L_{k}:L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
-and, if the claim holds for $k^{\prime}$, then
-$V$ $_{k-(k^{\prime}-1)}(L_{1},\dots,L_{k^{\prime}-1})$
-$=\sup\{L_{k^{\prime}}\oplus V_{k-k^{\prime}}(L_{1},\dots,L_{k^{\prime}}):L_{k^{\prime}}\in\mathcal{L},\oplus_{i=1}^{k^{\prime}}L_{i}\preceq B\}$
-$=\sup\{L_{k^{\prime}}\oplus\sup\{L_{k^{\prime}+1}\oplus\dots\oplus L_{k}:L_{k^{\prime}+1},\dots,L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}:L_{k^{\prime}}\in\mathcal{L},\oplus_{i=1}^{k^{\prime}}L_{i}\preceq B\}$
-$=\sup\{\sup\{L_{k^{\prime}}\oplus L_{k^{\prime}+1}\oplus\dots\oplus L_{k}:L_{k^{\prime}+1},\dots,L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}:L_{k^{\prime}}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
-$=\sup\{L_{k^{\prime}}\oplus\dots\oplus L_{k}:L_{k^{\prime}},\dots,L_{k}\in\mathcal{L},\oplus_{i=1}^{k}L_{i}\preceq B\}$
-and the claim holds by induction. In particular,
-$V_{k}(\lambda)=\sup\{L_{1}\oplus\dots\oplus L_{k}:L_{1},\dots,L_{k}\in\mathcal{L},L_{1}\oplus\dots\oplus L_{k}\preceq B\}\preceq B$
-and so the natural filter is free for $k$ queries in $\mathcal{L}$ with budget $B$ by Theorem˜1. ∎
-Our key result is that that this commutativity condition can essentially only be satisfied for family of queries $\mathcal{L}$ whose closure under composition $\mathcal{L}^{\infty}$ is well-ordered.
+Without loss of generality, by Lemma˜2, assume that $k$ is finite.
+First, we show that (i) implies (ii), so suppose that the natural filter is free for every budget $B$ and let $L\in\mathcal{L}$ and $\mathcal{L}^{\prime}\subseteq\mathcal{L}^{<k}$.
+For every $L^{\prime}\in\mathcal{L}^{\prime}$, we have $L\oplus L^{\prime}\preceq L\oplus\sup{\mathcal{L}^{\prime}}$ by Proposition˜2 and hence $\sup(L\oplus\mathcal{L}^{\prime})\preceq L\oplus\sup{\mathcal{L}^{\prime}}$ for free.
+As for the reverse inequality, consider the length-$2$ rule $\Gamma(\lambda):=\{L\},\Gamma(L):=\mathcal{L}^{\prime}$ as well as the budget $B:=\sup(L\oplus\mathcal{L}^{\prime})$.
+Now, for any $M\in\operatorname{Adv}(\Gamma)$, notice that we can transform $M$ into $M^{\prime}\in\operatorname{Adv}(\operatorname{Nat}_{\mathcal{L},B}^{k})$ such that $\operatorname{PLD}(M)=\operatorname{PLD}(M^{\prime})$.
+Indeed, let $M=M_{1}\otimes M_{2}\in\operatorname{Adv}(\Gamma)$ and fix $y_{1}\in\mathcal{Y}_{1}$ so that $\operatorname{PLD}(M_{2}(\cdot;y_{1}))=L_{2,1}\oplus\dots\oplus L_{2,t}\in\mathcal{L}^{<k}$.
+For each $L_{2,i}$, we can find a mechanism $M_{2,i}$ with $\operatorname{PLD}(M_{2,i})=L_{2,i}$.
+Now, set $M^{\prime}_{2}(\cdot;y_{1}):=M_{2,1}\otimes\dots\otimes M_{2,t}$ so that $\operatorname{PLD}(M^{\prime}_{2}(\cdot;y_{1}))=L_{2,1}\oplus\dots\oplus L_{2,t}=\operatorname{PLD}(M_{2}(\cdot;y_{1}))$.
+Clearly, $M^{\prime}:=M_{1}\otimes M^{\prime}_{2}\in\operatorname{Adv}(\operatorname{Nat}_{\mathcal{L},B}^{k})$ and, by Proposition˜3, we have $\operatorname{PLD}(M)=\operatorname{PLD}(M^{\prime})$, so by Lemma˜1 and since the natural filter is free for $B$ by assumption, we have
+$L\oplus\sup{\mathcal{L}^{\prime}}=\sup_{M\in\operatorname{Adv}(\Gamma)}\operatorname{PLD}(M)\preceq\sup_{M^{\prime}\in\operatorname{Adv}(\operatorname{Nat}_{\mathcal{L},B}^{k})}\operatorname{PLD}(M^{\prime})\preceq B=\sup(L\oplus\mathcal{L}^{\prime}).$
+That (ii) implies (iii) is immediate from the construction of the update rule $U$ by taking $\mathcal{L}^{\prime}:=\{L^{\prime}\in\mathcal{L}^{<k}:L\oplus L^{\prime}\preceq B\}$; likewise (iii) implies (iv) follows immediately from Theorem˜1, so what remains to show is that, if the residue filter $\operatorname{Update}_{U,B}$ is universally free, then so is $\operatorname{Nat}_{\mathcal{L},B}$, in which case the two filters coincide.
+To this end, first notice that $\operatorname{Update}_{U,B}^{k}$ is always at least as permissive a filter as $\operatorname{Nat}_{\mathcal{L},B}^{k}$.
+Indeed, for any $L\in\mathcal{L}$ and $L^{\prime}\in\mathcal{L}^{<\infty}$ such that $L\oplus L^{\prime}\preceq B$, it follows immediately from the construction of $U$ that $L^{\prime}\preceq U(B,L)$.
+Thus for any legal query sequence $L_{1},\dots,L_{t}$,
+$L_{t+1}\in\operatorname{Nat}_{\mathcal{L},B}^{k}(L_{1},\dots,L_{t})$ $\implies L_{1}\oplus\dots\oplus L_{t+1}\preceq B$
+$\implies L_{2}\oplus\dots\oplus L_{t+1}\preceq U(B,L_{1})$
+$\implies\dots$
+$\implies L_{t+1}\preceq U(\dots U(B,L_{1})\dots,L_{t})$
+$\implies L_{t+1}\in\operatorname{Update}_{U,B}^{k}(L_{1},\dots,L_{t}).$
+Therefore, if $\operatorname{Update}_{U,B}^{k}$ is universally free, then so is $\operatorname{Nat}_{\mathcal{L},B}^{k}$.
+Finally, notice that, as long as the residue condition for $U$ is satisfied, then $\operatorname{Nat}_{\mathcal{L},B}^{k}$ is at least as permissive a filter as $\operatorname{Update}_{U,B}^{k}$.
+Indeed, for any $L\in\mathcal{L}$ and $L^{\prime}\in\mathcal{L}^{<\infty}$ for which $L^{\prime}\preceq U(B,L)$, Proposition˜2 and the residue condition ensure that $L\oplus L^{\prime}\preceq L\oplus U(B,L)\preceq B$.
+Inductively, just as before, this implies that for any legal query sequence $L_{1},\dots,L_{t}$, we have that $L_{t+1}\in\operatorname{Update}(U,B)^{k}(L_{1},\dots,L_{t})\implies L_{t+1}\in\operatorname{Nat}_{\mathcal{L},B}^{k}(L_{1},\dots,L_{t})$ and so the two filters are indeed equivalent.
+∎
+Our key result is that that these conditions can essentially only be satisfied for family of queries $\mathcal{L}$ whose closure under composition $\mathcal{L}^{<k}$ is totally-ordered.
+For a distribution $A$ over $\overline{\mathbb{R}}$, we denote by $\operatorname{supp}(A):=\{t\in\mathbb{R}:\forall\text{open }U\ni t,A(U)>0\}$ the support of $A$, namely the finite points with locally positive probability mass.
+We say that $A$ is non-degenerate if $|\operatorname{supp}(A)|>1$, i.e.
+$A$ is supported at at least two finite points.
 Theorem 8.
-Fix a family of non-degenerate PLDs $\mathcal{L}$, i.e.
-every $L\in\mathcal{L}$ is supported at at least two points in $\mathbb{R}$.
-The natural filter Algorithm˜1 is free for every budget $B$ and every query capacity $k$ if and only if $\mathcal{L}^{\infty}$ is well-ordered with respect to $\preceq$.
-We prove here that families of queries that are well-ordered under composition satisfy the commutativity condition.
-The proof of the other direction is significantly more involved, so we delay it until Section˜6.
-Proof of Sufficient Condition for Theorem˜3.
-Assume that $\mathcal{L}^{\infty}$ is well-ordered and let $L\in\mathcal{L},\mathcal{L}^{\prime}\subseteq\mathcal{L}^{<k}\subseteq\mathcal{L}^{\infty}$.
-For convenience, we will denote by $h_{L}$ the unique hockey-stick curve associated to a PLD $L$, namely $h_{L}(x)=\mathop{\mathbb{E}}_{Z\sim L}[(1-xe^{-Z})_{+}]$ (see Proposition˜5).
-Now, as in the proof of Theorem˜2, we already have $\sup\{L\oplus\mathcal{L}^{\prime}\}\preceq L\oplus\sup{\mathcal{L}^{\prime}}$ for free, so due to Proposition˜11, we just need to show that $h_{L\oplus\sup{\mathcal{L}^{\prime}}}\preceq h_{\sup\{L\oplus\mathcal{L}^{\prime}\}}$.
-To that end, let $\gamma>0$.
-By well-ordering of $\mathcal{L}^{\prime}$ and by Proposition˜11, we can find some $L^{\prime}\in\mathcal{L}^{\prime}$ so that $h_{\sup{\mathcal{L}^{\prime}}}\preceq h_{L^{\prime}}+\gamma$.
-In this case, we have
-$h_{L\oplus\sup{\mathcal{L}^{\prime}}}(x)$ $=\mathop{\mathbb{E}}_{Z\sim L}[\mathop{\mathbb{E}}_{Z^{\prime}\sim\sup{\mathcal{L}^{\prime}}}[(1-xe^{-(Z+Z^{\prime})})_{+}]]$
-$=\mathop{\mathbb{E}}_{Z\sim L}[h_{\sup{\mathcal{L}^{\prime}}}(xe^{-Z})]$
-$\leq\mathop{\mathbb{E}}_{Z\sim L}[h_{L^{\prime}}(xe^{-Z})+\gamma]$
-$=\mathop{\mathbb{E}}_{Z\sim L}[\mathop{\mathbb{E}}_{Z^{\prime}\sim L^{\prime}}[(1-xe^{-(Z+Z^{\prime})})_{+}]]+\gamma$
-$=h_{L\oplus L^{\prime}}(x)+\gamma$
-$\leq h_{\sup\{L\oplus L^{\prime}\}}(x)+\gamma$
-for any $x\in\mathbb{R}^{\times}$, which yields the desired result as $\gamma$ was arbitrary. ∎
+Let $\mathcal{L}\subseteq\mathcal{L}_{\mathrm{all}}$ and $k\in\mathbb{Z}^{+}\cup\{\infty\}$.
+If $\mathcal{L}^{<k}$ is totally-ordered, then $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L},B}^{k})\preceq B$ for every $B\in\mathcal{L}_{\mathrm{all}}$.
+Conversely, if the non-degenerate members of $\mathcal{L}^{<k}$ are not totally-ordered, then $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L},B}^{2k-3})\npreceq B$ for some choice of $B\in\mathcal{L}_{\mathrm{all}}$.
+Proving the forward direction is easy. We show in general that the supremum commutes with convolution for totally-ordered families of queries.
+Lemma 3.
+Let $L$ be a PLD and let $\mathcal{L}$ be a totally-ordered set of PLDs. Then $L\oplus\sup{\mathcal{L}}=\sup(L\oplus\mathcal{L})$.
+Proof.
+As in the proof of Theorem˜7, we already have $\sup(L\oplus\mathcal{L})\preceq L\oplus\sup{\mathcal{L}}$ for free, so due to Proposition˜11, we just need to show that $h_{L\oplus\sup{\mathcal{L}}}\preceq h_{\sup(L\oplus\mathcal{L})}$.
+By total-ordering of $\mathcal{L}$ and by Proposition˜11, we can find a sequence $L_{1}\preceq L_{2}\preceq\dots\in\mathcal{L}$ so that $h_{L_{t}}\to h_{\sup{\mathcal{L}}}$ pointwise.
+For any $x\in\mathbb{R}^{\times}$, monotone convergence yields
+$h_{L\oplus\sup{\mathcal{L}}}(x)$ $=\mathop{\mathbb{E}}_{Z\sim L}[\mathop{\mathbb{E}}_{Z^{\prime}\sim\sup{\mathcal{L}}}[(1-xe^{-(Z+Z^{\prime})})_{+}]]$
+$=\mathop{\mathbb{E}}_{Z\sim L}[h_{\sup{\mathcal{L}}}(xe^{-Z})]$
+$=\mathop{\mathbb{E}}_{Z\sim L}[\lim_{t\to\infty}h_{L_{t}}(xe^{-Z})]$
+$=\lim_{t\to\infty}\mathop{\mathbb{E}}_{Z\sim L}[h_{L_{t}}(xe^{-Z})]$
+$=\lim_{t\to\infty}\mathop{\mathbb{E}}_{Z\sim L}[\mathop{\mathbb{E}}_{Z^{\prime}\sim L_{t}}[(1-xe^{-(Z+Z^{\prime})})_{+}]]$
+$=\lim_{t\to\infty}h_{L\oplus L_{t}}(x)$
+$\leq h_{\sup(L\oplus\mathcal{L})}(x).$
+Proving the other direction is substantially more challenging.
+The key lemma is that, given a pair of non-degenerate queries that are not totally-ordered, composing one of these queries with their supremum can amplify the gaps between them, leading to a violation of the commutativity condition.
+The proof of this technical result involves analyzing the topological structure of the hockey-stick curves of unordered queries.
+The details are given in Section˜8.
+Note that non-degeneracy is key as the following result fails for degenerate queries such as $L_{1}:=1/2\cdot 1_{-\ln{2}}+1/2\cdot 1_{\infty}$ and $L_{2}:=2/3\cdot 1_{0}+1/3\cdot 1_{\infty}$.
+Lemma 4.
+Let $L_{1}$ and $L_{2}$ be non-degenerate PLDs such that $L_{1}\npreceq L_{2}$ and $L_{2}\npreceq L_{1}$, i.e. $\{L_{1},L_{2}\}$ is not totally-ordered. Then, for either $i=1$ or $i=2$, we have
+$L_{i}\oplus\sup\{L_{1},L_{2}\}\succ\sup\{L_{i}\oplus L_{1},L_{i}\oplus L_{2}\}.$
+Proof of Theorem˜8.
+The forward direction now follows directly from Lemma˜3 and Theorem˜7.
+As for the reverse direction, suppose that the subfamily of non-degenerate queries in $\mathcal{L}^{<k}$ is not totally-ordered, i.e.
+there exist $L_{1},L_{2}\in\mathcal{L}^{<k}$ that are not ordered.
+By Lemma˜4, we have that $L_{i}\oplus\sup\{L_{1},L_{2}\}\succ\sup\{L_{i}\oplus L_{1},L_{i}\oplus L_{2}\}$.
+On the other hand, if the commutativity condition of Theorem˜7 were satisfied for $\mathcal{L}^{<2k-3}$, then, letting $L_{i}=L_{i,1}\oplus\dots\oplus L_{i,t}$ ($t<k$), we would get
+$L_{i}\oplus\sup\{L_{1},L_{2}\}$ $=L_{i,1}\oplus\dots\oplus L_{i,t}\oplus\sup\{L_{1},L_{2}\}$
+$=L_{i,1}\oplus\dots\oplus L_{i,t-1}\oplus\sup\{L_{i,t}\oplus L_{1},L_{i,t}\oplus L_{2}\}$
+$=\dots$
+$=\sup\{L_{i,1}\oplus\dots\oplus L_{i,t}\oplus L_{1},L_{i,1}\oplus\dots\oplus L_{i,t}\oplus L_{2}\}$
+$=\sup\{L_{i}\oplus L_{1},L_{i}\oplus L_{2}\},$
+which is a contradiction. ∎
 6 Notable Implications
-5.1 PLD and $f$-DP filters are NOT free in general
-Theorems˜2 and 3 give a necessary and sufficient condition for free PLD filters.
-We now switch our view to $f$-DP using Propositions˜10 and 11, for which free filters correspond to Algorithm˜2 in Appendix, to illustrate these results and their implications.
-The condition in Theorem˜2, which is necessary and sufficient for free privacy filters, translates as:
-$\forall f\in\mathcal{F},\ \text{and}\ \mathcal{F}^{\prime}\subseteq\mathcal{F}^{\infty},f\otimes\operatorname{conv}(\mathcal{F}^{\prime})=\operatorname{conv}(f\otimes\mathcal{F}^{\prime}),$ (1)
-where $\otimes$ is $f$-DP composition, $f\otimes\mathcal{F}^{\prime}=\{f\otimes f^{\prime}:f^{\prime}\in\mathcal{F}^{\prime}\}$, and $\operatorname{conv}$ is the lower convex envelope of a set of tradeoff curves.
-All examples also apply to PLD by Proposition˜10, but we visualize them using trade-off curves.
-For examples and visualizations, we work with piece-wise linear tradeoff curves, which can be composed exactly using the following result.
+6.1 Natural Filters Are NOT Free in General
+In this section, we demonstrate examples of queries and budgets that do and do not pass the characterizing conditions of the previous section.
+For visualization purposes, it will be convenient to pass to tradeoff curves.
+Recall that analogous operation for the PLD supremum is the lower convex envelope (c.f.
+Proposition˜11).
+Recall also that the order of domination for PLDs is reversed for their tradeoff curves (see Proposition˜10).
+In this section, we will generally discuss piecewise linear tradeoff curves, which can be composed exactly as follows.
+A proof is given in Appendix˜B.
 Proposition 12 (Exact composition of piecewise-linear tradeoff curves).
-Consider two piecewise-linear tradeoff curves $f_{1},f_{2}$.
-Denote $\alpha_{i,n_{i}}\in[0,1]$ the lengths of intervals over which $f_{i}$ is linear, and $f^{\prime}_{i,n_{i}}\leq 0$ the slope of $f_{i}$ on those intervals.
-Denote $\delta_{i}\triangleq 1-f_{i}(0)$.
-That is, $\sum_{j=1}^{n_{i}}\alpha_{i,j}=1$, and, noting $m_{i}(\alpha)=\arg\max_{n:\in\{1,n_{i}\}}\sum_{j=1}^{n}\alpha_{i,j}\leq\alpha$, we have:
-$f_{i}(\alpha)=1-\delta_{i}+\sum_{j=1}^{m_{i}(\alpha)}f^{\prime}_{i,j}\alpha_{i,j}+\big(\alpha-\sum_{j=1}^{m_{i}(\alpha)}\alpha_{i,j}\big)f^{\prime}_{i,j+1}.$
-Let $k\in[1,n_{1}n_{2}]$ index all combinations of $(j_{1,k},j_{2,k})\in[1,n_{1}]\times[1,n_{2}]$ in order of increasing values of $f^{\prime}_{1,j_{1,k}}f^{\prime}_{2,j_{2,k}}$.
-That is, $k$ indexes all products of a slope from $f_{1}$ and a slope for $f_{2}$ (which are now positive) sorted in decreasing order of magnitude.
-With ne minus sign, the first index is thus the steepest negative slope.
-Then the composition $f\triangleq f_{1}\otimes f_{2}$ is also a piecewise linear tradeoff curve, with $\delta=\delta_{1}+\delta_{2}-\delta_{1}\delta_{2}$, and $\alpha_{k}=\alpha_{1,k}\alpha_{2,k}$ and slopes $f^{\prime}_{k}=-f^{\prime}_{1,j_{1,k}}f^{\prime}_{2,j_{2,k}}$.
-That is, denoting $m(\alpha)=\arg\max_{k:\in\{1,n_{1}n_{2}\}}\sum_{k=1}^{n}\alpha_{k}\leq\alpha$, $f=f_{1}\otimes f_{2}$ is such that:
-$f(\alpha)=1-\delta_{1}-\delta_{2}+\delta_{1}\delta_{2}+\sum_{j=1}^{m(\alpha)}f^{\prime}_{k}\alpha_{k}+\big(\alpha-\sum_{j=1}^{m(\alpha)}\alpha_{k}\big)f^{\prime}_{k+1}.$
+Consider two piecewise linear tradeoff curves $f_{1}$ and $f_{2}$ corresponding to PLDs $L_{1}$ and $L_{2}$.
+Let $(w_{i,j})_{j=1}^{n_{i}}$ be the lengths of the intervals over which $f_{i}$ is linear, let $(f^{\prime}_{i,j})_{j=1}^{n_{i}}$ be the slope of $f_{i}$ over those intervals, and let $\delta_{i}:=1-f_{i}(0)$.
+Now, let $(j_{k},\ell_{k})_{k=1}^{n_{1}n_{2}}$ be an enumeration of $n_{1}]\times[n_{2}]$ such that the product of the slopes $f^{\prime}_{1,j_{k}}\cdot f^{\prime}_{2,\ell_{k}}$ is descending in $k$, i.e.
+$f_{1,j_{1}}\cdot f_{2,\ell_{1}}$ is the product of the steepest slopes of $f_{1}$ and $f_{2}$.
+Then, the composed tradeoff curve $f$ corresponding to $L_{1}\oplus L_{2}$ is also a piecewise linear tradeoff curve over intervals of length $w_{k}:=w_{1,j_{k}}w_{2,\ell_{k}}$ with slopes $f^{\prime}_{k}:=-f^{\prime}_{1,j_{k}}\cdot f^{\prime}_{2,\ell_{k}}$ such that $1-f(0)=\delta_{1}+\delta_{2}-\delta_{1}\delta_{2}$.
+An immediate implication of Theorem˜8 is that the natural filter is not universally free for general queries, even for a single adaptive round.
+This can be seen by noticing that $\mathcal{L}_{\mathrm{approx}}$ is not totally-ordered (recall Definition˜4).
+Figure˜2(a) demonstrates this visually by passing to the tradeoff curves of $\mathcal{L}_{\mathrm{approx}}$.
+Corollary 1.
+The natural filter for $\geq 2$ adaptive $(\varepsilon,\delta)$-DP queries is not universally free, i.e.
+there exists a budget $B\in\mathcal{L}_{\mathrm{all}}$ for which $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L}_{\mathrm{approx}},B}^{2})\npreceq B$.
+(a) Queries $L_{1},L_{2}\in\mathcal{L}_{\mathrm{approx}}$ such that $\{L_{1},L_{2}\}$ is not totally-ordered and a budget $B\in\mathcal{L}_{\mathrm{approx}}$ such that $L_{1}\oplus L_{1},L_{1}\oplus L_{2}\preceq B$, yet $L_{1}\oplus\sup\{L_{1},L_{2}\}\npreceq B$.
+(b) Queries $L_{1},L_{2}\in\mathcal{L}_{\mathrm{pure}}$ such that $\{L_{1},L_{2}\oplus L_{2}\}$ is not totally-ordered.
+Figure 2: Counterexamples showing that natural filters fail for $(\varepsilon,\delta)$-DP queries and $(\varepsilon,\delta)$-DP budgets as well as $\varepsilon$-DP queries and general budgets.
+Another interesting case is that of the $\varepsilon$-DP queries $\mathcal{L}_{\mathrm{pure}}$.
+Indeed, this family is totally-ordered and known to admit privacy filters under basic composition (Rogers et al., 2016).
+However, this family of queries is not closed under composition and $\mathcal{L}_{\mathrm{pure}}^{<3}$ is not totally-ordered anymore.
+Consequently, Theorem˜8 yields the following result, illustrated visually in Figure˜2(b).
+Corollary 2.
+The natural filter for queries $\mathcal{L}_{\mathrm{pure}}$ is universally free for $k=2$ queries but not $k\geq 3$ queries, i.e.
+$\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L}_{\mathrm{pure}},B}^{2})\preceq B$ for all $B\in\mathcal{L}_{\mathrm{all}}$ whereas $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L}_{\mathrm{pure}},B}^{3})\npreceq B$ for some $B\in\mathcal{L}_{\mathrm{all}}$.
+We have shown that natural filters are not universally free, even when restricted to $(\varepsilon,\delta)$-queries.
+These results are surprising, as all previous filter results were that filters are free (RDP/zCDP Feldman and Zrnic (2021); Lécuyer (2021), GDP Koskela et al.
+(2022); Smith and Thakurta (2022)) or asymptotically free ($(\varepsilon,\delta)$-DP Rogers et al.
+(2016); Whitehouse et al.
+(2023)).
+One might hope that restricting the budget would be sufficient to ensure free natural filters.
+Two natural candidates come to mind.
+First, restricting $B\in\mathcal{L}$—a practical condition of choosing privacy bounds from the legal query set itself.
+Second, we could restrict $B$ to a GDP bound $G_{\mu}$ or an approximate GDP bound $R_{0,\delta}\oplus G_{\mu}$ to support approximate $(\varepsilon,\delta)$-DP queries.
+These are good candidates, as the only known free natural filters Koskela et al.
+(2022); Smith and Thakurta (2022) are for $\mathcal{L}_{\mathrm{gdp}}$ queries and budgets.
+Unfortunately, neither restriction is sufficient.
+The first point, that restricting the budget to $\mathcal{L}$ does not ensure the freeness of the natural filter, can be seen by applying Theorem˜6.
+Figure˜2(a) shows an example of queries $L_{1},L_{2}\in\mathcal{L}_{\mathrm{approx}}$ and a budget $B\in\mathcal{L}_{\mathrm{approx}}$ such that $L_{1}\oplus L_{1},L_{1}\oplus L_{2}\preceq B$ yet $L_{1}\oplus\sup\{L_{1},L_{2}\}\npreceq B$.
+In this case, Theorem˜6 implies $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L}_{\mathrm{approx}},B}^{2})\npreceq B$.
+Corollary 3.
+The natural filter is not free for $(\varepsilon,\delta)$-DP queries, even for a $(\varepsilon,\delta)$-DP budget, i.e.
+for some choice of $B\in\mathcal{L}_{\mathrm{approx}}$, we have that $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L}_{\mathrm{approx}},B}^{2})\npreceq B$.
+(a) An example showing queries $L,L_{1},L_{2}\in\mathcal{L}_{\mathrm{all}}$ and a budget $G_{\mu}\in\mathcal{L}_{\mathrm{gdp}}$ such that $L\oplus L_{1},L\oplus L_{2}\preceq G_{\mu}$, yet $L\oplus\sup\{L_{1},L_{2}\}\npreceq G_{\mu}$.
+(b) An example showing queries $L_{1},L_{2}\in\mathcal{L}_{\mathrm{approx}}$ and a budget $(\mu,\delta)$ such that $L_{1}\oplus L_{1},L_{1}\oplus L_{2}\preceq G_{\mu}\oplus R_{0,\delta}$, yet $L_{1}\oplus\sup\{L_{1},L_{2}\}\npreceq G_{\mu}\oplus R_{0,\delta}$.
+Figure 3: Failure of natural filters for GDP and approximate GDP budgets.
+Similarly, restricting the budget to a well-behaved class does not generally lead to free filters either.
+That GDP budgets do not suffice for general queries can again be shown by appealing to Theorem˜6 in the same manner, shown visually in Figure˜3(a).
+Likewise, approximate GDP budgets do not give free natural filters for $(\varepsilon,\delta)$-DP queries, as illustrated in Figure˜3(b).
+Corollary 4.
+The natural filter is not free for general queries, even for a GDP budget, i.e. for some $\mu\geq 0$, $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L}_{\mathrm{all}},G_{\mu}}^{2})\npreceq G_{\mu}$.
+Furthermore, the natural filter is not free for $(\varepsilon,\delta)$-DP queries, even if the budget is chosen to be approximate GDP, i.e.
+for some $\mu\geq 0$ and $\delta\in[0,1]$, $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L}_{\mathrm{approx}},G_{\mu}\oplus R_{0,\delta}}^{2})\npreceq G_{\mu}\oplus R_{0,\delta}$.
+6.2 Families of Queries With Universally Free Natural Filters
+Theorem˜8 ensures that query families that are closed under composition and totally-ordered admit universally free natural filters.
+In particular, since $G_{\mu_{1}}\oplus G_{\mu_{2}}=G_{\sqrt{\mu_{1}^{2}+\mu_{2}^{2}}}$ and since $G_{\mu_{1}}\preceq G_{\mu_{2}}$ whenever $0\leq\mu_{1}\leq\mu_{2}$, we have:
+Corollary 5.
+The natural filter for GDP queries is universally free, i.e. $\operatorname{PLD}(\operatorname{Nat}_{\mathcal{L}_{\mathrm{gdp}},B})\preceq B$ for all $B\in\mathcal{L}_{\mathrm{all}}$.
+Likewise, as $R_{0,\delta_{1}}\oplus R_{0,\delta_{2}}=R_{0,\delta_{1}+\delta_{2}-\delta_{1}\delta_{2}}$ and as $R_{0,\delta_{1}}\preceq R_{0,\delta_{2}}$ for $0\leq\delta_{1}\leq\delta_{2}\leq 1$, we have:
+Corollary 6.
+The natural filter for $(0,\delta)$-DP queries is universally free, i.e.
+for every $B\in\mathcal{L}_{\mathrm{all}}$ we have $\operatorname{PLD}(\operatorname{Nat}_{\{R_{0,\delta:\delta\in[0,1]}\},B})\preceq B$.
+7 Upper Bound for the Natural $(\varepsilon,\delta)$-DP PLD Filter
+We consider the natural $(\varepsilon,\delta)$-DP PLD filter, which is simply an instantiation of Algorithm˜2 with the $(\varepsilon,\delta)$-DP budget $B=R_{\varepsilon,\delta}$ where, recalling Definition˜4,
+$R_{\varepsilon,\delta}=\begin{cases}\infty&\text{w.p. }\delta\\ \varepsilon&\text{w.p. }\frac{1-\delta}{e^{-\varepsilon}+1}\\ -\varepsilon&\text{w.p. }\frac{1-\delta}{e^{\varepsilon}+1}\end{cases}.$
+It is well known that a PLD $L$ is dominated by $R_{\varepsilon,\delta}$ iff $h_{L}(e^{\varepsilon})\leq\delta$.
+In other words, this filter uses the “natural” condition that checks whether $L_{1}\oplus\cdots\oplus L_{t}$ satisfies $(\varepsilon,\delta)$-DP.
+Our main result is that, while the above natural $(\varepsilon,\delta)$-DP PLD filter can fail, it cannot fail too badly; in the sense that we still get a DP guarantee with a blow-up in parameters that is only polylogarithmic in $k,1/\delta$, as stated below [^2] Recall the notation $\operatorname{Nat}_{\mathcal{L},B}^{k}$ from Section 5..
+We remark that our upper bound holds for the family $\mathcal{L}_{\mathrm{sym}}$ of all symmetric PLDs, which is a standard practical setting (see Appendix˜C).
+Theorem 9.
+Let $\varepsilon>0,\delta\in(0,1/4)$.
+Then, the natural $(\varepsilon,\delta)$-DP PLD filter for symmetric queries satisfies $\left({\chi\cdot(\varepsilon+1),\chi\cdot\delta}\right)$}-DP; in other words,
+$\operatorname{PLD}{\left({\operatorname{Nat}_{\mathcal{L}_{\mathrm{sym}},R_{\varepsilon,\delta}}^{k}}\right)}\preceq R_{\chi\cdot(\varepsilon+1),\chi\cdot\delta},$
+where
+$\chi=O{\left({{\left({\log k+\log{\left({1+\frac{1}{\varepsilon}}\right)}}\right)}\cdot\log(1/\delta)}\right)}.$
+When $\varepsilon=\Omega(1)$, the above guarantee is simply $(\chi^{\prime}\cdot\varepsilon,\chi^{\prime}\cdot\delta)$-DP where $\chi^{\prime}=O(\log k\cdot\log(1/\delta))$.
+Remark 10.
+In our current proof, the parameters are quite flexible and we can tradeoff the blow-up factors of the two parameters $\varepsilon,\delta$.
+For example, by changing $\varepsilon_{i}$ in Section˜7.1 to $\frac{2\varepsilon}{2^{n-1-i}}$., we can prove a guarantee of $\left(\chi\cdot\varepsilon,\chi\cdot\frac{\delta}{\varepsilon}\right)$-DP for $\chi=O(\log k\cdot\log(1/\delta))$.
+However, we choose to keep only a single setting of parameters in Theorem˜9 for simplicity.
+For the same reason, we do not attempt to optimize $\chi$.
+In particular, it is plausible that our techniques can achieve $\chi=\tilde{O}(\log k+\log(1/\delta))$ with a more refined argument.
+Nevertheless, our proof cannot eliminate the dependency of $k$ or $\delta$ in $\chi$, and this remains an interesting open question.
+Proof Overview.
+We start with an informal overview of our proof, which consists of three steps:
+  • Privacy Loss Discretization.
+Before we analyze the privacy filter, we start by providing a discretization argument for PLDs.
+Namely, we pick discretization points $\varepsilon_{0},\dots,\varepsilon_{n}$.
+Then, we “round” the PLD $L$ up or down to be these values (or their negations).
+We show that this process results in $L$ being “sandwiched” between two instances of “generalized randomized response”, which has the following form: Randomly select index $i$ from a certain probability distribution and then output $i$ together with the $\varepsilon_{i}$-DP randomized response.
+We show that $L$ is dominated by the “rounded up” (aka “pessimistic”) version of such randomized response while it dominates the “rounded down” (aka “optimistic”) version.
+We remark that our discretization argument is different than previous work (e.g.
+Meiser and Mohammadi (2018); Koskela et al.
+(2020); Doroshenko et al.
+(2022); Gopi et al.
+(2021); Ghazi et al.
+(2022)) which do not produce a generalized randomized response instance (especially for the pessimistic version).
+Indeed, we exploit the symmetry of $L$ to achieve our discretization.
+  • Translation to Additive Filter.
+We consider the following “additive” filter, in which there is a non-negative real value budget associated with each $\varepsilon_{i}$ and each time we simply subtract from the budget; the filter rejects if we run out of budget for any $\varepsilon_{i}$.
+With the above discretization, we can show that the $(\varepsilon,\delta)$-PLD filter is “stronger” than such an additive filter (where the budget spent in each step is from the discretization).
+In other words, if one passes the PLD filter, then one also passes the additive filter.
+This means that, for the purpose of privacy guarantee, we may analyze the latter.
+  • Analysis of Additive Filter via Post-Processing.
+In the final step, we provide the privacy analysis of the additive filter.
+Roughly speaking, we show that the adversary’s view of the additive filter can be shown to be a post-processing of a certain pre-specified number of randomized response (“with high probability”).
+Due to the post-processing property of DP, we can thus conclude that this is as private as composition of randomized response, which we then show to be $(\varepsilon^{\prime},\delta^{\prime})$-DP.
+The above overview skips multiple subtle details, and the remainder of this section is devoted to carry out the above steps in full.
+7.1 Discretization of PLD
+Our discretization is based on a variant of randomized response where with a different prescribed probability, we run a randomized response with a different value of $\varepsilon$.
+We formalize this below.
+Definition 11 (Randomized Response).
+We use $\mathrm{RR}_{\varepsilon}$ to denote the $\varepsilon$-DP binary randomized response. That is, for $x\in\{0,1\}$,
+$\mathrm{RR}_{\varepsilon}(x)=\begin{cases}x&\text{ with probability }\frac{1}{e^{-\varepsilon}+1},\\ 1-x&\text{ with probability }\frac{1}{e^{\varepsilon}+1}.\end{cases}$
+Note that for $\varepsilon=\infty$, $\mathrm{RR}_{\varepsilon}(x)=x$ with probability one.
+Let $\Delta_{n-1}:=\{\mathbf{p}=(p_{1},\dots,p_{n})\in\mathbb{R}_{\geq 0}\mid p_{1}+\cdots+p_{n}=1\}$ be the (n1)(n-1)-simplex, and let $\mathcal{E}_{n}:=\{\bm{\varepsilon}=(\varepsilon_{0},\dots,\varepsilon_{n})\in\mathbb{R}_{\geq 0}^{n+1}\mid 0=\varepsilon_{0}\leq\varepsilon_{1}\leq\cdots\leq\varepsilon_{n}=\infty\}$.
+Definition 12 (Generalized Randomized Response).
+For any $\mathbf{p}\in\Delta_{n-1}$ and $\bm{\varepsilon}\in\mathcal{E}_{n}$, we write $\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}$ to denote the following mechanism (where $x$ denotes the input):
+  • Sample index $i\in[n]$ according to $\mathbf{p}$,
+  • Sample $y\sim\mathrm{RR}_{\varepsilon_{i}}(x)$,
+  • Output $(i,y)$.
+We also define the following “rounded down” version $\mathrm{RR}^{\downarrow}_{\mathbf{p},\bm{\varepsilon}}$ of the above mechanism as follows where we use $\varepsilon_{0}$ to denote 0 throughout:
+  • Sample index $i\in[n]$ according to $\mathbf{p}$,
+  • Sample $y\sim\mathrm{RR}_{\varepsilon_{i-1}}(x)$,
+  • Output $(i,y)$.
+Note that $\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}$ and $\mathrm{RR}^{\downarrow}_{\mathbf{p},\bm{\varepsilon}}$ only differ on the second step, where the former invokes $\mathrm{RR}_{\varepsilon_{i}}$ but the latter invokes $\mathrm{RR}_{\varepsilon_{i-1}}$.
+In addition to Generalized Randomized Response, another notion we will use for the discretization step is the notion of discretized privacy loss profile, which is a vector $\mathbf{p}=(p_{1},\dots,p_{n})$ where $p_{i}$ is the probability that the privacy loss falls into $\varepsilon_{i-1},\varepsilon_{i})$ or $(-\varepsilon_{i},\varepsilon_{i}]$.
+This is formalized below.
+Definition 13 (Discretized Privacy Loss Profile).
+For any PLD $L$, the discretized privacy loss profile of $L$ and $\bm{\varepsilon}\in\mathcal{E}_{n}$, denoted by $\mathrm{disc}_{\bm{\varepsilon}}(L)$, is the vector $\mathbf{p}=(p_{1},\dots,p_{n})\in\Delta_{n-1}$ where
+$p_{i}=\begin{cases}\Pr_{Z\sim L}\left[\varepsilon_{i-1}\leq|Z|<\varepsilon_{i}\right]&\text{ if }i\neq n,\\ \Pr_{Z\sim L}\left[\varepsilon_{n-1}\leq|Z|\right]&\text{ if }i=n.\end{cases}$
+The “optimistic” and “pessimistic” discretization of $L$ is exactly based on $\mathrm{RR}^{\downarrow}_{\mathbf{p},\bm{\varepsilon}}$ and $\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}$, and we have the following theorems which gives the aforementioned “sandwich” around the pair $L$.
+Theorem 11 (Pessimistic Estimate).
+Suppose that $L$ is a symmetric PLD, and let $\mathbf{p}$ be its discretized privacy loss profile.
+Then, $\operatorname{PLD}(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0),\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(1))\succeq L$.
+Theorem 12 (Optimistic Estimate).
+Suppose that $L$ is a symmetric PLD, and let $\mathbf{p}$ be its discretized privacy loss profile.
+Then, $\operatorname{PLD}(\mathrm{RR}^{\downarrow}_{\mathbf{p},\bm{\varepsilon}}(0),\mathrm{RR}^{\downarrow}_{\mathbf{p},\bm{\varepsilon}}(1))\preceq L$.
+The remainder of this subsection are devoted to the proofs of these two theorems.
+7.1.1 Useful Lemmata
+Before we can prove Theorems˜11 and 12, we will first prove a couple of technical lemmata that would be useful.
+First is the symmetry of $\mathrm{RR}$ and $\mathrm{RR}^{\downarrow}$, which are simple to prove.
+Lemma 5.
+For any $\mathbf{p},\bm{\varepsilon}$, the pairs $(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0),\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(1))$ and $(\mathrm{RR}^{\downarrow}_{\mathbf{p},\bm{\varepsilon}}(0),\mathrm{RR}^{\downarrow}_{\mathbf{p},\bm{\varepsilon}}(1))$ are symmetric.
 Proof.
-See Section˜A.1. ∎
-A first implication Theorems˜2 and 3 is that there are no free $f$-DP filters in general:
-Corollary 1 ($f$-DP filters are NOT free).
-Consider $\mathcal{T}$ the family of all tradeoff curves, and $\mathcal{F}=\{f_{\varepsilon,\delta},\ \varepsilon\geq 0,\ \delta\in[0,1]\}$ the set of tradeoff curves for all $(\varepsilon,\delta)$-DP mechanisms.
-Neither $\mathcal{T}$ nor $\mathcal{F}$ have free $f$-DP filters. Precisely, Algorithm˜2 is not $f$-DP.
+Consider the Markov kernel $\kappa$ that takes in $(i,y)$ and outputs (i,1y)(i,1-y).
+It is obvious to see that $\kappa(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0))=\kappa(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(1))$ and $\kappa(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(1))=\kappa(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0))$.
+Thus, $(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0),\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(1))$ dominates $(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(1),\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0))$ and vice versa.
+In other words, $(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0),\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(1))$ is symmetric.
+The proof for $(\mathrm{RR}^{\downarrow}_{\mathbf{p},\bm{\varepsilon}}(0),\mathrm{RR}^{\downarrow}_{\mathbf{p},\bm{\varepsilon}}(1))$ is analogous. ∎
+For the proofs below, it will be more convenient to deal with pairs rather than privacy loss. For this, we give an alternate definition of discretized privacy loss profile below.
+Definition 14.
+Let $(P,Q)$ be any pair of distributions. For any $0\leq\varepsilon<\varepsilon^{\prime}$, let
+$I_{\varepsilon,\varepsilon^{\prime}}:=\begin{cases}[\varepsilon,\varepsilon^{\prime})&\text{ if }\varepsilon^{\prime}\neq\infty,\\ [\varepsilon,\varepsilon^{\prime})\cup\{\infty\}&\text{ if }\varepsilon^{\prime}=\infty.\end{cases}$
+Moreover, let
+$\Omega^{P,Q}_{\varepsilon,\varepsilon^{\prime}}=\left\{\omega~\middle|~\log\left(\frac{dP}{dQ}(\omega)\right)\in I_{\varepsilon,\varepsilon^{\prime}}\right\},$
+and similarly
+$\Omega^{Q,P}_{\varepsilon,\varepsilon^{\prime}}=\left\{\omega~\middle|~\log\left(\frac{dQ}{dP}(\omega)\right)\in I_{\varepsilon,\varepsilon^{\prime}}\right\}.$
+The following observation follows directly from the definition of PLD.
+Observation 13.
+For any $L=\operatorname{PLD}(P,Q)$ and $\bm{\varepsilon}\in\mathcal{E}_{n}$, the discretized privacy loss profile of $L$ is equal to $\mathbf{p}=(p_{1},\dots,p_{n})$ where
+$p_{i}=P{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}\cup\Omega^{Q,P}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}.$
+The next lemma is a type of symmetry with regards to $\Omega^{P,Q}_{\varepsilon,\varepsilon^{\prime}}$ and $\Omega^{P,Q}_{\varepsilon^{\prime},\varepsilon}$:
+Lemma 6.
+For any symmetric pair $(P,Q)$ and $0\leq\varepsilon<\varepsilon^{\prime}$,
+$P{\left({\Omega^{P,Q}_{\varepsilon,\varepsilon^{\prime}}}\right)}=Q{\left({\Omega^{Q,P}_{\varepsilon,\varepsilon^{\prime}}}\right)}$ and $P{\left({\Omega^{Q,P}_{\varepsilon,\varepsilon^{\prime}}}\right)}=Q{\left({\Omega^{P,Q}_{\varepsilon,\varepsilon^{\prime}}}\right)}.$
 Proof.
-$\mathcal{F}\subset\mathcal{T}$, and $\mathcal{F}$ has crossing tradeoff functions, which are thus not well ordered.
-By Theorem˜3, Algorithm˜1 is not PLD for budget $B$, and so by Propositions˜10 and 11 Algorithm˜1 is not $f$-DP.
+We only prove the first equality as the second is analogous.
+Let $F_{P,Q}$ and $F_{Q,P}$ denote the CDFs of $\operatorname{PLD}(P,Q)$ and $\operatorname{PLD}(Q,P)$, respectively.
+Since the hockey-stick curves $H_{x}(P\parallel Q)$ and $H_{x}(Q\parallel P)$ are equal, (Zhu et al., 2022, Lemma 23) implies that $F_{P,Q}$ and $F_{Q,P}$ are exactly the same.
+Thus, for $\varepsilon^{\prime}\neq\infty$ we have
+$P{\left({\Omega^{P,Q}_{\varepsilon,\varepsilon^{\prime}}}\right)}=\lim_{\tilde{\varepsilon}\to(\varepsilon)^{-},\tilde{\varepsilon}^{\prime}\to(\varepsilon^{\prime})^{-}}F_{P,Q}(\tilde{\varepsilon}^{\prime})-F_{P,Q}(\tilde{\varepsilon})=\lim_{\tilde{\varepsilon}\to(\varepsilon)^{-},\tilde{\varepsilon}^{\prime}\to(\varepsilon^{\prime})^{-}}F_{Q,P}(\tilde{\varepsilon}^{\prime})-F_{Q,P}(\tilde{\varepsilon})=Q{\left({\Omega^{Q,P}_{\varepsilon,\varepsilon^{\prime}}}\right)}.$
+For $\varepsilon^{\prime}=\infty$, we similarly have
+$P{\left({\Omega^{P,Q}_{\varepsilon,\varepsilon^{\prime}}}\right)}=\lim_{\tilde{\varepsilon}\to(\varepsilon)^{-}}1-F_{P,Q}(\tilde{\varepsilon})=\lim_{\tilde{\varepsilon}\to(\varepsilon)^{-}}1-F_{Q,P}(\tilde{\varepsilon})=Q{\left({\Omega^{Q,P}_{\varepsilon,\varepsilon^{\prime}}}\right)}.$
+7.1.2 Pessimistic Estimate Proof
+We are now ready to prove Theorem˜11.
+Proof of Theorem˜11.
+Suppose that $L=\operatorname{PLD}(P,Q)$.
+Since both $(P,Q)$ and $(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0),\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(1))$ are symmetric (Lemma˜5), by Lemma˜8, it suffices to show that $H_{x}(P\parallel Q)\leq H_{x}(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0)\parallel\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(1))$ for any $x\geq 1$.
+For $i\in[n]$, let $P_{i}$ (resp.
+$Q_{i}$) denote the distribution of $P$ (resp.
+$Q$) conditioned on the value being in $\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}\cup\Omega^{Q,P}_{\varepsilon_{i-1},\varepsilon_{i}}$.
+From Lemma˜6, we have that $P$ and $Q$ can be written as mixture distributions $\sum_{i\in[n]}p_{i}\cdot P_{i}$ and $\sum_{i\in[n]}p_{i}\cdot Q_{i}$, respectively.
+Since $P_{i}$’s support are disjoint and $Q_{i}$’s support are disjoint (and the supports of $P_{i},Q_{j}$ are disjoint for all $i\neq j$), we simply have
+$H_{x}(P\parallel Q)=\sum_{i\in[n]}p_{i}\cdot H_{x}(P_{i}\parallel Q_{i}).$
+Similarly, we have
+$H_{x}(\mathrm{RR}_{\mathbf{p},\bm{\varepsilon}}(0)\parallel\mathrm{RR}_{\mathbf{p},\varepsilon}(1))=\sum_{i\in[n]}p_{i}\cdot H_{x}(\mathrm{RR}_{\varepsilon_{i}}(0)\parallel\mathrm{RR}_{\varepsilon_{i}}(1)).$
+Thus, it suffices to show that [^3] This essentially follows from Kairouz et al.
+(2015) but we repeat the full proof here for completeness.
+$H_{x}(P_{i}\parallel Q_{i})\leq H_{x}(\mathrm{RR}_{\varepsilon_{i}}(0)\parallel\mathrm{RR}_{\varepsilon_{i}}(1))$ for all $i\in[n]$ and $x\geq 1$.
+To see this is true, consider three cases:
+  • Case I: $x\geq e^{\varepsilon_{i}}$. Then, $H_{x}(P_{i}\parallel Q_{i})=0=H_{x}(\mathrm{RR}_{\varepsilon_{i}}(0)\parallel\mathrm{RR}_{\varepsilon_{i}}(1))$
+  • Case II: $1\leq x\leq e^{\varepsilon_{i-1}}$.
+In this case, we have $H_{x}(\mathrm{RR}_{\varepsilon_{i}}(0)\parallel\mathrm{RR}_{\varepsilon_{i}}(1))=\frac{1-x\cdot e^{-\varepsilon_{i}}}{1+e^{-\varepsilon_{i}}}.$ Meanwhile, we have
+    $H_{x}(P_{i}\parallel Q_{i})$ $=P_{i}{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}-x\cdot Q_{i}{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}$
+    $=\frac{1}{p_{i}}\cdot\left(P{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}-x\cdot Q{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}\right)$
+    $=\frac{P{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}-x\cdot Q{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}}{P{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}+P{\left({\Omega^{Q,P}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}}$
+[…]
+    $=\frac{x+1}{1+\frac{Q{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}}{P{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}}}-x$
+    $\leq\frac{x+1}{1+e^{-\varepsilon_{i}}}-x$
+    $=\frac{1-x\cdot e^{-\varepsilon_{i}}}{1+e^{-\varepsilon_{i}}},$
+    where the inequality is due to the definition of $\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}$.
+  • Case III: $x\in{\left({e^{\varepsilon_{i-1}},e^{\varepsilon_{i}}}\right)}$. Since $H_{x}$ is convex in $x$, we have
+    $H_{x}(P_{i}\parallel Q_{i})$ $\leq\frac{x-e^{\varepsilon_{i-1}}}{e^{\varepsilon_{i}}-e^{\varepsilon_{i-1}}}\cdot H_{e^{\varepsilon_{i}}}(P_{i}\parallel Q_{i})+\frac{e^{\varepsilon_{i}}-x}{e^{\varepsilon_{i}}-e^{\varepsilon_{i-1}}}\cdot H_{e^{\varepsilon_{i-1}}}(P_{i}\parallel Q_{i})$
+    $\leq\frac{x-e^{\varepsilon_{i-1}}}{e^{\varepsilon_{i}}-e^{\varepsilon_{i-1}}}\cdot H_{e^{\varepsilon_{i}}}(\mathrm{RR}_{\varepsilon_{i}}(0)\parallel\mathrm{RR}_{\varepsilon_{i}}(1))+\frac{e^{\varepsilon_{i}}-x}{e^{\varepsilon_{i}}-e^{\varepsilon_{i-1}}}\cdot H_{e^{\varepsilon_{i-1}}}(\mathrm{RR}_{\varepsilon_{i}}(0)\parallel\mathrm{RR}_{\varepsilon_{i}}(1))$
+    $=H_{\alpha}(\mathrm{RR}_{\varepsilon_{i}}(0)\parallel\mathrm{RR}_{\varepsilon_{i}}(1)),$
+    where the second inequality follows from Cases I and II above, and the last equality is due to linearity of $H_{x}(\mathrm{RR}_{\varepsilon_{i}}(0)\parallel\mathrm{RR}_{\varepsilon_{i}}(1))$ for $x\in I_{e^{\varepsilon_{i-1}},e^{\varepsilon_{i}}}$.
 ∎
-Figure˜1 directly illustrates a counter-example to Theorem˜2 (Equation˜1) visually.
-We show that there exists tradeoff curves in $\mathcal{F}$ (approximate DP mechanisms) such that $\alpha\ \textrm{s.t.}\ g_{1}\otimes\operatorname{conv}(g_{1},g_{2})(\alpha)<\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})(\alpha)$.
-That is, the exists $f=g1$ and $\mathcal{H}=\{g_{1},g_{2}\}$ such that $f\otimes\operatorname{conv}(\mathcal{H})\neq\operatorname{conv}(f\otimes\mathcal{H})$, violating (Equation˜1).
-By Propositions˜10 and 11, there exists $L$ and $\mathcal{L}^{\prime}$ such that $L\oplus\sup{\mathcal{L}^{\prime}}\neq\sup\{L\oplus L^{\prime}:L^{\prime}\in\mathcal{L}^{\prime}\}$, violating Theorem˜2.
-In fact, Figure˜1 illustrates the two cases to account for, to directly prove a special case of Theorem˜3 for $(\varepsilon,\delta)$-DP:
-Proposition 13 ($(\varepsilon,\delta)$-DP curves that cross always violate Theorem˜2 (Equation˜1)).
-Consider $f\in\mathcal{T}$, and any set of tradeoff curves $\mathcal{F}$ such that $f_{\varepsilon_{1},\delta_{1}}$ and $f_{\varepsilon_{2},\delta_{2}}$ cross, and $f_{\varepsilon_{1},\delta_{1}},f_{\varepsilon_{2},\delta_{2}}\in\mathcal{F}$.
-$\mathcal{F}$ does not have free $f$-DP filters. Precisely, Algorithm˜2 is not $f$-DP.
+7.1.3 Optimistic Estimate Proof
+We next prove Theorem˜12. This proof is in fact simpler than the previous one and only two cases are required.
+Proof of Theorem˜12.
+Similar to the proof of Theorem˜11, it suffices to show that
+$H_{x}(P_{i}\parallel Q_{i})\geq H_{x}(\mathrm{RR}_{\varepsilon_{i-1}}(0)\parallel\mathrm{RR}_{\varepsilon_{i-1}}(1))$
+for all $i\in[n]$ and $x\geq 1$. To see this is true, consider two cases:
+  • Case I: $x\geq e^{\varepsilon_{i-1}}$. In this case, we simply have $H_{x}(\mathrm{RR}_{\varepsilon_{i-1}}(0)\parallel\mathrm{RR}_{\varepsilon_{i-1}}(1))=0\leq H_{x}(P_{i}\parallel Q_{i})$.
+  • Case II: $x\in[1,e^{\varepsilon_{i-1}})$.
+In this case, we have $H_{x}(\mathrm{RR}_{\varepsilon_{i-1}}(0)\parallel\mathrm{RR}_{\varepsilon_{i-1}}(1))=\frac{1-x\cdot e^{-\varepsilon_{i-1}}}{1+e^{-\varepsilon_{i-1}}}.$ Meanwhile, similar to Case II in the proof of Theorem˜11, we have
+    $H_{x}(P_{i}\parallel Q_{i})=\frac{x+1}{1+\frac{Q{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}}{P{\left({\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}}\right)}}}-x\geq\frac{x+1}{1+e^{-\varepsilon_{i-1}}}-x=\frac{1-x\cdot e^{-\varepsilon_{i-1}}}{1+e^{-\varepsilon_{i-1}}},$
+    where the inequality is due to the definition of $\Omega^{P,Q}_{\varepsilon_{i-1},\varepsilon_{i}}$. ∎
+7.2 From PLD Filter to Additive Filter
+Having established fundamental results on discretization, we can now state our “additive” privacy filter.
+In this filter, we have a budget $\bm{\ell}=(\ell_{1},\dots,\ell_{n})\in\mathbb{R}^{n}_{\geq 0}$ where informally $\ell_{i}$ corresponds to the budget for $\mathrm{RR}_{\varepsilon_{i}}$.
+When a mechanism comes, subtract its discretized privacy profile $\mathbf{p}$ from the budget.
+If the budget remains non-negative in all coordinates (denoted by [^4] Here we write $\mathbf{0}$ to denote the all-zero vector, and $\mathbf{u}\geq\mathbf{v}$ for vectors $\mathbf{u},\mathbf{v}$ if the inequality holds coordinate-wise.
+$\bm{\ell}\geq\mathbf{0}$), we subtract it from the budget and return the response to the adversary.
+This is formalized in Algorithm˜3.
+Algorithm 3 Additive Discretized PL Profile Filter
+0: Discretization points $\bm{\varepsilon}\in\mathcal{E}_{n}$, budgets $\bm{\ell}\in\mathbb{R}^{n}_{\geq 0}$, analyst $\mathcal{A}$, query capacity $k$, dataset $D$
+for $t\leftarrow 1,\dots,k$ do
+  $\mathcal{A}(Y_{1},\dots,Y_{t-1})$ gives mechanism $M_{t}$ that has PLD $L_{t}\in\mathcal{L}_{\mathrm{sym}}$
+  $\mathbf{p}^{t}\leftarrow$ Discretized privacy loss profile of $L_{t}$ w.r.t. $\bm{\varepsilon}$
+  $\bm{\ell}\leftarrow\bm{\ell}-\mathbf{p}^{t}$
+  if $\bm{\ell}\geq\mathbf{0}$ then
+   yield $Y_{t}\sim M_{t}(D)$
+  else
+   break
+Setting of Parameters.
+For the remainder of this section, we set the parameter $\bm{\varepsilon}$ as follows:
+  • Let $n=1+\left\lceil\log_{2}{\left({\frac{k(\varepsilon+2)}{\varepsilon}}\right)}\right\rceil$.
+  • Let $\varepsilon_{n}=\infty$
+  • For all $i\in[n-1]$, let $\varepsilon_{i}=\frac{\varepsilon+2}{2^{n-1-i}}$.
+Furthermore, let us define $\bm{\ell}^{*}=(\ell^{*}_{1},\dots,\ell^{*}_{n})$ as follows:
+$\ell^{*}_{i}=\begin{cases}k&\text{ if }i=1\\ 8\cdot\max\left\{\ell\in\mathbb{N}~\middle|~H_{e^{\varepsilon}}(\mathrm{RR}_{\varepsilon_{i-1}}^{\otimes(\ell-1)}(0)\parallel\mathrm{RR}_{\varepsilon_{i-1}}^{\otimes(\ell-1)}(1))\leq 2\delta\right\}&\text{ if }1<i<n,\\ 4\delta&\text{ if }i=n\end{cases}$ (1)
+Additive Profile Filter as Adversary.
+Following notation in Sections˜3 and 5, we write the Additive Discretized PL Profile Filter (Algorithm˜3) can be viewed as the rule
+$\operatorname{Discrete}_{\bm{\varepsilon},\bm{\ell}}^{k}(L_{1},\dots,L_{t})=\{L_{t+1}\in\mathcal{L}_{\mathrm{sym}}:\mathrm{disc}_{\bm{\varepsilon}}(L_{1})+\cdots+\mathrm{disc}_{\bm{\varepsilon}}(L_{t+1})\leq\bm{\ell}\}.$
+Our main theorem in this subsection is that the Additive Discretized PL Profile Filter–when initialized with the above discretization points and privacy budgets–is less strict than the natural $(\varepsilon,\delta)$-DP PLD filter:
+Theorem 14.
+[…]
+$\operatorname{Nat}_{\mathcal{L}_{\mathrm{sym}},R_{\varepsilon,\delta}}^{k}(L_{1},\dots,L_{t})\subseteq\operatorname{Discrete}_{\bm{\varepsilon},\bm{\ell}}^{k}(L_{1},\dots,L_{t}).$ (2)
 Proof.
-Direct corollary of Theorem˜3, self standing direct proof in Section˜A.2. ∎
-(a) Case 1: $g_{1}\otimes g_{1}$ and $g_{1}\otimes g_{2}$ cross.
-(b) Case 2: $g_{1}\otimes g_{1}\geq g_{1}\otimes g_{2}$
-Figure 1: Counter examples to Theorem 2 (Equation˜1), showing that $f$-DP filters are NOT free in general (for the set of all tradeoff curves $\mathcal{T}$), or when choosing adaptively among the set of $(\varepsilon,\delta)$-DP mechanisms.
-This is true even if $f\in\mathcal{F}$.
-Another interesting case is that of pure DP, or $\varepsilon$-DP, mechanisms.
-Indeed, this family is well ordered ($\varepsilon_{1}>\varepsilon_{2}\geq 0\Rightarrow f_{\varepsilon_{1}}\preceq f_{\varepsilon_{2}}$, illustrated on Figure˜2(a)), and known to admit privacy filters under basic composition [RRUV16].
-However, this family of DP mechanisms is not closed under $f$-DP or PLD composition, and its self composition is not well ordered anymore, as illustrated on Figure˜2(b).
-As we can see, this is sufficient to violate Theorems˜2 and 3, and hence rule out free $f$-DP/PLD filters for the family of $\varepsilon$-DP under $f$-DP/PLD composition.
-(a) $\varepsilon$-DP admits two-step filters (example)
-(b) $\varepsilon$-DP general filter counter-example
-Figure 2: $\varepsilon$-DP mechanisms are well ordered, but not closed under composition.
-If $\mathcal{F}=\{f_{\varepsilon}:\ \varepsilon\geq 0\}$, then $\mathcal{F}^{2}\subset\mathcal{F}^{\infty}$ is not well ordered (Figure˜2(a)), and by Theorem˜3 pure DP mechanisms do not admit an $f$-DP/PLD filter (Figure˜2(b)).
-We have shown that $f$-DP/PLD filters are not free in general (when $\mathcal{F}$ is the set of all tradeoff curves), or for the set of $(\varepsilon,\delta)$-DP mechanisms.
-These results are surprising, as all previous filter results were that filters are free (RDP/zCDP [feldman2021individual, lecuyer2021practical], GDP [koskelaindividual, smith2022fullyadaptivecompositiongaussian]) or asymptotically free ($(\varepsilon,\delta)$-DP [rogers2016privacy, whitehouse2023fully]).
-One might hope that restricting $f$, the bound on privacy loss used as input to the $f$-DP filter, would be sufficient to ensure free filters.
-Two natural candidates come to mind.
-First, restricting $f\in\mathcal{F}$, a practical condition of choosing privacy bounds in the set of tradeoff curves used in the filter.
-Second, we could restrict $f$ to a GDP curve $G_{\mu}$ (the tradeoff curve of a Gaussian mechanism), or an approximate GDP curve $f_{0,\delta}\otimes G_{\mu}$ to support approximate DP $f_{\varepsilon,\delta}$ tradeoff curves in $\mathcal{F}$.
-These are good candidates, as $\{G_{\mu}:\mu\in\mathbb{R}^{+}\}$ is the only known set of tradeoff curves to have free $f$-DP filters [koskelaindividual, smith2022fullyadaptivecompositiongaussian].
-Unfortunately, neither restriction is sufficient:
-Corollary 2.
-Consider the set of $(\varepsilon,\delta)$-DP tradeoff curves $\mathcal{F}=\{f_{\varepsilon,\delta}:\varepsilon\geq 0,\ \delta\in[0,1]\}$.
-$\mathcal{F}$ does not have free $f$-DP filters even for $f\in\mathcal{F}$.
+Consider any $L_{1},\dots,L_{t+1}\in\mathcal{L}_{\mathrm{sym}}$, and let $\mathbf{p}^{1}=\mathrm{disc}_{\bm{\varepsilon}}(L_{1}),\dots,\mathbf{p}^{t+1}=\mathrm{disc}_{\bm{\varepsilon}}(L_{t+1})$.
+Note that (2) is equivalent to showing the following: if $h_{L_{1}\oplus\cdots\oplus L_{t+1}}({e^{\varepsilon}})\leq\delta$, then we must have $p^{1}_{i}+\cdots+p^{t+1}_{i}\leq\ell^{*}_{i}$ for all $i\in[n]$.
+The case $i=1$ is obvious since $\ell^{*}_{1}=k\geq t$.
+We will henceforth only consider $i>1$.
+We will show the contrapositive: Suppose that $p^{1}_{i}+\cdots+p^{t+1}_{i}>\ell^{*}_{i}$ for some $i\in\{2,\dots,n\}$.
+We will show that $h_{L_{1}\oplus\cdots\oplus L_{t+1}}({e^{\varepsilon}})>\delta$.
+Recall (from Theorem˜12) that $L_{j}$ dominates $(\mathrm{RR}^{\downarrow}_{\mathbf{p}^{j},\varepsilon}(0),\mathrm{RR}^{\downarrow}_{\mathbf{p}^{j},\varepsilon}(1))$.
+Thus, it suffices to show that
+$H_{e^{\varepsilon}}\left(\mathrm{RR}^{\downarrow}_{\mathbf{p}^{1},\varepsilon}(0)\times\cdots\times\mathrm{RR}^{\downarrow}_{\mathbf{p}^{t+1},\varepsilon}(0)~\middle\|~\mathrm{RR}^{\downarrow}_{\mathbf{p}^{1},\varepsilon}(1)\times\cdots\times\mathrm{RR}^{\downarrow}_{\mathbf{p}^{t+1},\varepsilon}(1)\right)>\delta.$ (3)
+For $X\in\{0,1\}$, consider post-processing $(y^{1},\dots,y^{t+1})\sim\mathrm{RR}^{\downarrow}_{\mathbf{p}^{1},\varepsilon}(X)\times\cdots\times\mathrm{RR}^{\downarrow}_{\mathbf{p}^{t+1},\varepsilon}(X)$ as follows, where $\tau\in\mathbb{N}$ will be specified below.
+  • Recall that each $y^{j}=(y^{j}_{1},y^{j}_{2})$ is a tuple of two numbers: $y^{j}_{1}\in[n]$ and $y^{j}_{2}\in\{0,1\}$. Let $j_{1}<\dots<j_{m}$ be indices $j$ such that $y^{j}=i$.
+  • If $m<\tau$, then output $\bot$. Otherwise, output $(y^{j_{1}},\dots,y^{j_{\tau}})$.
+Let $R(X)$ denote the distribution of the output of the above process, and let $q$ denote the probability that the above process does not output $\bot$.
+(Note that $q$ does not depend on $X$.) Observe that, conditioned on not outputting $\bot$, the distribution $R(X)$ is exactly the same as $\mathrm{RR}_{\varepsilon_{i-1}}^{\otimes\tau}(X)$.
+Thus,
+$H_{e^{\varepsilon}}\left(\mathrm{RR}^{\downarrow}_{\mathbf{p}^{1},\varepsilon}(0)\times\cdots\times\mathrm{RR}^{\downarrow}_{\mathbf{p}^{t+1},\varepsilon}(0)~\middle\|~\mathrm{RR}^{\downarrow}_{\mathbf{p}^{1},\varepsilon}(1)\times\cdots\times\mathrm{RR}^{\downarrow}_{\mathbf{p}^{t+1},\varepsilon}(1)\right)$
+$\geq q\cdot H_{e^{\varepsilon}}\left(\mathrm{RR}_{\varepsilon_{i-1}}^{\otimes\tau}(0)~\middle\|~\mathrm{RR}_{\varepsilon_{i-1}}^{\otimes\tau}(1)\right).$ (4)
+We will next consider two cases separately: $i=n$ and $1<i<n$.
+Case I: $i=n$.
+In this case, set $\tau=1$.
+Notice that $H_{e^{\varepsilon}}(\mathrm{RR}_{\varepsilon_{n-1}}^{\otimes\tau}(0)\parallel\mathrm{RR}_{\varepsilon_{n-1}}^{\otimes\tau}(1))=\frac{1-e^{\varepsilon-\varepsilon_{n-1}}}{1+e^{-\varepsilon_{n-1}}}$.
+Due to our choice $\varepsilon_{n-1}=\varepsilon+2$, we have $H_{e^{\varepsilon}}(\mathrm{RR}_{\varepsilon_{n-1}}^{\otimes\tau}(0)\parallel\mathrm{RR}_{\varepsilon_{n-1}}^{\otimes\tau}(1))>0.5$.
+Observe also that
+$q$ $=1-(1-p^{1}_{n})\cdots(1-p^{t+1}_{n})\geq 1-e^{-p^{1}_{n}-\cdots-p^{t+1}_{n}}\geq 1-e^{-4\delta}\geq 2\delta,$
+where the first and last inequalities follow from the well-known [^5] See e.g.
+Topsøe (2007).
+identity $\frac{x}{1+x}\leq\ln(1+x)\leq x$ for all x>1x>-1 (and from $\delta<1/4$).
+Combining the two inequalities and (4), we can conclude that (3) holds (for $i=n$) as desired.
+Case II: $1<i<n$.
+In this case, we set $\tau=\max\left\{\ell\in\mathbb{N}~\middle|~H_{e^{\varepsilon}}\left(\mathrm{RR}_{\varepsilon_{i-1}}^{\otimes(\ell-1)}(0)~\middle\|~\mathrm{RR}_{\varepsilon_{i-1}}^{\otimes(\ell-1)}(1)\right)\leq 2\delta\right\}$.
+By the choice of $\tau$, we have
+$H_{e^{\varepsilon}}(\mathrm{RR}_{\varepsilon_{i-1}}^{\otimes\tau}(0)\parallel\mathrm{RR}_{\varepsilon_{i-1}}^{\otimes\tau}(1))>2\delta.$
+We will next bound $q$.
+Observe that $m$ is distributed exactly as $Z=Z_{1}+\cdots+Z_{t+1}$ where $Z_{1}\sim\mathrm{Ber}(p^{1}_{i}),\dots,Z_{t+1}\sim\mathrm{Ber}(p^{t+1}_{i})$.
+Thus, we have $\mathop{\mathbb{E}}[Z]=p^{1}_{i}+\cdots+p^{t+1}_{i}>\ell^{*}_{i}$ and $\mathrm{Var}[Z]\leq\mathop{\mathbb{E}}[Z]$.
+Note that we set $\tau<\ell^{*}_{i}/2$.
+Thus, we can thus apply Chebyshev’s inequality to conclude that
+$1-q=\Pr[Z<\tau]\leq\Pr[Z<\mathop{\mathbb{E}}[Z]/2]\leq\frac{\mathrm{Var}[Z]}{(\mathop{\mathbb{E}}[Z]/2)^{2}}\leq\frac{4}{\mathop{\mathbb{E}}[Z]}<\frac{4}{\ell^{*}_{i}}<\frac{1}{2}.$
+Combining the two inequalities and (4), we can conclude that (3) holds as desired. ∎
+7.3 Analyzing Additive Discretized PL Profile Filter
+Due to Theorem˜14, it suffices to directly analyze the Additive Discretized PL Profile Filter. In particular, we will show the following:
+Theorem 15.
+For any $\varepsilon>0,\delta\in(0,1/4)$, let $\bm{\varepsilon},\bm{\ell}^{*}$ be as in (1). Then,
+$\operatorname{PLD}{\left({\operatorname{Discrete}_{\bm{\varepsilon},\bm{\ell}^{*}}^{k}}\right)}\preceq R_{\chi\cdot(\varepsilon+1),\chi\cdot\delta},$
+where
+$\chi=O{\left({{\left({\log k+\log{\left({1+\frac{1}{\varepsilon}}\right)}}\right)}\cdot\log(1/\delta)}\right)}.$
+Before we prove Theorem˜15, let us first note that it immediately implies the main theorem of this section (Theorem˜9).
+Proof of Theorem˜9.
+Combining Theorems˜14 and 15 respectively, we have
+$\operatorname{PLD}{\left({\operatorname{Nat}_{\mathcal{L}_{\mathrm{sym}},R_{\varepsilon,\delta}}^{k}}\right)}\preceq\operatorname{PLD}{\left({\operatorname{Discrete}_{\bm{\varepsilon},\bm{\ell}^{*}}^{k}}\right)}\preceq R_{\chi\cdot(\varepsilon+1),\chi\cdot\delta}.$
+[…]
+$b^{*}_{i}=\begin{cases}0&\text{ if }i=n\\ 2\lceil\log(1/\delta)\rceil\cdot\ell^{*}_{i}&\text{ if }i\neq n.\end{cases}$ (5)
+7.3.1 Useful Lemmata
+We will also use the following version of “triangle inequality”.
+Note that the assumption here is stronger than just assumption that the TV distance between $P,P^{\prime}$ and $Q,Q^{\prime}$ are smaller than $\gamma$.
+This stronger assumption allows us to derive an additive bound that is only $\gamma$, rather than a weaker bound $(x+1)\cdot\gamma$ (if we were to only assume a bound on the TV distance).
+This is crucial for our proof of Theorem˜15, which requires a large value of $x$.
+Lemma 7.
+Let $x,\gamma\geq 0$, and let $P,Q$ be distributions on domain $\Omega$.
+Furthermore, let $P^{\prime},Q^{\prime}$ be distributions on domain $\Omega\cup\{\omega^{*}\}$ such that the following holds: $P(E)\geq P^{\prime}(E)$ and $Q(E)\geq Q^{\prime}(E)$ for all $E\subseteq\Omega$, and $P^{\prime}(\omega^{*})\leq\gamma$.
+Then, $H_{x}(P\parallel Q)\leq H_{x}(P^{\prime}\parallel Q^{\prime})+\gamma$.
 Proof.
-Consider two tradeoff curves that cross $g_{1}=f_{\varepsilon_{1},\delta_{1}}\in\mathcal{F}$ and $g_{2}=f_{\varepsilon_{2},\delta_{2}}\mathcal{F}$, that fall in case 1 in the proof of Proposition˜13 (this is possible, but the proof is identical for case 2, swapping the proper slope).
-Pick $f=f_{\varepsilon,\delta}$ with $\varepsilon=\log\big(e^{2\varepsilon_{1}}-(1+e^{\varepsilon_{1}})\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}\big)$ (that is, $e^{-\varepsilon}$ is the first slope of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$) and $\delta=\delta_{1}+\delta_{2}-\delta_{1}\delta_{2}$ (that is, the $\delta$ of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$).
-Consider an adversary playing Algorithm˜2 with a $g_{1}$ mechanism first, and then an adaptive choice between $g_{1},g_{2}$.
-By construction, $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})\geq f$, so $g_{1}\otimes g_{1}\geq f$ and $g_{1}\otimes g_{2}\geq f$: Algorithm˜2 always accepts the query.
-From the proof of Theorem˜2 we know that a two-steps filter can play $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$.
-By construction, we have that $f(\alpha)=\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})(\alpha)$ on $\alpha\in(0,\alpha_{1}^{\star 2}]$ From Figure˜1(a) (and the proof of Proposition˜13) we know that on $\alpha\in(0,\alpha_{1}^{\star 2}]$, we have $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})(\alpha)<\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})(\alpha)=f(\alpha)$.
-Algorithm˜2 is thus not $f$-DP, concluding the proof.
+We have
+$H_{x}(P\parallel Q)$ $=\sup_{E\subseteq\Omega}P(E)-x\cdot Q(E)$
+$=\sup_{E\subseteq\Omega}1-P(\Omega\setminus E)-x\cdot Q(E)$
+$\leq\sup_{E\subseteq\Omega}1-P^{\prime}(\Omega\setminus E)-x\cdot Q^{\prime}(E)$
+$=\sup_{E\subseteq\Omega}P^{\prime}(\omega^{*})+P^{\prime}(E)-x\cdot Q^{\prime}(E)$
+$\leq P^{\prime}(\omega^{*})+H_{x}(P^{\prime}\parallel Q^{\prime})$
+$\leq\gamma+H_{x}(P^{\prime}\parallel Q^{\prime}).\qed$
+The next lemma asserts that, to check if a symmetric PLD dominates another, it suffices to check the hockey-divergence of order $x\geq 1$.
+This is unlike the (asymmetric) case where we need to check $x\in(0,1)$ as well.
+This will be convenient for our proofs.
+We defer its (simple) proof to Appendix˜C.
+Lemma 8.
+For any symmetric PLDs $^{\prime}, LLL\succeq L^{\prime}$ iff $h_{x}(L)\geq h_{x}(L^{\prime})$ for all $x\geq 1$.
+We will use the following “Multiplicative Azuma’s Inequality” from (Kuszmaul and Qi, 2021, Corollary 11).
+This is an extension of the classic multiplicative Chernoff bound, with the difference that the distributions $\mathcal{D}_{i}$ can depend on the previous outcomes $W_{1},\dots,W_{i-1}$.
+The only constraint required is that $\sum_{i\in[k]}\mathop{\mathbb{E}}_{W_{i}\sim\mathcal{D}_{i}}[W_{i}]\leq\mu$ where $\mu$ is a prespecified parameter.
+Theorem 16 (Multiplicative Azuma’s Inequality Kuszmaul and Qi (2021)).
+Let $\mu>0$.
+Suppose that an adversary constructs random variables $W_{1},\dots,W_{k}\in[0,1]$ via the following process: When the outcomes of $W_{1},\dots,W_{i-1}$ are determined, the adversary can select the distribution $\mathcal{D}_{i}$ with mean $\mu_{i}$ for $W_{i}$ under the constraint $\sum_{i\in[k]}\mu_{i}\leq\mu$.
+If $W=\sum_{i}W_{i}$, then for any $\zeta>0$,
+$\Pr[W\geq(1+\zeta)\mu]\leq\exp\left(-\frac{\zeta^{2}}{2+\zeta}\mu\right).$
+It will also be convenient to state the following version of the tail bound for a similar setting as above but when $\mu$ is very small.
+Lemma 9.
+Let the setting be exactly the same as in Theorem˜16. Then, $\Pr[W\geq\mu/\zeta]\leq\zeta$.
+Proof.
+This is simply an application of Markov’s inequality, since $\mathop{\mathbb{E}}[W]\leq\mu$. ∎
+7.3.2 Privacy Analysis of RR: Comparing Different $\varepsilon$’s
+Before we proceed to the main argument, we need some technical results.
+The main lemma below states that, if we take $\tilde{\varepsilon}-$DP randomized response and compose it 12 times, then its pair dominates that of $2\tilde{\varepsilon}-$DP randomized response.
+This result seems interesting on its own and might be useful beyond the context of this work.
+We remark that we did not attempt to optimize the constant “12” here; nevertheless, it is not hard to see that any constant less than 6 does not work (by taking $\varepsilon\to 0$ and looking at the hockey-stick divergence of order $x=1$, i.e.
+the TV distance).
+Lemma 10.
+For any $\tilde{\varepsilon}\geq 0$, $(\mathrm{RR}^{\otimes 12}_{\tilde{\varepsilon}}(0),\mathrm{RR}^{\otimes 12}_{\tilde{\varepsilon}}(1))\succeq(\mathrm{RR}_{2\tilde{\varepsilon}}(0),\mathrm{RR}_{2\tilde{\varepsilon}}(1))$.
+Proof.
+Let $t=12$.
+Since both $(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0),\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))$ and $(\mathrm{RR}_{2\tilde{\varepsilon}}(0),\mathrm{RR}_{2\tilde{\varepsilon}}(1))$ are symmetric, by Lemma˜8, it suffices to show that $H_{x}(\mathrm{RR}_{2\tilde{\varepsilon}}(0)\parallel\mathrm{RR}_{2\tilde{\varepsilon}}(1))\leq H_{x}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0)\parallel\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))$ for any $x\geq 1$.
+To see this is true, consider three cases:
+  • Case I: $x\geq e^{2\tilde{\varepsilon}}$.
+Then, $H_{x}(\mathrm{RR}_{2\tilde{\varepsilon}}(0)\parallel\mathrm{RR}_{2\tilde{\varepsilon}}(1))=0\leq H_{x}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0)\parallel\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))$.
+  • Case II: $x=1$.
+Note that $H_{x}$ is simply the TV distance.
+Thus, $H_{x}(\mathrm{RR}_{2\tilde{\varepsilon}}(0)\parallel\mathrm{RR}_{2\tilde{\varepsilon}}(1))=\frac{e^{2\tilde{\varepsilon}}-1}{e^{2\tilde{\varepsilon}}+1}.$
+    To lower bound $H_{x}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0)\parallel\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))$, let $S_{i}$ denote the set of all tuples in $\{0,1\}^{t}$ such that the number of ones is exactly $i$.
+Let $P=\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0),Q=\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1)$ We thus have
+    $H_{x}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0)\parallel\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))$ $\geq\sum_{i=0}^{t/2-1}{\left({P(S_{i})-Q(S_{i})}\right)}$
+    $=\sum_{i=0}^{t/2-1}\binom{t}{i}\left(\frac{e^{(t-i)\tilde{\varepsilon}}-e^{i\tilde{\varepsilon}}}{(e^{\tilde{\varepsilon}}+1)^{t}}\right)$
+    $=\sum_{i=0}^{t/2-1}\binom{t}{i}\left(\frac{e^{i\tilde{\varepsilon}}{\left({e^{(t-2i)\tilde{\varepsilon}}-1}\right)}}{(e^{\tilde{\varepsilon}}+1)^{t}}\right)$
+    $\geq\frac{e^{2\tilde{\varepsilon}}-1}{(e^{\tilde{\varepsilon}}+1)^{t}}{\left({\binom{t}{t/2-1}e^{(t/2-1)\tilde{\varepsilon}}+\sum_{i=0}^{t/2-2}\binom{t}{i}{\left({e^{i\tilde{\varepsilon}}{\left({e^{(t-2i-2)\tilde{\varepsilon}}+1}\right)}}\right)}}\right)}.$
+    Thus, we have
+    $\frac{H_{x}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0)\parallel\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))}{H_{x}(\mathrm{RR}_{2\tilde{\varepsilon}}(0)\parallel\mathrm{RR}_{2\tilde{\varepsilon}}(1))}$
+    $=\frac{e^{2\tilde{\varepsilon}}+1}{(e^{\tilde{\varepsilon}}+1)^{t}}{\left({\binom{t}{t/2-1}e^{(t/2-1)\tilde{\varepsilon}}+\sum_{i=0}^{t/2-2}\binom{t}{i}{\left({e^{i\tilde{\varepsilon}}{\left({e^{(t-2i-2)\tilde{\varepsilon}}+1}\right)}}\right)}}\right)}$
+    $=\frac{1}{(e^{\tilde{\varepsilon}}+1)^{t}}{\left({\binom{t}{t/2-1}{\left({e^{(t/2+1)\tilde{\varepsilon}}+e^{(t/2-1)\tilde{\varepsilon}}}\right)}+\sum_{i=0}^{t/2-2}\binom{t}{i}{\left({e^{(t-i)\tilde{\varepsilon}}+e^{i\tilde{\varepsilon}}+e^{(t-i-2)\tilde{\varepsilon}}+e^{(i+2)\tilde{\varepsilon}}}\right)}}\right)}$
+    $\geq\frac{1}{(e^{\tilde{\varepsilon}}+1)^{t}}{\left({\binom{t}{t/2-1}{\left({e^{(t/2+1)\tilde{\varepsilon}}+e^{(t/2-1)\tilde{\varepsilon}}}\right)}+\sum_{i=0}^{t/2-2}\binom{t}{i}{\left({e^{(t-i)\tilde{\varepsilon}}+e^{i\tilde{\varepsilon}}}\right)}+2\binom{t}{t/2-2}e^{(t/2)\tilde{\varepsilon}}}\right)}$
+    $=\frac{1}{(e^{\tilde{\varepsilon}}+1)^{t}}{\left({(e^{\tilde{\varepsilon}}+1)^{t}+{\left({2\binom{t}{t/2-2}-\binom{t}{t/2}}\right)}e^{(t/2)\tilde{\varepsilon}}}\right)}$
+    $\geq 1,$
+    where the last inequality is due to our choice $t=12$.
+  • Case III: $x\in{\left({1,e^{2\tilde{\varepsilon}}}\right)}$.
+Since the privacy loss values for both $\operatorname{PLD}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0),\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))$ and $\operatorname{PLD}(\mathrm{RR}_{2\tilde{\varepsilon}}(0),\mathrm{RR}_{2\tilde{\varepsilon}}(1))$ are multiples [^6] Note that for $\operatorname{PLD}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0),\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))$, this follows from the fact that $t$ is even.
+of $2\tilde{\varepsilon}$, the hockey stick curves $H_{x}$ of both pairs are linear for $x\in\left[1,e^{2\tilde{\varepsilon}}\right$].
+Thus, we have
+    $H_{x}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0)\parallel\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))$
+    $=\frac{x-1}{e^{2\tilde{\varepsilon}}-1}\cdot H_{e^{2\tilde{\varepsilon}}}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0)\parallel\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))+\frac{e^{2\tilde{\varepsilon}}-x}{e^{2\tilde{\varepsilon}}-1}\cdot H_{1}(\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(0)\parallel\mathrm{RR}^{\otimes t}_{\tilde{\varepsilon}}(1))$
+    $\geq\frac{x-1}{e^{2\tilde{\varepsilon}}-1}\cdot H_{e^{2\tilde{\varepsilon}}}(\mathrm{RR}_{2\tilde{\varepsilon}}(0)\parallel\mathrm{RR}_{2\tilde{\varepsilon}}(1))+\frac{e^{2\tilde{\varepsilon}}-x}{e^{2\tilde{\varepsilon}}-1}\cdot H_{1}(\mathrm{RR}_{2\tilde{\varepsilon}}(0)\parallel\mathrm{RR}_{2\tilde{\varepsilon}}(1))$
+    $=H_{x}(\mathrm{RR}_{2\tilde{\varepsilon}}(0)\parallel\mathrm{RR}_{2\tilde{\varepsilon}}(1)),$
+    where the inequality follows from Cases I and II above. ∎
+Using the above lemma together with our setting of parameters, it is simple to derive the following, which will be more convenient in a subsequent usage.
+Lemma 11.
+For every $i\in[n]$, $\mathrm{RR}_{\varepsilon_{i}}^{\otimes b^{*}_{i}}$ is $\left(2\lceil\log(1/\delta)\rceil\cdot(9\varepsilon+2),32\lceil\log(1/\delta)\rceil\cdot\delta\right)$-DP.
+Proof.
+Consider three cases:
+  • Case I: $i=n$. This case is obvious since $b^{*}_{i}=0$.
+  • Case II: $i=1$.
+In this case, $\ell^{*}_{i}=k$ and $\varepsilon_{i}\leq\frac{\varepsilon}{k}$.
+Thus, by our setting of $b^{*}_{i}=2\lceil\log(1/\delta)\rceil\cdot\ell^{*}_{i}$ and basic composition, $\mathrm{RR}^{\otimes b_{i}}_{\varepsilon_{i}}$ is $\left({2\lceil\log(1/\delta)\rceil\cdot\varepsilon}\right)$}-DP as desired.
+  • Case III: $1<i<n$.
+From the setting of $\ell^{*}_{i}$ in (1), $\mathrm{RR}_{\varepsilon_{i}}^{\otimes\frac{\ell^{*}_{i}-1}{8}}$ is $(\varepsilon,2\delta)$-DP.
+Since $\varepsilon_{i}\leq\varepsilon+2$, by basic composition, $\mathrm{RR}_{\varepsilon_{i}}^{\otimes\ell^{*}_{i}}$ is $(9\varepsilon+2,16\delta)$-DP.
+Thus, by our setting of $b^{*}_{i}=2\lceil\log(1/\delta)\rceil\cdot\ell^{*}_{i}$ and basic composition, $\mathrm{RR}^{\otimes b_{i}}_{\varepsilon_{i}}$ is $\left(2\lceil\log(1/\delta)\rceil\cdot(9\varepsilon+2),32\lceil\log(1/\delta)\rceil\cdot\delta\right)$-DP as desired.
 ∎
-(a) GDP $G_{\mu}$ filter bound
-(b) Approximate GDP $f_{0,\delta}\otimes G_{\mu}$ filter bound
-Figure 3: Counter example to free $f$-DP filters when $f$ is GDP (a) or approximate GDP (b).
-Corollary 3.
-The set of all tradeoff curves $\mathcal{T}$ does not have free $f$-DP filters even for $f\in\{G_{\mu}:\mu\geq 0\}$.
-The set of $(\varepsilon,\delta)$-DP tradeoff curves $\mathcal{F}=\{f_{\varepsilon,\delta}:\varepsilon\geq 0,\ \delta\in[0,1]\}$ does not have free $f$-DP filters even for $f\in\{f_{0,\delta}\otimes G_{\mu}:\mu\geq 0,\delta\in[0,1]\}$.
+7.3.3 Formulation as Post-Processing of RR
+As alluded to earlier, we will prove Theorem˜15 by showing that, with high probability, the filter’s output is a post-processing of randomized response answers.
+To do this, recall from Section˜2 that it suffices to consider two neighbouring datasets $D_{0},D_{1}$.
+Our post-processing procedure is described in full in Algorithm˜4.
+It takes as input the randomized response outputs on a bit $h\in\{0,1\}$, and its output (approximately) matches the output of the additive filter.
+To formalize this, let $M^{\mathrm{discrete}}_{\mathcal{A}}(D)$ denote the output distribution of the Additive Discretized Privacy Loss Filter (Algorithm˜3) on input dataset $D$ with analyst $\mathcal{A}$ (where other parameters are omitted from the notation for readability).
+Algorithm 4 Post-processing Algorithm
+0: Discretization points $\bm{\varepsilon}\in\mathcal{E}_{n}$, budgets $\bm{\ell}\in\mathbb{R}^{n}_{\geq 0}$, analyst $\mathcal{A}$, query capacity $k$, neighbouring datasets $D_{0},D_{1}$, independent samples $y_{i}^{j}\sim\mathrm{RR}_{\varepsilon_{i}}(h)$ (where $h\in\{0,1\}$)
+$\mathbf{j}\leftarrow\mathbf{0}$ $\triangleright$ $n$-dimensional all-zero vector; indices for current $R$ samples
+for $t\leftarrow 1,\dots,k$ do
+  $\mathcal{A}(Y_{1},\dots,Y_{t-1})$ gives mechanism $M_{t}$ that has PLD $L_{t}\in\mathcal{L}_{\mathrm{sym}}$
+  $\mathbf{p}^{t}\leftarrow$ Discretized privacy loss profile of $L_{t}$ w.r.t. $\bm{\varepsilon}$
+  $\bm{\ell}\leftarrow\bm{\ell}-\mathbf{p}^{t}$
+  $\kappa_{t}\leftarrow$ Markov kernel with $\kappa_{t}{\left({\mathrm{RR}_{\mathbf{p}^{t},\bm{\varepsilon}}(h^{\prime})}\right)}=M_{t}(D_{h^{\prime}})$ for $h^{\prime}\in\{0,1\}$ $\triangleright$ exists due to Theorem˜11
+  if $\bm{\ell}\geq\mathbf{0}$ then
+   $i_{t}\leftarrow$ sample according to $\mathbf{p}^{t}$
+   $j_{i_{t}}\leftarrow j_{i_{t}}+1$
+   if $j_{i_{t}}>b_{i_{t}}$ then
+    return FAIL $\triangleright$ Exceed budget; fail and terminate the algorithm
+   Sample $Y_{t}\sim\kappa_{t}{\left({i_{t},y^{j_{i_{t}}}}\right)}$
+  else
+   $Y_{t}\leftarrow\bot$
+return $(Y_{1},\ldots,Y_{k})$
+Our post-processing procedure (Algorithm˜4) has the parameter $\mathbf{b}=(b_{1},\dots,b_{n})$ representing the “sample budgets” (i.e.
+how many randomized response samples are allowed to be used).
+To analyze the algorithm, first note that when there is no sample budgets (i.e.
+$b_{1}=\cdots=b_{n}=\infty$), the output is exactly the same as the Additive Discretized Privacy Loss Filter:
+Observation 17.
+When we set $b_{i}=\infty$ for all $i\in[n]$, the output of Algorithm˜4 has exactly the same distribution as $M^{\mathrm{discrete}}_{\mathcal{A}}(D_{h})$.
 Proof.
-Following the reasoning of Corollary˜2, we show that we can construct a set of two tradeoff curves $g_{1},g_{2}\in\mathcal{F}$, and a bound $f$ in the correct family, such that $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})\geq f$ but there exists values of $\alpha$ s.t.
-$g_{1}\otimes\operatorname{conv}(g_{1},g_{2})(\alpha)<f$.
-Figure˜3(a) shows such a construction for $f\in\{G_{\mu}:\mu\geq 0\}$ and Figure˜3(b) for $f\in\{f_{0,\delta}\otimes G_{\mu}:\mu\geq 0,\delta\in[0,1]\}$.
-All trade-off curves are (compositions of) piecewise linear functions, or Gaussian trade-offs.
-They can thus be computed exactly in closed-form.
+Since $b_{i}=\infty$ for all $i\in[n]$, we never output FAIL.
+Thus, for each $t\in[k]$, $(i_{t},y^{j_{i_{t}}})$ has exactly the same distribution as $\mathrm{RR}_{\mathbf{p}^{t},\bm{\varepsilon}}(h)$.
+From the definition of $\kappa_{t}$, $Y_{t}\sim\kappa_{t}(i_{t},y^{j_{i_{t}}})$ has exactly the same distribution as $M_{t}(D_{h})$.
+Thus, the entire process is exactly the same as $M^{\mathrm{discrete}}_{\mathcal{A}}(D_{h})$.
 ∎
-5.2 Families of mechanisms with free filters
-Theorem˜3, combined with Proposition˜10, show that a family of DP mechanisms that is closed under composition, and with well ordered tradeoff curves, admits free $f$-DP (and PLD by Proposition˜10) filters.
-Corollary 4.
-The set of all Gaussian tradeoff curves $G=\{G_{\mu}:\mu\geq 0\}$ has free filters. That is Algorithm˜2 is $f$-DP.
+We next show that, even when we set the budget as in (1), the algorithm barely fails:
+Lemma 12.
+When we set $\ell_{i}=\ell^{*}_{i}$ and $b_{i}=b^{*}_{i}$ for all $i\in[n]$ (where $\bm{\ell}^{*},\mathbf{b}^{*}$ are as specified in (1) and (5), respectively), Algorithm˜4 outputs FAIL with probability at most $4n\cdot\delta$.
 Proof.
-For any $\mu_{1},\mu_{2}\geq 0$, $G_{\mu_{1}}\otimes G_{\mu_{2}}=G_{\sqrt{\mu_{1}^{2}+\mu_{2}^{2}}}\in G$.
-If $\mu_{1}>\mu_{2}$, then $G_{\mu_{1}}\preceq G_{\mu_{2}}$.
-Gaussian mechanisms are closed under composition and well ordered: by Theorem˜3, they have free $f$-DP/PLD filters, as already shown in [koskelaindividual, smith2022fullyadaptivecompositiongaussian].
-∎
-Corollary 5.
-The set of all $0,\delta$-DP tradeoff curves $\mathcal{H}=\{f_{0,\delta}:\delta\in[0,1]\}$ has free filters. That is Algorithm˜2 if $f$-DP.
-Proof.
-For any $\delta_{1},\delta_{2}\geq 0$, $f_{0,\delta_{1}}\otimes f_{0,\delta_{2}}=f_{0,\delta_{1}+\delta_{2}-\delta_{1}\delta_{2}}\in\mathcal{H}$.
-If $\delta_{1}>\delta_{2}$, then $f_{0,\delta_{1}}\preceq f_{0,\delta_{2}}$.
-Pure $\delta$ mechanisms are closed under composition and well ordered: by Theorem˜3, they have free $f$-DP/PLD filters.
-∎
-6 A Proof of Theorem˜3
-In this section, we give a proof of Theorem˜3 that goes through the topology of hockey-stick curves.
-As we have seen, Theorem˜3 can be proved by showing that, for any family $\mathcal{F}$ that is not well-ordered, we can find $g_{1},g_{2}\in\mathcal{F}^{\infty}$ for which
-$g_{1}\otimes\operatorname{conv}\{g_{1},g_{2}\}\prec\operatorname{conv}\{g_{1}\otimes g_{1},g_{1}\otimes g_{2}\}$
-Equivalently, in the language of PLDs, we would like to show that if $L_{1}$ and $L_{2}$ are not well-ordered, then $L_{1}\oplus\sup\{L_{1},L_{2}\}\succ\sup\{L_{1}\oplus L_{1},L_{1}\oplus L_{2}\}$.
+The algorithm outputs FAIL only when $j_{i}>b_{i}$ for some $i\in[n]$.
+Thus, by the union bound, it suffices to show that $\Pr[j_{i}\geq b_{i}]\leq 4\delta$ for every $i\in[n]$.
+To do this, we consider two cases:
+  • Case I: $i<n$.
+Let $\mathcal{D}_{t}$ denote the distribution $\mathrm{Ber}(p^{t}_{i})$ and $W_{t}$ denote $\mathbf{1}[i_{t}=i]$.
+For notational convenience, we also assume that, if the algorithm outputs FAIL in the $t_{0}$-th iteration, then we simply set $p^{t}_{i}=0,\mathcal{D}_{t}=\mathrm{Ber}(0)$ for $t\in\{t_{0}+1,\dots,k\}$.
+    Under these notations, we have $W_{t}\sim\mathcal{D}_{t}$ and $W=W_{1}+\cdots+W_{k}$ is exactly $j_{i}$.
+Finally, due to the constraint of the filter, we have $\mathop{\mathbb{E}}[W_{1}]+\cdots+\mathop{\mathbb{E}}[W_{k}]=p^{1}_{i}+\cdots+p^{k}_{i}\leq\ell^{*}_{i}$.
+Thus, applying Theorem˜16 with $\zeta=6\lceil\log(1/\delta)\rceil-1$ and observing that $\frac{\zeta^{2}}{2+\zeta}\geq\frac{\zeta+1}{6}\geq\lceil\log(1/\delta)\rceil$, we have
+    $\Pr\left[j_{i}>b^{*}_{i}\right]=\Pr[W>(1+\zeta)\cdot\ell^{*}_{i}]\leq\exp(-\log(1/\delta))=\delta.$
+  • Case II: $i=n$. Using the same notation as before and recalling that $b_{i}^{*}=0$ and $\ell^{*}_{i}=4\delta$ in this case, we have
+    $\Pr\left[j_{i}>b^{*}_{i}\right]=\Pr[j_{i}\geq 1]=\Pr[W\geq 1]$ $\leq\ell^{*}_{i}=4\delta,$
+    where the inequality is due to Lemma˜9. ∎
+7.3.4 Putting Things Together: Proof of Theorem˜15
+Combing all the pieces together, we can now prove Theorem˜15.
+Proof of Theorem˜15.
+Let $\varepsilon^{\prime}=n\cdot 2\lceil\log(1/\delta)\rceil\cdot(9\varepsilon+2)$ and $\delta^{\prime}=n\cdot 32\lceil\log(1/\delta)\rceil\cdot\delta+4n\delta$.
+We will show that the $\operatorname{PLD}{\left({\operatorname{Discrete}_{\bm{\varepsilon},\bm{\ell}^{*}}^{k}}\right)}\preceq R_{\varepsilon^{\prime},\delta^{\prime}}$.
+Let $D_{0},D_{1}$ be any neighbouring input datasets.
+From Observation˜17, $M^{\mathrm{discrete}}_{\mathcal{A}}(D_{h})$ is exactly the distribution of the output of Algorithm˜4 with $b_{i}=\infty$.
+Consider also the output distribution of Algorithm˜4 with $b_{i}=b^{*}_{i}$ as defined in (5); let us call the resulting distribution $P^{\prime}(D_{h})$.
+By Lemma˜12, $(P,Q)=(M^{\mathrm{discrete}}_{\mathcal{A}}(D_{0}),M^{\mathrm{discrete}}_{\mathcal{A}}(D_{1}))$ and $(P^{\prime},Q^{\prime})=(P^{\prime}(D_{0}),P^{\prime}(D_{1}))$ satisfies the conditions in Lemma˜7 with $\omega^{*}=\text{FAIL}$ and $\gamma=4n\delta$.
+Thus, Lemma˜7 implies that
+$H_{e^{\varepsilon^{\prime}}}(M^{\mathrm{discrete}}_{\mathcal{A}}(D_{0})\parallel M^{\mathrm{discrete}}_{\mathcal{A}}(D_{1}))\leq H_{e^{\varepsilon^{\prime}}}(P^{\prime}(D_{0})\parallel P^{\prime}(D_{1}))+4n\delta.$
+To bound $H_{e^{\varepsilon^{\prime}}}(P^{\prime}(D_{0})\parallel P^{\prime}(D_{1}))$, notice that, since we set $b_{i}=b^{*}_{i}$ for all $i\in[n]$, $P^{\prime}(D_{h})$ is simply a post-processing of $\mathrm{RR}_{\varepsilon_{1}}^{\otimes b^{*}_{1}}(h)\otimes\cdots\otimes\mathrm{RR}_{\varepsilon_{n}}^{\otimes b^{*}_{n}}(h)$.
+Thus, we have
+$H_{e^{\varepsilon^{\prime}}}(P^{\prime}(D_{0})\parallel P^{\prime}(D_{1}))$ $\leq H_{e^{\varepsilon^{\prime}}}\left(\mathrm{RR}_{\varepsilon_{1}}^{\otimes b^{*}_{1}}(0)\otimes\cdots\otimes\mathrm{RR}_{\varepsilon_{n}}^{\otimes b^{*}_{n}}(0)~\middle\|~\mathrm{RR}_{\varepsilon_{1}}^{\otimes b^{*}_{1}}(1)\otimes\cdots\otimes\mathrm{RR}_{\varepsilon_{n}}^{\otimes b^{*}_{n}}(1)\right).$
+Finally, from Lemma˜11 and basic composition, we can conclude that $\mathrm{RR}_{\varepsilon_{1}}^{\otimes b^{*}_{1}}\otimes\cdots\otimes\mathrm{RR}_{\varepsilon_{n}}^{\otimes b^{*}_{n}}$ is $\left({\varepsilon^{\prime},n\cdot 32\lceil\log(1/\delta)\rceil\cdot\delta}\right)$}-DP.
+In other words,
+$H_{e^{\varepsilon^{\prime}}}\left(\mathrm{RR}_{\varepsilon_{1}}^{\otimes b^{*}_{1}}(0)\otimes\cdots\otimes\mathrm{RR}_{\varepsilon_{n}}^{\otimes b^{*}_{n}}(0)~\middle\|~\mathrm{RR}_{\varepsilon_{1}}^{\otimes b^{*}_{1}}(1)\otimes\cdots\otimes\mathrm{RR}_{\varepsilon_{n}}^{\otimes b^{*}_{n}}(1)\right)\leq n\cdot 32\lceil\log(1/\delta)\rceil\cdot\delta.$
+Combining the three inequalities, we get
+$H_{e^{\varepsilon^{\prime}}}(M^{\mathrm{discrete}}_{\mathcal{A}}(D_{0})\parallel M^{\mathrm{discrete}}_{\mathcal{A}}(D_{1}))\leq n\cdot 32\lceil\log(1/\delta)\rceil\cdot\delta+4n\delta=\delta^{\prime}.$
+Thus, $\operatorname{PLD}{\left({\operatorname{Discrete}_{\bm{\varepsilon},\bm{\ell}^{*}}^{k}}\right)}\preceq R_{\varepsilon^{\prime},\delta^{\prime}}$ as claimed. ∎
+8 A Proof of Lemma˜4
+In this section, we give a proof of the key technical Lemma˜4 that goes through the topology of hockey-stick curves.
+That is, we are given non-degenerate PLDs $L_{1}$ and $L_{2}$ that are not ordered.
+Our goal is to argue that, for one of these two PLDs $L_{i}$, $L_{i}\oplus\sup\{L_{1},L_{2}\}\succ\sup\{L_{i}\oplus L_{1},L_{i}\oplus L_{2}\}$.
 Our first step is to show that gaps between PLDs can be amplified in some sense by composition.
-For a distribution $A$ in the reals, we denote by $\operatorname{supp}(A):=\{t\in\mathbb{R}:\forall\text{open }U\ni t,A(U)>0\}$ the support of $A$, namely the points with locally positive probability mass.
+Recall that the support of a distribution $A$ over $\overline{\mathbb{R}}$, is $\operatorname{supp}(A):=\{t\in\mathbb{R}:\forall\text{open }U\ni t,A(U)>0\}$, i.e.
+the finite points with locally positive probability.
 It turns out that any PLD whose support can be shifted to hit gaps between other PLDs causes the gaps to align when composed.
 Proposition 13 (PLD Gap Amplification).
-Let $A,B,B_{1},B_{2}$ be PLDs so that $B_{i}\prec B$ with $h_{B_{i}}(e^{\varepsilon_{i}+u})<h_{B}(e^{\varepsilon_{i}+u})$ for some $\varepsilon_{1},\varepsilon_{2}\in\operatorname{supp}(A)$ and some $u\in\mathbb{R}$.
+Let $A,L,L_{1},L_{2}$ be PLDs so that $L_{i}\prec L$ with $h_{L_{i}}(e^{\varepsilon_{i}+u})<h_{L}(e^{\varepsilon_{i}+u})$ for some $\varepsilon_{1},\varepsilon_{2}\in\operatorname{supp}(A)$ and some $u\in\mathbb{R}$.
 Then
-$A\oplus B\succ\sup\{A\oplus B_{1},A\oplus B_{2}\}.$
+$A\oplus L\succ\sup\{A\oplus L_{1},A\oplus L_{2}\}.$
 Proof.
-First, it is clear that $A\oplus B\succeq\sup\{A\oplus B_{1},A\oplus B_{2}\}$, so we just need to find a single gap.
-To that end, by continuity, we can find a constant $\rho>0$ s.t $\inf_{\varepsilon\in(\varepsilon_{i}+u\pm\rho)}h_{B}(e^{\varepsilon})-h_{B_{i}}(e^{\varepsilon})>0$.
+First, it is clear that $A\oplus L\succeq\sup\{A\oplus L_{1},A\oplus L_{2}\}$, so we just need to find a single gap.
+To that end, by continuity, we can find a constant $\rho>0$ s.t $\inf_{\varepsilon\in(\varepsilon_{i}+u\pm\rho)}h_{L}(e^{\varepsilon})-h_{L_{i}}(e^{\varepsilon})>0$.
 Now, set $\varepsilon^{\star}:=\varepsilon_{1}+\varepsilon_{2}+u$ and notice that
-$h_{A\oplus B}(e^{\varepsilon^{\star}})-h_{A\oplus B_{i}}(e^{\varepsilon^{\star}})$ $=\mathbb{E}_{Z\sim A\oplus B}[(1-e^{\varepsilon^{\star}-Z})_{+}]-\mathbb{E}_{Z\sim A\oplus B_{i}}[(1-e^{\varepsilon^{\star}-Z})_{+}]$
-$=\mathbb{E}_{Z_{1}\sim A}[\mathbb{E}_{Z_{2}\sim B}[(1-e^{(\varepsilon^{\star}-Z_{1})-Z_{2}})_{+}]-\mathbb{E}_{Z_{2}\sim B_{i}}[(1-e^{(\varepsilon^{\star}-Z_{1})-Z_{2}})_{+}]]$
-$=\mathbb{E}_{Z\sim A}[h_{B}(e^{\varepsilon^{\star}-Z})-h_{B_{i}}(e^{\varepsilon^{\star}-Z})]$
-$\geq\mathbb{P}_{Z\sim A}(Z\in(\varepsilon_{2-i}\pm\rho))\cdot\inf_{\varepsilon\in(\varepsilon_{2-i}\pm\rho)}h_{B}(e^{\varepsilon_{1}+\varepsilon_{2}+u-\varepsilon})-h_{B_{i}}(e^{\varepsilon_{1}+\varepsilon_{2}+u-\varepsilon})$
-$=\mathbb{P}_{Z\sim A}(Z\in(\varepsilon_{2-i}\pm\rho))\cdot\inf_{\varepsilon\in(\varepsilon_{i}+u\pm\rho)}h_{B}(e^{\varepsilon})-h_{B_{i}}(e^{\varepsilon})$
+$h_{A\oplus L}(e^{\varepsilon^{\star}})-h_{A\oplus L_{i}}(e^{\varepsilon^{\star}})$ $=\mathbb{E}_{Z\sim A\oplus L}[(1-e^{\varepsilon^{\star}-Z})_{+}]-\mathbb{E}_{Z\sim A\oplus L_{i}}[(1-e^{\varepsilon^{\star}-Z})_{+}]$
+$=\mathbb{E}_{Z_{1}\sim A}[\mathbb{E}_{Z_{2}\sim L}[(1-e^{(\varepsilon^{\star}-Z_{1})-Z_{2}})_{+}]-\mathbb{E}_{Z_{2}\sim L_{i}}[(1-e^{(\varepsilon^{\star}-Z_{1})-Z_{2}})_{+}]]$
+$=\mathbb{E}_{Z\sim A}[h_{L}(e^{\varepsilon^{\star}-Z})-h_{L_{i}}(e^{\varepsilon^{\star}-Z})]$
+$\geq\mathbb{P}_{Z\sim A}(Z\in(\varepsilon_{2-i}\pm\rho))\cdot\inf_{\varepsilon\in(\varepsilon_{2-i}\pm\rho)}h_{L}(e^{\varepsilon_{1}+\varepsilon_{2}+u-\varepsilon})-h_{L_{i}}(e^{\varepsilon_{1}+\varepsilon_{2}+u-\varepsilon})$
+$=\mathbb{P}_{Z\sim A}(Z\in(\varepsilon_{2-i}\pm\rho))\cdot\inf_{\varepsilon\in(\varepsilon_{i}+u\pm\rho)}h_{L}(e^{\varepsilon})-h_{L_{i}}(e^{\varepsilon})$
 $>0.$
-In particular, $h_{\sup\{A\oplus B_{1},A\oplus B_{2}\}}(e^{\varepsilon^{\star}})=\max\{h_{A\oplus B_{1}}(e^{\varepsilon^{\star}}),h_{A\oplus B_{2}}(e^{\varepsilon^{\star}})\}<h_{A\oplus B}(e^{\varepsilon^{\star}})$.
+In particular, $h_{\sup\{A\oplus L_{1},A\oplus L_{2}\}}(e^{\varepsilon^{\star}})=\max\{h_{A\oplus L_{1}}(e^{\varepsilon^{\star}}),h_{A\oplus L_{2}}(e^{\varepsilon^{\star}})\}<h_{A\oplus L}(e^{\varepsilon^{\star}})$.
 ∎
-In order to apply gap amplification to PLDs that are not well-ordered i.e.
+In order to apply gap amplification to PLDs that are not totally-ordered i.e.
 $L_{1},L_{2}\prec\sup\{L_{1},L_{2}\}$, we would like to reduce the problem to geometric and topological properties of convex curves over $\mathbb{R}^{\times}:=(0,\infty)$.
@@
 Our problem is now purely geometric.
-Given decreasing, convex, unordered curves $h_{1},h_{2}:\mathbb{R}^{\times}\to[0,1]$ that meet in the limit $x\to 0$, we would like to show that one of them can be scaled along the x-axis so that its elbows hit a region where $h_{1}<h_{2}$ and another region where $h_{1}>h_{2}$.
+Given decreasing, convex, unordered curves $h_{1},h_{2}:\mathbb{R}^{\times}\to[0,1]$ that meet in the limit $x\to 0$ (recall Proposition˜4), we show that one of them can be scaled along the x-axis so that its elbows hit a region where $h_{1}<h_{2}$ and another region where $h_{1}>h_{2}$.
 We proceed by analysis of these regions’ topology within the ambient space $\mathbb{R}^{\times}$.
@@
 ∎
-Critically, non-trivial closed and divisible subsets of $\mathbb{R}^{\times}$ have a highly regular multiplicative structure enclosed byt their divisors.
+Critically, non-trivial closed and divisible subsets of $\mathbb{R}^{\times}$ have a highly regular multiplicative structure enclosed by their divisors.
 Proposition 15.
@@
 ∎
-We now can prove our main result of the section, which implies Theorem˜3.
+We now can prove our main result of the section, which implies Lemma˜4.
 We note the small caveat that the queries must be non-degenerate, namely their PLDs must be supported at at least two finite points.
@@
 Theorem 18.
-Let $h_{1},h_{2}:\mathbb{R}^{\times}\to\mathbb{R}$ be non-degenerate hockey-stick curves that are not well-ordered.
+Let $h_{1},h_{2}:\mathbb{R}^{\times}\to\mathbb{R}$ be non-degenerate hockey-stick curves that are not totally-ordered.
 Then one of $\operatorname{supp}(h_{1}),\operatorname{supp}(h_{2})$ links $X(h_{1},h_{2})$ and $X(h_{2},h_{1})$.
@@
 By Proposition 16, $\operatorname{supp}(h_{2})$ meets $\partial X(h_{1},h_{2})\subseteq\overline{X(h_{1},h_{2})}$.
-Furthermore, since $h_{1}$ and $h_{2}$ are not well-ordered, $X(h_{2},h_{1})\neq\varnothing,\mathbb{R}^{\times}$ and thus $\operatorname{supp}(h_{2})$ meets $X(h_{2},h_{1})$ as well by Proposition 18.
+Furthermore, since $h_{1}$ and $h_{2}$ are not totally-ordered, $X(h_{2},h_{1})\neq\varnothing,\mathbb{R}^{\times}$ and thus $\operatorname{supp}(h_{2})$ meets $X(h_{2},h_{1})$ as well by Proposition 17.
 By Lemma 13, $\operatorname{supp}(h_{2})$ must link $X(h_{1},h_{2})$ and $X(h_{2},h_{1})$.
@@
 On the other hand, a homogeneous linear recurrence over the integers admits a bounded non-constant solution if and only its characteristic polynomial has a non-unit root lying on the complex unit circle (see e.g.
-Section 2.3 of [elaydi2005introduction]).
-Therefore either $y_{k}=h_{1}(a^{k}b_{1})=h_{2}(a^{k}b_{1})$ is unbounded, which is impossible because hockey-stick curves are bounded, or it must be constant, which is also impossible because hockey-stick curves are monotone, so this would force $h_{1}$ and $h_{2}$ to be constant and thus well-ordered.
+Section 2.3 of Elaydi (2005)).
+Therefore either $y_{k}=h_{1}(a^{k}b_{1})=h_{2}(a^{k}b_{1})$ is unbounded, which is impossible because hockey-stick curves are bounded, or it must be constant, which is also impossible because hockey-stick curves are monotone, so this would force $h_{1}$ and $h_{2}$ to be constant and thus totally-ordered.
 ∎
-Theorem˜3 now follows immediately by combining Propositions 14 and 15 with Theorem 4 by taking $B:=\sup\{B_{1},B_{2}\}$.
-Appendix A $f$-DP Results, Proofs and Special Cases
-We present additional results focusing on $f$-DP, which we use to instantiate our main results Theorems˜1, 2 and 3 on special cases, that may help with intuition.
-For completeness, Algorithm˜2 states the privacy filter algorithm for $f$-DP, and is a direct analogy to Algorithm˜1.
-Algorithm 2 NaturalFilter ($f$-DP)
-0: Privacy bound $f$, adversary (or analyst) $\mathcal{A}$, family of tradeoff curves $\mathcal{F}$, query capacity $k$
-$i=0$
-while $\mathcal{A}$ continues and $i<k$ do
-  $i\leftarrow i+1$
-  $\mathcal{A}$ gives mechanism $\mathcal{M}_{i}$ that is $f_{i}$-DP, with $f_{i}\in\mathcal{F}$ $\triangleright$ can depend on previous results $Y_{<i}$
-  if $f_{1}\otimes\ldots\otimes f_{i}\geq f$ then
-   $\mathcal{A}$ receives $Y_{i}\sim\mathcal{M}_{i}$
-  end if
-end while
-return $(Y_{1},\ldots,Y_{i})$
-A.1 Proof of Proposition˜12
-Proof.
+Lemma˜4 now follows immediately by combining Propositions 13 and 14 with Theorem 18 by taking $L:=\sup\{L_{1},L_{2}\}$.
+Appendix A Proofs from Preliminaries
+Proof of Proposition˜5.
+For any pair $(P,Q)$ defined on $\Omega$ with likelihood ratio $\ell$, we have that
+$H_{x}(P\parallel Q)$ $=\sup_{E\subseteq\Omega}P(E)-xQ(E)$
+$=\sup_{E\subseteq\Omega}\int_{E}\frac{d(P-xQ)}{dP}(\omega)\,dP(\omega)$
+$=\sup_{E\subseteq\Omega}\int_{E}1-x/\frac{dP}{dQ}(\omega)\,dP(\omega)$
+$=\int_{\Omega}\left(1-x/\frac{dP}{dQ}(\omega)\right)_{+}\,dP(\omega)$
+$=\int_{\Omega}(1-xe^{-\log(\ell(\omega))})_{+}\,dP(\omega)$
+$=\mathop{\mathbb{E}}_{Z\sim\operatorname{PLD}(P\parallel Q)}[(1-xe^{-Z})_{+}].$
+Proof of Proposition˜6.
+Let $\ell$ denote the likelihood ratio for $M_{1}\otimes M_{2}$ and let $\ell^{2}_{y_{1}}$ denote the likelihood ratio for $M_{2}(\cdot;y_{1})$.
+By Bayes’ rule, we have $\ell(y_{1},y_{2})=\ell^{1}(y_{1})\cdot\ell^{2}_{y_{1}}(y_{2})$, so we have
+$\operatorname{PLD}(M_{1}\otimes M_{2})\equiv\log(\ell^{1}(Y_{1}))+\log(\ell^{2}_{Y_{1}}(Y_{2})),Y_{1}\sim M_{1}(D_{1}),Y_{2}\sim M_{2}(D_{1};Y_{1}).$
+By Proposition˜5 and the law of total expectation, it follows that
+$h_{M_{1}\otimes M_{2}}(x)$ $=\mathop{\mathbb{E}}_{Z\sim\operatorname{PLD}(M_{1}\otimes M_{2})}[(1-xe^{-Z})_{+}]$
+$=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[\mathop{\mathbb{E}}_{Y_{2}\sim M_{2}(D_{1};Y_{1})}[(1-xe^{-(\log(\ell^{1}(Y_{1}))+\log(\ell^{2}_{Y_{1}}(Y_{2})))})_{+}]]$
+$=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[\mathop{\mathbb{E}}_{Y_{2}\sim M_{2}(D_{1};Y_{1})}[(1-x/\ell^{1}(Y_{1})\cdot e^{-\log(\ell^{2}_{Y_{1}}(Y_{2}))})_{+}]]$
+$=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[\mathop{\mathbb{E}}_{Z_{2}\sim\operatorname{PLD}(M_{2}(\cdot;Y_{1})}[(1-x/\ell^{1}(Y_{1})\cdot e^{-Z_{2}})_{+}]]$
+$=\mathop{\mathbb{E}}_{Y_{1}\sim M_{1}(D_{1})}[h_{M_{2}(\cdot;Y_{1})}(x/\ell^{1}(Y_{1}))].$
+Proof of Proposition˜9.
+First, let $L$ denote the PLD of $(P,Q)$ and let $F$ be its CDF.
+For $\alpha\in[0,1]$, let $\psi_{\alpha}$ denote the threshold as in Proposition˜8.
+For any $x\in\mathbb{R}^{\times}$, Proposition˜5 yields
+$h(x)$ $=\mathop{\mathbb{E}}_{Z\sim L}[1(Z>\log{x})\cdot(1-xe^{-Z})]$
+$=\sup_{\alpha\in[0,1]}\mathop{\mathbb{E}}_{Z\sim L}[\psi_{\alpha}(Z)\cdot(1-xe^{-Z})]$ ($\alpha=F(\log{x})$)
+$=\sup_{\alpha\in[0,1]}1-\alpha-x\tau(\alpha)$
+$=\sup_{\beta\in[0,1]}1-\tau^{-1}(\beta)-x\beta$ ($\alpha=\tau^{-1}(\beta),\beta=\tau(\alpha)$)
+$=1+(\tau^{-1})^{*}(-x).$
+Appendix B Proofs for Section 6
+Proof of Proposition˜12.
 For each $i\in\{1,2\}$, let $f_{i}$ be a tradeoff curve with $n_{i}$ linear segments and $f_{i}(0)=1-\delta_{i}$.
@@
 The width of interval $j$ is denoted by $w_{i,j}$.
-We also assume $f^{\prime}_{i,1}\geq-f^{\prime}_{i,2}\geq\cdots\geq-f^{\prime}_{i,n_{i}}$.
+By convexity, we have $f^{\prime}_{i,1}\geq-f^{\prime}_{i,2}\geq\cdots\geq-f^{\prime}_{i,n_{i}}$.
 Note that we have $\alpha_{i,1}=0$, $\alpha_{i,n_{i}+1}=1$, $f_{i}(1)=0$, and $\sum_{j=1}^{n_{i}}w_{i,j}=1$.
@@
     $\begin{array}[]{l}\alpha=\sum_{s=1}^{j}\tilde{w}_{s}+c\cdot\tilde{w}_{j+1}\quad\Rightarrow c\cdot\tilde{w}_{j+1}=\alpha-\sum_{s=1}^{j}\tilde{w}_{s},\\ T(P\times P,Q_{1}\times Q_{2})(\alpha)=1-\delta_{1}-\delta_{2}-\delta_{1}\delta_{2}-\sum_{s=1}^{j}\tilde{w}_{s}\cdot\tilde{f}^{\prime}_{s}-c\cdot\tilde{w}_{j+1}\cdot\tilde{f}^{\prime}_{j+1}\\ =1-\delta_{1}-\delta_{2}-\delta_{1}\delta_{2}-\sum_{s=1}^{j}\tilde{w}_{s}\cdot\tilde{f}^{\prime}_{s}-(\alpha-\sum_{s=1}^{j}\tilde{w}_{s})\cdot\tilde{f}^{\prime}_{j+1}.\end{array}$
-A.2 A Simple Derivation of Theorem˜3 for Approximate DP
-Definition 8 (Crossing tradeoff curves).
-We say that two tradeoff curves $f,g\in\mathcal{T}$ cross if there exists $\alpha_{1},\alpha_{2}\in[0,1]$ s.t.:
-$f(\alpha_{1})>g(\alpha_{1})\ \textrm{and}\ f(\alpha_{2})<g(\alpha_{2}).$
-By continuity, there has to be $\alpha\in(\alpha_{1},\alpha_{2})$ s.t.
-$f(\alpha)=g(\alpha)$.
-For any two $(\varepsilon,\delta)$-DP with tradeoff curves that cross, there cannot be free $f$-DP filters.
-We can now prove Proposition˜13, a special case of Theorem˜3 for $(\varepsilon,\delta)$-DP.
+Appendix C On Symmetric PLDs and Dominating Pairs
+For the purpose of accurate privacy accounting, we often need to identify the “worst” pairs of datasets and its corresponding privacy loss.
+This concept can be formalized via the notion of (worst-case) dominating pairs as follows:
+Definition 16 (Dominating Pair [58]).
+$(P,Q)$ is a dominating pair for a mechanism $\mathcal{M}$ if $(P,Q)\succeq(\mathcal{M}(D),\mathcal{M}(D^{\prime}))$ for all neighbouring datasets $D,D^{\prime}$.
+Furthermore, it is said to be a worst-case dominating pair if additionally $P=\mathcal{M}(D),Q=\mathcal{M}(D^{\prime})$ for some neighbouring datasets $D,D^{\prime}$.
+In practice, a typical privacy accounting is done by taking each $L_{i}$ to be the PLD of a worst-case dominating pair of the mechanism $\mathcal{M}_{i}$.
+We observe below that such PLD is symmetric.
+Thus, the assumption in Section˜7 is quite a natural one that is already satisfied in practice.
+Lemma 15.
+If $(P,Q)$ is a worst-case dominating pair of $\mathcal{M}$, then $(P,Q)$ is symmetric.
 Proof.
-Consider $g_{1}\triangleq f_{\varepsilon_{1},\delta_{1}}$ and $g_{2}\triangleq f_{\varepsilon_{2},\delta_{2}}$ that cross.
-By symmetry w.r.t.
-the $y=x$ axis, they have to cross at an $\alpha<\min\{\alpha_{1}^{\star},\alpha_{2}^{\star}\}$, where $\alpha_{i}^{\star}\ \textrm{s.t.}\ \alpha_{i}^{\star}=g_{i}(\alpha_{i}^{\star})$ (otherwise they would never cross).
-This means that their linear segments cross: one function starts strictly lower, and then ends-up strictly higher by $\min\{\alpha_{1}^{\star},\alpha_{2}^{\star}\}$ (again, otherwise they would never cross).
-Without loss of generality, we reorder functions so that $g_{2}$ starts strictly lower.
-We now know that $\delta_{2}>\delta_{1}\geq 0$ ($g_{2}$ starts lower); $\varepsilon_{1}>\varepsilon_{2}\geq 0$ ($g_{1}$ has a steeper slope to catch-up to $g_{2}$ so that they can cross); and $\alpha_{2}^{\star}>\alpha_{1}^{\star}$.
-The last inequality holds because $g_{1}$ is strictly lower by the time is crosses $y=x$ at $\alpha_{1}^{\star}$, so at that point $g_{2}(\alpha_{1}^{\star})>\alpha_{1}^{\star}$, and by continuity of $g_{2}$ we have that $\alpha_{2}^{\star}>\alpha_{1}^{\star}$.
-Since we are dealing with approximate DP curves, we also know that $\alpha_{i}^{\star}=\frac{1-\delta_{i}}{1+e^{\varepsilon_{i}}}$.
-Now, in Theorem˜2, pick $\mathcal{H}=\{g_{1},g_{2}\}\subseteq\mathcal{F}$ and $g=g_{1}\in\mathcal{H}\subseteq\mathcal{F}$.
-We make a few useful observations about the first segment of $g_{1}\otimes g_{1}$, $g_{1}\otimes g_{2}$, and $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$, by applying Proposition˜12:
-The first segment of $g_{1}\otimes g_{1}$ starts at $(1-\delta_{1})^{2}$, with slope $e^{2\varepsilon_{1}}$ for $\alpha\in[0,\alpha_{1}^{\star 2}]$.
-The second slope is $e^{\varepsilon_{1}}e^{-\varepsilon_{1}}=-1$.
-The first segment of $g_{1}\otimes g_{2}$ starts at $(1-\delta_{1})(1-\delta_{2})$, with slope $e^{\varepsilon_{1}+\varepsilon_{2}}$ for $\alpha\in[0,\alpha_{1}^{\star}\alpha_{2}^{\star}]$.
-The second slope is $e^{\varepsilon_{1}-\varepsilon_{2}}<-1$.
-The first segment of $\operatorname{conv}(g_{1},g_{2})$ starts at $(1-\delta_{2})$ (the lowest point at $\alpha=0$), and since $g_{1}$ and $g_{2}$ cross, it meets $g_{1}$ at $\alpha=\alpha_{1}^{\star}$, where its value is $g_{1}(\alpha_{1}^{\star})=1-\delta_{1}-e^{\varepsilon_{1}}\alpha_{1}^{\star}$.
-As a result, the first slope of $\operatorname{conv}(g_{1},g_{2})$ is $\frac{1-\delta_{1}-e^{\varepsilon_{1}}\alpha_{1}^{\star}-(1-\delta_{2})}{\alpha_{1}^{\star}}=\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}-e^{\varepsilon_{1}}$.
-The first segment of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$ starts at $(1-\delta_{1})(1-\delta_{2})$, with first slope $\frac{\delta_{1}-\delta_{2}}{\alpha_{1}^{\star}}e^{\varepsilon_{1}}-e^{2\varepsilon_{1}}$ for $\alpha\in[0,\alpha_{1}^{\star 2}]$.
-Let us study the first segment of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$ and compare it to that of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$.
-We first reason about the possible shapes of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$.
-We saw that $g_{1}\otimes g_{1}(0)=(1-\delta_{1})^{2}>(1-\delta_{1})(1-\delta_{2})=g_{1}\otimes g_{2}(0)$, so $g_{1}\otimes g_{2}$ starts lower.
-Since $e^{2\varepsilon_{1}}<-e^{\varepsilon_{1}+\varepsilon_{2}}<-e^{\varepsilon_{1}-\varepsilon_{2}}$, $g_{1}\otimes g_{1}<-1$, we know that $g_{1}\otimes g_{1}$ starts with a steepest slope, which stops first (at $\alpha_{1}^{\star 2}<\alpha_{1}^{\star}\alpha_{2}^{\star}$) and then gets the smallest second slope (until at least $\alpha\geq\alpha_{g_{1}\otimes g_{2}}^{\star}$).
-As a result, either that first slope is steep enough such that $g_{1}\otimes g_{1}(\alpha_{1}^{\star 2})<g_{1}\otimes g_{2}(\alpha_{1}^{\star 2})$ and the composed curves cross, or they will never cross and $g_{1}\otimes g_{2}\leq g_{1}\otimes g_{1}$.
-We treat each case in turn.
-Case 1: $g_{1}\otimes g_{1}(\alpha_{1}^{\star 2})<g_{1}\otimes g_{2}(\alpha_{1}^{\star 2})$ and the composed curves cross (Figure˜1(a)).
-In that case, we know that $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})(\alpha_{1}^{\star 2})=g_{1}\otimes g_{1}(\alpha_{1}^{\star 2})=(1-\delta_{1})^{2}-e^{2\varepsilon_{1}}\alpha_{1}^{\star 2}$.
-As a result, the slope of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$ in $\alpha\in[0,\alpha_{1}^{\star 2}]$ is $\frac{(1-\delta_{1})^{2}-e^{2\varepsilon_{1}}\alpha_{1}^{\star 2}-(1-\delta_{1})(1-\delta_{2})}{\alpha_{1}^{\star 2}}=\frac{(1-\delta_{1})(\delta_{2}-\delta_{1})}{\alpha_{1}^{\star 2}}-e^{2\varepsilon_{1}}=(1+e^{\varepsilon_{1}})\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}-e^{2\varepsilon_{1}}$ (the last step uses $\alpha_{i}^{\star}=\frac{1-\delta_{i}}{1+e^{\varepsilon_{i}}}$).
-Remember that $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})(0)=(1-\delta_{1})(1-\delta_{2})=\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$, and the first slope of with first slope of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})=\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}e^{\varepsilon_{1}}-e^{2\varepsilon_{1}}<(1+e^{\varepsilon_{1}})\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}-e^{2\varepsilon_{1}}$ since $\delta_{2}-\delta_{1}>0$.
-We have shown that in this case, $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})>\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$ at least on $(0,\alpha_{1}^{\star 2}]$, so by Theorem˜2, and there are no $f$-DP filters.
-Case 2: $g_{1}\otimes g_{1}(\alpha_{1}^{\star 2})\geq g_{1}\otimes g_{2}(\alpha_{1}^{\star 2})$ and $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})=g_{1}\otimes g_{2}$ (Figure˜1(b)).
-Intuitively, we know that the first slope of $\operatorname{conv}(g_{1},g_{2})$ is steeper than that of $g_{2}$, since $g_{2}$ starts lower and crosses with $g_{1}$.
-Formally, the first slope of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$ is $\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}e^{\varepsilon_{1}}-e^{2\varepsilon_{1}}=-e^{\varepsilon_{1}}\big(e^{\varepsilon_{1}}-\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}\big)$.
-We know that $1-\delta_{1}-e^{\varepsilon_{1}}\alpha_{1}^{\star}=g_{1}(\alpha_{1}^{\star})<g_{2}(\alpha_{1}^{\star 2})=1-\delta_{2}-e^{\varepsilon_{2}}\alpha_{1}^{\star}$, so $\delta_{2}-\delta_{1}<\alpha_{1}^{\star}(e^{\varepsilon_{1}}-e^{\varepsilon_{2}})$ and $0<e^{\varepsilon_{2}}<e^{\varepsilon_{1}}-\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}$.
-That is, $e^{\varepsilon_{1}}\big(e^{\varepsilon_{1}}-\frac{\delta_{2}-\delta_{1}}{\alpha_{1}^{\star}}\big)<-e^{\varepsilon_{1}}e^{\varepsilon_{2}}$.
-In this case as well, the first slope of $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})$ is strictly more steep than that of $\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$, so $g_{1}\otimes\operatorname{conv}(g_{1},g_{2})>\operatorname{conv}(g_{1}\otimes g_{1},g_{1}\otimes g_{2})$ at least on $(0,\alpha_{1}^{\star 2}]$.
-By Theorem˜2, and there are no free $f$-DP filters.
+Since $(P,Q)$ is a worst-case dominating pair, we have $P=\mathcal{M}(D)$ and $Q=\mathcal{M}(D^{\prime})$ for some neighbouring datasets $D,D^{\prime}$ and $(P,Q)\succeq(\mathcal{M}(D^{\prime}),\mathcal{M}(D))=(Q,P)$.
+By definition, there exists a Markov kernel $\kappa$ such that $\kappa P=Q$ and $\kappa Q=P$, but this also implies that $(Q,P)\succeq(P,Q)$.
 ∎
+We will next prove Lemma˜8. To do so, let us start by stating a relationship between the hockey-stick divergence at $x$ and $x^{-1}$, which will be convenient for our proof.
+Lemma 16 ([58]).
+For any pair $(P,Q)$ and $x\in\mathbb{R}^{\times}$, we have $H_{x}(Q\parallel P)=x\cdot H_{x^{-1}}(P\parallel Q)-x+1$.
+Proof of Lemma˜8.
+Let $L=\operatorname{PLD}(P,Q),L^{\prime}=\operatorname{PLD}(P^{\prime},Q^{\prime})$ where $(P,Q),(P^{\prime},Q^{\prime})$ are symmetric pairs.
+The forward direction is obvious (from definition of dominating).
+For the backward direction, if $0<x<1$, then Lemma˜16 implies that
+$H_{x}(P\parallel Q)\geq H_{x}(P^{\prime}\parallel Q^{\prime})\Leftrightarrow H_{x^{-1}}(Q\parallel P)\geq H_{x^{-1}}(Q^{\prime}\parallel P^{\prime})\Leftrightarrow H_{x^{-1}}(P\parallel Q)\geq H_{x^{-1}}(P^{\prime}\parallel Q^{\prime}),$
+where the second equivalence is due to the symmetry assumption. ∎